### v1

# On Geometric Evolution and Microlocal Regularity of the Navier–Stokes Equations

Abstract
We develop a geometric and microlocal framework for the Navier–Stokes equations by lifting the dynamics to the cosphere bundle of a Riemannian manifold. In this formulation, the velocity field and vorticity are represented as microlocal distributions whose evolution is governed by a transport–dissipation system generated by a canonical dynamical vector field described with differential forms. We introduce microlocal amplitudes, directional energy functionals, and monotone volume invariants on the compact phase space, which quantify directional concentration and alignment mechanisms associated with potential loss of regularity. The viscous term induces effective geometric diffusion on the cosphere bundle, yielding closed differential inequalities in a geometric setting. To capture the interaction between fluid deformation and geometry, we define an effective connection and curvature tensor encoding the influence of the symmetric velocity gradient. This structure gives rise to a Ricci-type geometric evolution at the microlocal level, controlled by a dissipative functional that excludes extreme directional concentration and persistent transverse stretching. Our results do not resolve the global regularity problem in full generality, but provide a coherent geometric mechanism that severely restricts blow-up scenarios compatible with viscous Navier–Stokes dynamics, reformulating the analytic regularity question as a problem of dissipative stability on a compact phase space.
Keywords
Navier–Stokes equations, microlocal analysis, cosphere bundle, geometric regularity criteria, effective Ricci curvature, dissipative dynamical systems, differential forms.
preprint: APS/123-QED
I Introduction: The Regularity Problem and Geometric Flows
The question about the existence and uniqueness of the solution space for the Navier-Stokes equations is one of the most significant and open research topics in mathematics and theoretical physics [11, 40], with a vast range of applications. Geometric and covariant formulations on Riemannian manifolds provide a coherent framework independent of coordinates [7, 13]. This clearly explains differential operators, the energy structure, the role of curvature, and the symmetries of the system through Killing fields [6]. This is because Hamiltonian formulations and differential forms, using weighted and covariant tensors with exterior derivatives [30], ensure the method works naturally, making diffusion and boundary terms appear more clearly in components with non-zero curvature.
For cases of two spatial dimensions with certain geometries and specific boundary conditions, existence, uniqueness, and continuity are well understood through frequency decomposition and invariant methods [14]. On compact differential manifolds with boundaries, there are strong and weak solutions [39]. Both are critical and converge exponentially to an equilibrium state given a certain evolution regime of the system, matching the Killing vector fields. Consequently, they are associated with invariant quantities that describe asymptotes, providing a model of the solution space near the origin. Similarly, the definition in spectral submanifolds is established in compact domains of the flow [5], with results showing the continuity of these invariants associated with the original differential manifold [36, 32].
In contrast, for methods with three or more spatial dimensions in critical spaces, there is strong evidence of non-uniqueness and geometric indefinitions in the solution space due to vorticities [2, 10]. More specifically, if the vorticity directions stay within a double cone in high-vorticity regions, the solutions remain regular; near singularities, the vorticity directions must spread across the unit sphere [31]. The quantitative classification of potential singularities and Carleman-type inequalities [28] provide testable lower and upper bounds near candidate blow-up times [8]. Classic explicit constructions show a blow-up in finite time in special stagnation-point forms (Euler and Navier-Stokes), driven by unlimited vorticity and infinite domains [9]. As a result, Fourier analysis and microlocal tools adapted to differential manifolds [27, 29] support the proofs of global regularity, geometric trapping [4], and hybrid discretizations in 3D domains [1]. These are based on eigenfunction and multilinear estimates adapted to curved spaces, and are considered methods to provide regularity and restrictions on possible singularities of the system [41, 26].
In this work, we reformulate the incompressible Navier–Stokes equations as a dissipative dynamical system defined globally on the cosphere bundle $S^{*}M$. Our goal is to classify blow-up mechanisms and obtain geometric criteria for conditional regularity using variational and Fourier methods. In this way, we find inequalities and invariants to show how the physical solution is reconstructed from the dynamics in $S^{*}M$.
II Covariant Formulation on Manifolds
To analyze the interaction between fluid dynamics and geometry, it is necessary to reformulate the Navier-Stokes equations in a fully covariant way on an oriented Riemannian differential manifold $(M,g)$. This manifold can be compact without a boundary or have a boundary with appropriate physical conditions, such that its curvatures and derivatives are bounded and the manifold is uniformly injective, equipped with a metric tensor. We denote its tangent and cotangent bundles as $TM$ and $T^{*}M$, respectively. In this section, we leave behind the traditional notation of vector calculus in favor of differential forms and the Hodge Laplacian operator, $\Delta_{H}$, which naturally adapts to the curvature of the space [13, 30]. This reformulation allows us to express momentum conservation and the incompressibility condition as geometric constraints that depend explicitly on the metric tensor. Therefore, the fluid no longer moves through a passive background but instead interacts with a dynamic geometric structure. Usually, for an incompressible fluid, we have the following system of equations using the covariant derivative $\nabla$ given a certain arbitrary connection $\Gamma$ on the Riemannian manifold $(M,g)$ induced by the metric [7].
$\left\{\begin{array}[]{cc}\partial_{t}u+\nabla_{u}u+\nabla p-\nu\Delta u-f=0\\ div(u)=0\end{array}\right$. (1)
Now, let’s consider the 1-forms associated with the velocity field and the external forces acting on the system, denoted as $\phi=\phi^{k}dx_{k}=u^{k}dx_{k}$ and $F=F^{k}dx_{k}$ respectively, with indices such that $k=1,2,\dots,\dim(M,g)$. This allows us to rewrite the original system (1) as follows
$\left\{\begin{array}[]{cc}\partial_{t}\phi+\mathcal{L}_{u}\phi+dP-\nu\Delta_{H}\phi-F=0\\ \delta\phi=0\end{array}\right$. (2)
where
  • $\mathcal{L}_{u}\phi=\iota_{u}d\phi+d(\iota_{u}\phi)=u^{j}(\nabla_{j}\phi_{i})+(\nabla_{i}u^{j})\phi_{j}$ is the Lie derivative of $\phi$ along the direction of $u$.
  • $\Delta_{H}=d\delta+\delta d$ is the Hodge Laplacian, which holds the Ricci tensor and the metric tensor within its definition.
  • $\delta:\Omega^{h}(M)\longmapsto\Omega^{h-1}(M)$ is the Hodge delta. It acts as the adjoint operator to the exterior derivative, such that $\delta=(-1)^{n(p+1)+1}\star d\ \star$ for a differential p-form. When applied to $\phi$, it results in $\delta\phi=-\nabla^{k}\phi_{k}=-g^{kj}\nabla_{j}\phi_{j}$.
Given these definitions and the system of equations we are looking at, if we use the incompressibility equation $\delta\phi=0$, the term for the Lie derivative regroups like this
$\mathcal{L}_{u}\phi=\nabla_{u}u^{k}+\frac{1}{2}d\|u\|^{2}_{L^{2}}$ (3)
Since the norm of $u$ appears here, it is helpful for establishing convergence criteria. As for the Hodge Laplacian, in local coordinates, we can write it as
$\Delta_{H}=(\nabla^{\star}\nabla\phi)_{k}+R^{k}_{ikj}g^{jk}\phi_{k}=\nabla^{i}\nabla_{i}\phi_{k}+R^{k}_{ikj}g^{jk}\phi_{k}$ (4)
By substituting this into the Navier-Stokes equations, we get
$\left\{\begin{array}[]{cc}\partial_{t}\phi+d(\frac{1}{2}\|u\|^{2}_{L^{2}}+P)-\nu(\nabla^{i}\nabla_{i}\phi_{k}+R^{k}_{ikj}g^{jk}\phi_{k})-F=0\\ g^{ik}\nabla_{i}\phi_{i}=0\end{array}\right$. (5)
This formulation makes the system of equations topologically more natural and clearer to analyze. Because of this, we define the covariant gradient of the velocity field
$A:=A_{ik}(dx^{i}\otimes dk^{k})=\nabla_{i}u_{k}(dx^{i}\otimes dk^{k})\in\Gamma(T^{\star}M\otimes TM)$ (6)
as a canonical decomposition induced by $g_{ik}$
$A=S+\Omega$ (7)
where
  • $S\in\Gamma(\text{Sym}^{2}T^{\star}M)$, such that $S_{ik}=\tfrac{1}{2}(\nabla_{i}u_{k}+\nabla_{k}u_{i})$, is a symmetric object by definition, representing the deformation tensor.
  • $\Omega\in\Omega^{2}(M)$ corresponds to the 2-form associated with the fluid’s rotation, defined as $\Omega_{ik}=\tfrac{1}{2}(\nabla_{i}u_{k}-\nabla_{k}u_{i})$.
The antisymmetric part $\Omega$ induces a differential 2-form which coincides, up to a constant, with the exterior derivative of the velocity 1-form. Indeed, by defining
$\omega:=d\phi\in\Omega^{2}(M)$ (8)
we get locally
$\omega_{ik}=\nabla_{i}u_{k}-\nabla_{k}u_{i}=2\Omega_{ik}$ (9)
To explicitly verify the relationship between the antisymmetric part of the gradient and the vorticity, consider the velocity 1-form
$\phi=u_{i}dx^{i}$ (10)
By the definition of the exterior derivative
$d\phi=\partial_{i}u_{j}\,dx^{i}\wedge dx^{j}=\tfrac{1}{2}(\partial_{i}u_{k}-\partial_{k}u_{i})\,dx^{i}\wedge dx^{k}$ (11)
By replacing partial derivatives with covariant derivatives, the connection terms cancel out due to antisymmetry, yielding
$d\phi=\tfrac{1}{2}(\nabla_{i}u_{k}-\nabla_{k}u_{i})\,dx^{i}\wedge dx^{k}$ (12)
Comparing this with the definition of $\Omega$, we conclude that,
$\omega:=d\phi=2\Omega\in\Omega^{2}(M)$ (13)
In arbitrary dimensions, the vorticity is a differential 2-form whose Hodge dual is
$\star\omega\in\Omega^{dim(M)-2}(M)$ (14)
Therefore, in three dimensions, the vorticity can be identified with a 1-form through the Hodge operator
$\xi:=\star\omega\in\Omega^{1}(M)$ (15)
which corresponds to the classical vorticity field. Consequently, the vorticity coincides with the dual of the antisymmetric part of the velocity gradient, confirming that the vortex stretching is fully encoded within the tensor $\Omega$.
Thus, the evolution of the vorticity field can be derived directly from the covariant formulation of the Navier–Stokes equations. By applying the exterior derivative to the momentum equation expressed in terms of the velocity 1-form $\phi$, and utilizing the identity $d^{2}\phi=0$, the evolution equation for the geometric vorticity is obtained [30].
$\partial_{t}\omega+\mathcal{L}_{u}\omega=\nu\Delta_{H}\omega+dF$ (16)
In the absence of external forces, or for conservative forces, the last term vanishes, and the equation adopts a purely geometric form.
In three dimensions, by applying the Hodge operator, this equation translates into an evolution equation for the classical vorticity field $\xi$
$\partial_{t}\xi+\mathcal{L}_{u}\xi=\nu\Delta\xi+(\nabla u)\cdot\xi$ (17)
where the last term represents the vortex stretching, induced exclusively by the symmetric part of the velocity gradient. To describe the dissipative structure of this dynamics, we introduce the vortex energy functional
$\mathcal{E}_{\omega}[\phi]:=\frac{1}{2}\int_{M}\|\omega\|^{2}_{L^{2}}dV=\frac{1}{2}\int_{M}\|d\phi\|^{2}_{L^{2}}dV$ (18)
This functional measures the total intensity of the fluid’s rotation and is invariant under gauge transformations $\phi\mapsto\phi+d\psi$ compatible with the incompressibility condition. Let $\phi_{\varepsilon}=\phi+\varepsilon\psi$, with $\psi\in\Omega^{1}(M)$ being smooth such that $\delta\psi=0$. Then,
$\frac{d}{d\varepsilon}\mathcal{E}_{\omega}[\phi_{\varepsilon}]\Big|_{\varepsilon=0}=\int_{M}\langle d\phi,d\psi\rangle\,dV=\int_{M}\langle\delta d\phi,\psi\rangle\,dV$ (19)
Therefore, the variational gradient of the functional is given by
$\frac{\delta\mathcal{E}_{\omega}}{\delta\phi}=\delta d\phi=\Delta_{H}\phi$ (20)
Consequently, the viscous term of the Navier–Stokes equations can be interpreted as the gradient descent flow of the vortex energy functional with respect to the $L^{2}$ metric induced by $g$. In this sense, the dynamics of an incompressible fluid is presented as a dissipative system whose regularization mechanism is governed by a geometric variational principle. By differentiating with respect to time and using the evolution equation for $\omega$, we obtain the dissipation identity
$\frac{d}{dt}\mathcal{E}_{\omega}(t)=-\nu\int_{M}\|\nabla\omega\|^{2}_{L^{2}}dV\leq\gamma\Big|_{\gamma=0}=0$ (21)
which shows that viscosity imposes a strict monotonicity on the vortex energy, suggesting that the regularity of the system is intimately linked to the existence of dissipative geometric quantities that control the complexity of the flow.
Although the covariant formulation clarifies the geometric structure of the system, the regularity problem remains open in three dimensions. In particular, the vorticity equation reveals that potential amplification is associated with directional stretching, suggesting that the analysis must incorporate not only the magnitude of the vorticity but also its directional structure.
This motivates lifting the problem to a space where cotangent directions can be treated geometrically in a global manner, leading us to the cosphere bundle $S^{*}M$ through the cotangent bundle [20].
III Microlocal Lifting and the Cosphere Bundle
As discussed in the previous section, the natural space for this lift is the cotangent bundle $T^{*}M$, equipped with its canonical symplectic structure given by a projection. Let
$\pi:T^{*}M\longmapsto M,\qquad\pi(x,\xi)=x$ (22)
denote the natural projection, where each fiber $T_{x}^{*}M$ represents the space of co-vectors at the point $x\in M$.
Since the system is invariant under homogeneous rescaling of the cotangent variables, it is convenient to restrict our attention to the co-sphere bundle, defined as
$S^{*}M:=\{(x,\xi)\in T^{*}M:\|\xi\|_{g}=1\}$ (23)
which can be interpreted as the space of unit cotangent directions over $M$. There exists an isomorphism that does not change the homotopy class of the differential manifold considered:
$S^{*}M\cong M\times S^{dim(M)-1}$ (24)
From this perspective, a point in $S^{*}M$ simultaneously encodes a position in physical space and a privileged direction in frequency space. Therefore, the cotangent bundle $T^{*}M$ possesses a canonical symplectic structure given by the 2-form
$\Lambda=d\lambda$ (25)
where $\lambda=\xi_{i}\,dx^{i}$ is a 1-form. This structure is independent of the metric and encodes the microlocal geometry of the system, allowing for the description of the propagation of directional structures, such as vorticity, in a global and geometrically coherent manner.
Consequently, since fluid dynamics is implicitly related to geometry, it is logical to argue that the fluid couples to the metric or forms part of it, shaping the evolution of the manifold. Thus, a natural decomposition of the tangent bundle of $T^{*}M$ into horizontal and vertical subspaces is induced:
$TT^{*}M\simeq H\oplus V$ (26)
where the horizontal distribution is determined by the corresponding connection and the vertical distribution corresponds to purely cotangent variations.
This decomposition allows for the definition of coframes adapted to the microlocal lift and provides a natural geometric basis for lifting differential operators defined on $M$ to the cosphere bundle.
The vertical distribution $V$ is given by
${V}_{(x,\xi)}=\ker(d\pi)_{(x,\xi)}\simeq T_{\xi}(T_{x}^{*}M)$ (27)
While for the horizontal distribution, the connection $\Gamma$ defines it such that $H\subset TT^{*}M$ is characterized by the property that for every $X_{x}\in T_{x}M$, there exists a unique vector $X^{H}_{(x,\xi)}\in T_{(x,\xi)}T^{*}M$ such that
$d\pi\big[X^{H}_{(x,\xi)}\big]=X_{x},\qquad\Gamma_{X_{x}}\xi=0$ (28)
In this way, the decomposition is invariant under isometric diffeomorphisms and provides a natural geometric basis for lifting operators defined on $M$ to $T^{*}M$ and $S^{*}M$. To prove this, let $\varphi:M\longmapsto M$ be an isometric diffeomorphism, that is,
$\varphi^{*}g=g$ (29)
Then the horizontal distribution $H\subset TT^{*}M$ is invariant under the induced diffeomorphism $\tilde{\varphi}$, which implies that
$d\tilde{\varphi}\big[H_{(x,\xi)}\big]=H_{\tilde{\varphi}(x,\xi)}$ (30)
Given $X^{H}_{(x,\xi)}\in H_{(x,\xi)}$, we have $d\pi(X^{H})=X_{x}$ and $\Gamma_{X_{x}}\xi=0$. Since $\varphi$ is an isometry, it preserves the connection, so
$\Gamma_{d\varphi(X_{x})}\big[(d\varphi^{-1})^{*}\xi\big]=0$ (31)
Furthermore,
$d\pi\big[d\tilde{\varphi}(X^{H})\big]=d\varphi(X_{x})$ (32)
By the uniqueness of the horizontal lift, it follows that $d\tilde{\varphi}(X^{H})\in H_{\tilde{\varphi}(x,\xi)}$. Thus, the invariance under isometric diffeomorphisms is proven, as originally intended. Following this, the velocity field $u\in\Gamma(TM)$ canonically induces a dynamical field on $T^{*}M$, defined by
$X_{u}(x,\xi)=u^{i}(x)\,X_{i}^{H}-(\nabla_{k}u^{i})(x)\,\xi_{i}\,\frac{\partial}{\partial\xi_{k}}$ (33)
which simultaneously describes spatial transport and the directional evolution of covectors under the action of the flow [29].
From a geometric perspective, the dynamical vector field $X_{u}$ endows the cotangent bundle with a structure closely analogous to the configuration space of an unquantized field theory [34, 38]. In this interpretation, the microlocal variables $(x,\xi)\in T^{*}M$ play the role of generalized phase-space coordinates, while the flow generated by $X_{u}$ encodes both spatial transport and internal directional evolution.
This viewpoint naturally leads to the introduction of an effective connection adapted to the flow. In particular, the velocity field $u$ induces a deformation of the connection through
$(\Gamma_{u})^{j}_{ik}:=\Gamma^{j}_{ik}+g^{\prime}_{ik}\,u^{j}$ (34)
which incorporates the influence of the fluid transport into the geometric structure of the bundle [3]. Associated with this deformation, we consider an effective metric of the form
$g^{\prime}_{ik}=g_{ik}+\alpha\,\int_{t_{0}}^{t}S_{ik}(s)\ ds,\qquad S_{ik}=\frac{1}{2}\bigl(\nabla_{i}u_{k}+\nabla_{k}u_{i}\bigr)$ (35)
where $\alpha$ is a fixed geometric coupling parameter [21]. This metric deformation encodes the interaction between the background geometry and the fluid deformation tensor.
Under this construction, the curvature tensor associated with the effective connection $\Gamma^{u}$ is given in local coordinates by
$(R^{u})^{l}{}_{ijk}$ $=\partial_{i}(\Gamma^{l}_{jk}+g^{\prime}_{jk}u^{l})-\partial_{j}(\Gamma^{l}_{ik}+g^{\prime}_{ik}u^{l})$
$\quad+(\Gamma^{m}_{jk}+g^{\prime}_{jk}u^{m})(\Gamma^{l}_{im}+g^{\prime}_{im}u^{l})$
$\quad-(\Gamma^{m}_{ik}+g^{\prime}_{ik}u^{m})(\Gamma^{l}_{jm}+g^{\prime}_{jm}u^{l}).$ (36)
This decomposes as
$(R^{u})^{l}{}_{ijk}=R^{l}{}_{ijk}+\mathcal{A}^{l}{}_{ijk}+\mathcal{Q}^{l}{}_{ijk}$, (37)
where $R^{l}{}_{ijk}$ is the Riemann curvature of $g$
$\mathcal{A}^{l}{}_{ijk}=(\nabla_{i}g^{\prime}_{jk}-\nabla_{j}g^{\prime}_{ik})u^{l}+g^{\prime}_{jk}\nabla_{i}u^{l}-g^{\prime}_{ik}\nabla_{j}u^{l}$ (38)
and $\mathcal{Q}^{l}{}_{ijk}$ denotes quadratic terms in $u$,
$\mathcal{Q}^{l}{}_{ijk}=g^{\prime}_{jk}g^{\prime}_{im}u^{m}u^{l}-g^{\prime}_{ik}g^{\prime}_{jm}u^{m}u^{l}$ (39)
The effective Ricci tensor is obtained by contraction,
$R^{u}_{jk}=(R^{u})^{i}{}_{jik}$ (40)
Explicitly,
$R^{u}_{jk}$ $=R_{jk}+\nabla^{i}g^{\prime}_{jk}\,u_{i}-\nabla^{i}g^{\prime}_{ik}\,u_{j}+g^{\prime}_{jk}\nabla^{i}u_{i}-g^{\prime}_{ik}\nabla^{i}u_{j}$
$\quad+g^{\prime}_{jk}g^{\prime}_{im}u^{i}u^{m}-g^{\prime}_{ik}g^{\prime}_{jm}u^{i}u^{m}$ (41)
Under the incompressibility condition $\nabla^{i}u_{i}=0$ and at leading order in the metric deformation, this reduces to
$R^{u}_{jk}\simeq R_{jk}+\alpha\,\nabla^{i}\bigl(\nabla_{j}S_{ik}+\nabla_{k}S_{ij}-\nabla_{i}S_{jk}\bigr).$ (42)
making explicit the coupling between curvature and deformation. In terms of the effective Ricci tensor $R^{u}_{ik}$, the metric evolution over time is
$\partial_{t}g^{\prime}_{i}k=\partial_{t}g_{ik}+\alpha S_{ik}$ (43)
may be formally rewritten as a Ricci-type flow,
$\partial_{t}g^{\prime}_{ik}+2R^{u}_{ik}=\mathcal{E}_{ik}$ (44)
where $\mathcal{E}_{ik}$ represents higher-order and transport-induced correction terms controlled by the deformation tensor $S$. Then, by definition $\mathcal{E}_{ik}$ may be regarded as
$\mathcal{E}_{ik}=\partial_{t}g_{ik}+\alpha S_{ik}+2R^{u}_{ik}$ (45)
This comparison does not assert that the physical metric evolves by Ricci flow, but rather that the Navier–Stokes dynamics induces a Ricci-type geometric structure at the microlocal level, where viscous dissipation enforces curvature control and prevents the formation of singular geometric configurations compatible with blow-up [4]. Then, the correction tensor $\mathcal{E}_{ik}$ admits the canonical decomposition
$\mathcal{E}_{ik}=\mathcal{E}^{(I)}_{ik}+\mathcal{E}^{(II)}_{ik}+\mathcal{E}^{(III)}_{ik}+\mathcal{E}^{(IV)}_{ik}$ (46)
where
$\mathcal{E}^{(I)}_{ik}$ $=\partial_{t}g_{ik}+2R_{ik}$ (47)
$\mathcal{E}^{(II)}_{ik}$ $=\alpha S_{ik}$ (48)
$\mathcal{E}^{(III)}_{ik}$ $=2\alpha\nabla^{j}(\nabla_{i}S_{jk}+\nabla_{k}S_{ji}-\nabla_{j}S_{ik})$ (49)
$\mathcal{E}^{(IV)}_{ik}$ $=2\mathcal{Q}_{ik}(u)$ (50)
Therefore, the tensor $\mathcal{E}_{ik}$ measures the deviation of the effective metric evolution from a pure Ricci flow, encoding transport, deformation, and higher-order nonlinear effects induced by the fluid dynamics. The first term of $\mathcal{E}_{ik}$ goes for geometric background, the second one for instantaneous deformation, the third one for geometric diffusion and the fourth one for non-linear transport.
IV Geometric Evolution Operators and Microlocal Representation of Solutions
By homogeneity, the field $X_{u}(x,\xi)$ is tangent to $S^{*}M$ and therefore defines a well-defined dynamics on the cosphere bundle, where $\|\xi\|_{g^{\prime}}$ is preserved. This will be useful because, at this point, it is beneficial to perform a Fourier transform on the system [5, 6]. Since the classical Fourier transform is not globally defined on compact manifolds, we employ a microlocal transform defined locally in normal coordinates and extended globally as a function on $T^{*}M$. For a 1-form $\phi\in\Omega^{1}(M)$, we define
$\tilde{\phi}(x,\xi)=\int_{T_{x}M}e^{-i\langle\xi,y\rangle}\phi(\exp_{x}y)\,\chi(y)\,dy$ (51)
where $\chi$ is a smooth cutoff function. Restricting by homogeneity to $S^{*}M$, we obtain $\tilde{\phi}\in C^{\infty}(S^{*}M)$.
Applying the microlocal transform to the incompressible covariant system and projecting onto transverse directions to eliminate the pressure term, we obtain the microlocal equation
$\partial_{t}\tilde{\phi}+\mathcal{L}_{X_{u}}\tilde{\phi}+\nu\|\xi\|_{g}^{2}\,\tilde{\phi}=\tilde{R}+\tilde{F},\qquad\langle\tilde{\phi},\xi\rangle_{g}=0$ (52)
Equation (52) is a linear non-autonomous dissipative system on $S^{*}M$. Here, the term corresponding to the pressure $dP$ does not contribute to the system’s dynamics, acting as a Lagrange multiplier that imposes the incompressibility condition, existing only in the radial direction of the cotangent bundle. Therefore, we define the incompressible microlocal projector
$\Pi_{\xi}:T^{\star}_{x}M\longmapsto\xi^{\perp}\subset T^{\star}_{x}M$ (53)
given by
$\Pi_{\xi}(\eta)=\eta-\frac{\langle\eta,\xi\rangle_{g}}{\|\xi\|_{g}^{2}}\xi$ (54)
which yields the conditions
$\Pi_{\xi}(\xi)=0,\quad\Pi_{\xi}(dP)=0$ (55)
This causes the pressure term to vanish when restricting the equation to the cosphere bundle and projecting onto transverse directions. In fact, the condition of the divergence transform also justifies the defined projection and confirms that the dynamics evolve naturally in $S^{*}M$.
In order to solve the microlocal system on the cosphere bundle $S^{*}M$, we define the Green operator $G(t,s)$ as the fundamental solution of the non-autonomous dissipative equation [25, 38]. Let $Q(t)=\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2}$ be the differential operator acting on sections of the bundle. The Green operator satisfies
$\left(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2}\right)G(t,s)=\delta(t-s)\mathbb{I}$ (56)
where $\mathbb{I}$ is the identity operator on $C^{\infty}(S^{*}M)$.
Let $\Psi_{t,s}:S^{*}M\longmapsto S^{*}M$ be the flow generated by the dynamical vector field $X_{u}$. This flow represents the lifting of the fluid’s trajectories to the phase space. For any initial microlocal distribution $\tilde{\phi}_{0}$, the pure transport part without dissipation is given by the pull-back $(\Psi_{t,s}^{*}\tilde{\phi}_{0})(x,\xi)=\tilde{\phi}_{0}(\Psi_{s,t}(x,\xi))$.
In the general case, let $\tilde{f}\in C^{\infty}(S^{*}M\times N)$ represent the total microlocal source term, encompassing both external forces and non-linear residuals with N{N} is a general manifold representing additional degrees of freedom, parameters or internal symmetries of the system. In the relativistic limit, $N$ is identified with the timelike component of the worldline, and the physical fields are recovered by the push-forward integrating over the fibers. Then, the Green operator is constructed by coupling the geometric transport with the exponential dissipation factor induced by the viscosity and the metric.
$G(t,s)[\tilde{f}](x,\xi)=\exp\left(-\int_{s}^{t}\nu\|\xi(\tau)\|_{g}^{2}\,d\tau\right)\tilde{f}(\Psi_{s,t}(x,\xi))$ (57)
where
  • $\xi(\tau)$ denotes the evolution of the covector along the integral curves of $X_{u}$ in the cotangent bundle.
  • $\tilde{f}$ is understood as a section whose values in N{N} are preserved along the flow $\Psi_{s,t}$ unless acted upon by internal dynamics, being part of the phase space for the system, describing the Fourier transform of the external forces and non-linear residuals of curvature.
To verify that $G(t,s)$ is the fundamental solution, we observe, at $t=s$, the flow $\Psi_{s,s}$ is the identity and the integral in the exponent vanishes, thus $G(s,s)=\mathbb{I}$. Applying the operator $\partial_{t}+\mathcal{L}_{X_{u}}$ to the transported term, we have by the property of the Lie derivative and the flow
$(\partial_{t}+\mathcal{L}_{X_{u}})\left[\tilde{f}(\Psi_{s,t}(x,\xi))\right]=0$ (58)
which follows from the fact that $\Psi_{s,t}$ is the solution to the transport equation associated with $X_{u}$. Differentiating the exponential factor with respect to $t$ yields the term $\nu\|\xi\|_{g}^{2}$. Summing these contributions, we obtain
$(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2})G(t,s)\\ =(-\nu\|\xi\|_{g}^{2}+\nu\|\xi\|_{g}^{2}+\mathcal{L}_{X_{u}}-\mathcal{L}_{X_{u}})G(t,s)=0$ (59)
for $t>s$ and it converges for the delta distribution at $t=s$. The operator thus acts as a low-pass filter in the phase space, where the dissipation follows the geometric trajectories.
Thus, the solution to the transformed microlocal system can be formally written by solving for $\tilde{\phi}$ over the time variable as
$\tilde{\phi}(t)=\mathcal{T}\bigg[\exp\left(-\int_{t_{0}}^{t}\big(\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2}\big)\,ds\right)\tilde{\phi}_{0}\bigg]\\ +\int_{t_{0}}^{t}G(t,s)\big[\tilde{R}(s)+\tilde{F}(s)\big]\,ds$ (60)
where
  • $\mathcal{T}$ is the time-ordering operator.
  • $G(t,s)$ is the associated Green’s operator.
  • $\tilde{\phi}_{0}$ is the initial microlocal data.
Finally, the physical solution is recovered by integrating over the cotangent directions
$\phi(x,t)=\int_{S_{x}^{\star}M}\tilde{\phi}(x,t,\xi)\,d\xi$ (61)
which defines a geometrically canonical reconstruction of the velocity field.
This formulation allows for the study of the system’s regularity through the microlocal behavior of $\tilde{\phi}$, reducing the global analysis of Navier-Stokes to the dynamical properties of the flow induced on $S^{*}M$. Analogously, we follow the same procedure for vorticity, obtaining the following equation
$\partial_{t}\tilde{\omega}+\mathcal{L}_{X_{u}}\tilde{\omega}=-\nu\|\xi\|_{g}^{2}\tilde{\omega}+\tilde{R}_{\omega}+i\xi\wedge\tilde{F},\qquad\xi\wedge\tilde{\omega}=0$ (62)
The condition $\xi\wedge\tilde{\omega}=0$ is the microlocal version of $d\omega=0$. Equation (62) is linear and dissipative, and its solution is formally written as
$\tilde{\omega}(t)=\mathcal{T}\bigg[\exp\left[-\int_{t_{0}}^{t}(\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2})\,ds\right]\tilde{\omega}_{0}\bigg]\\ +\int_{t_{0}}^{t}G(t,s)\big[\tilde{R}\omega(s)+i\xi\wedge\tilde{F}(s)\big]ds$ (63)
If $\tilde{\phi}(t)$ is a solution to (60), then $\tilde{\omega}(t):=d\tilde{\phi}(t)$ satisfies the microlocal vorticity equation (62). This is because the exterior derivative commutes with the differential operators used in the system for smooth differential forms, and $d(dP)=0$; therefore, applying the exterior derivative to (52) yields exactly (62) with a solution reconstructed also from
$\omega(x,t)=\int_{S_{x}^{*}M}\tilde{\omega}(x,t,\xi)\,d\xi$ (64)
It is inferred that the microlocal lift of the Navier-Stokes equations produces a well-defined dissipative dynamical system on the cosphere bundle, whose solution admits an explicit representation through evolution operators. The exact compatibility between the velocity and vorticity equations guaranties the consistency of the formalism. This approach reduces the global regularity problem to the control and convergence of geometric microlocal invariants, the analysis of which is addressed in the following section.
V Stability, Energy and Geometric Criteria for Regularity and Smoothness
In the previous sections, the incompressible Navier–Stokes equations were reformulated as a linear, non-autonomous, dissipative dynamical system acting on microlocal lifts of differential forms over the compact phase space $S^{*}M$. We now exploit this structure to derive quantitative stability estimates and geometric criteria for regularity [28]. The central idea of this section is that any potential singular behavior of the physical solution must manifest itself as a geometric instability of the microlocal dynamics on $S^{*}M$ [29, 7]. Viscosity, compactness, and the geometric nature of the lift combine to impose strong restrictions on such instabilities.
Throughout this section, $\lambda\in\{\phi,\omega\}$ denotes either the velocity $1$-form or the vorticity $2$-form on $M$, and $\tilde{\lambda}$ its microlocal lift to $S^{*}M$. All norms and contractions are taken with respect to the effective metric $g^{\prime}$ introduced earlier. Since $S^{*}M$ is compact and endowed with a canonical contact structure, all integrals below are well-defined without further assumptions.
We begin by defining the fundamental local geometric quantity governing the analysis. The microlocal amplitude of $\tilde{\lambda}$ is defined pointwise by the contraction
$|\tilde{\lambda}|_{g^{\prime}}^{2}:=\langle\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}=\iota_{\tilde{\lambda}^{\sharp}}\tilde{\lambda}$, (65)
where $\tilde{\lambda}^{\sharp}$ denotes the metric dual with respect to $g^{\prime}$. This quantity measures the intensity of the lifted differential form along a given cotangent direction. Unlike classical norms, it retains full directional information encoded in the cosphere bundle.
The corresponding global geometric functional is obtained by pairing this density with the canonical Liouville volume form on $S^{*}M$,
$\mu_{S^{*}M}=\alpha\wedge(d\alpha)^{n-1}$ (66)
where $\alpha$ is the contact form induced from the Liouville $1$-form on $T^{*}M$. We define the microlocal energy functional
$\mathcal{V}_{\lambda}(t):=\int_{S^{*}M}|\tilde{\lambda}(x,\xi,t)|_{g^{\prime}}\,\mu_{S^{*}M}$ (67)
Geometrically, $\mathcal{V}_{\lambda}$ measures the total directional spread of the field $\lambda$ in phase space. Physically, it quantifies how much the solution concentrates along preferred cotangent directions, independently of its spatial localization.
We now derive a balance law for $\mathcal{V}_{\lambda}$. The microlocal evolution equation satisfied by $\tilde{\lambda}$ is
$\partial_{t}\tilde{\lambda}+\mathcal{L}_{X_{u}}\tilde{\lambda}+\nu\|\xi\|_{g}^{2}\,\tilde{\lambda}=\tilde{R}[\tilde{\lambda}]+\tilde{C}[\tilde{\lambda}]+\tilde{F}$ (68)
where $X_{u}$ is the canonical lifted transport field on $S^{*}M$. We take the pointwise inner product with $\tilde{\lambda}$ and use bilinearity to obtain
$\frac{1}{2}\partial_{t}|\tilde{\lambda}|_{g^{\prime}}^{2}+\frac{1}{2}\mathcal{L}_{X_{u}}|\tilde{\lambda}|_{g^{\prime}}^{2}+\nu\|\xi\|_{g}^{2}|\tilde{\lambda}|_{g^{\prime}}^{2}\\ =\langle\tilde{R}[\tilde{\lambda}],\tilde{\lambda}\rangle+\langle\tilde{C}[\tilde{\lambda}],\tilde{\lambda}\rangle+\langle\tilde{F},\tilde{\lambda}\rangle$ (69)
Each term on the right-hand side admits a sharp geometric bound. Since the curvature term is zeroth order in $\xi$,
$|\langle\tilde{R}[\tilde{\lambda}],\tilde{\lambda}\rangle|\leq\|\mathrm{Ric}\|_{L^{\infty}(M)}\,\|\tilde{\lambda}\|_{g^{\prime}}^{2}$ (70)
The commutator term, arising from non-exact microlocalization, satisfies
$|\langle\tilde{C}[\tilde{\lambda}],\tilde{\lambda}\rangle|\leq C\,\|\nabla u\|_{L^{\infty}(M)}\,\|\tilde{\lambda}\|_{g^{\prime}}^{2}$ (71)
Finally, the forcing term is controlled linearly,
$|\langle\tilde{F},\tilde{\lambda}\rangle|\leq\|\tilde{F}\|\,\|\tilde{\lambda}\|_{g^{\prime}}$ (72)
We now integrate the balance identity over $S^{*}M$. Since the Liouville measure is invariant under the flow generated by $X_{u}$, that is,
$\mathcal{L}_{X_{u}}\mu_{S^{*}M}=0$ (73)
the transport term drops out after integration. We obtain
$\frac{d}{dt}\mathcal{V}_{\lambda}(t)+\nu\int_{S^{*}M}\|\xi\|_{g}^{2}\|\tilde{\lambda}\|_{g^{\prime}}^{2}\,\mu_{S^{*}M}\leq C\|\nabla u\|_{L^{\infty}}\mathcal{V}_{\lambda}(t)\\ +\|\tilde{F}\|_{L^{1}(S^{*}M)}$ (74)
This inequality expresses a geometric fact, microlocal energy cannot accumulate without being dissipated, unless the deformation of the flow becomes uncontrolled. In particular, in the absence of forcing, $\mathcal{V}_{\lambda}$ is monotone up to a deformation-controlled factor.
To make this mechanism more explicit, recall that the velocity gradient decomposes as $\nabla u=S+\Omega$, where only the symmetric part $S$ contributes to stretching. Restricting the evolution to integral curves of $X_{u}$ yields the pointwise inequality
$\frac{d}{dt}\|\tilde{\lambda}\|\leq\|S(t)\|_{L^{\infty}}\,\|\tilde{\lambda}\|-\nu\|\tilde{\lambda}\|$ (75)
with solution
$\|\tilde{\lambda}(t)\|\leq\|\tilde{\lambda}(0)\|\exp\!\left[\int_{t_{0}}^{t}(\|S(s)\|_{L^{\infty}}-\nu)\,ds\right]$ (76)
showing that any amplification of microlocal amplitude requires a persistent failure of temporal integrability of $\|S\|_{L^{\infty}}$.
We now specialize to the vorticity $\tilde{\omega}$, whose geometric structure allows for finer control. Since $\tilde{\omega}$ is a closed microlocal $2$-form, we introduce the functional
$W[\tilde{\omega}]=\int_{S^{*}M}\left(\tau\,|d^{\perp}\log|\tilde{\omega}||^{2}+\log|\tilde{\omega}|\right)\mu_{S^{*}M}$ (77)
where $d^{\perp}$ denotes the vertical exterior derivative induced by the horizontal–vertical splitting of $TS^{*}M$, and $\tau>0$ is a dimensional parameter. This functional penalizes sharp angular gradients of vorticity and quantifies directional irregularity.
A direct differentiation in time, combined with integration by parts along the fibers [30] and the compactness of $S^{*}M$, yields the dissipation inequality
$\frac{d}{dt}W[\tilde{\omega}]\leq\int_{S^{*}M}\left(\tau\,|d^{\perp}\log|\tilde{\omega}||^{2}-\tau\,\|S\|\,|\tilde{\omega}|-\nu\right)\mu_{S^{*}M}$ (78)
This inequality shows that viscosity and geometric diffusion suppress directional oscillations of vorticity, unless sustained stretching dominates dissipation.
Since $S^{*}M$ is compact, the above estimates imply uniform control of $\tilde{\omega}$ in $L^{\infty}(S^{*}M)$, and therefore of the reconstructed physical vorticity [4]. Consequently, any loss of regularity must be accompanied by a breakdown of one of the assumptions used above.
We arrive at a geometric classification of admissible blow-up mechanisms [8]. A singularity can occur only if at least one of the following phenomena takes place,
  • Extreme directional concentration of microlocal vorticity in $S^{*}M$.
  • Persistent loss of vortex alignment encoded in the symmetric deformation tensor.
  • Failure of temporal integrability of $\|\nabla u\|_{L^{\infty}}$.
Viscous dissipation excludes all other scenarios by enforcing monotonicity of microlocal geometric invariants.
In conclusion, the Navier-Stokes equations admit a rigid geometric reformulation in which regularity is equivalent to the stability of a dissipative dynamical system on a compact phase space. Blow-up is not a purely analytic accident, but a geometric instability that must overcome strong dissipative and structural constraints.
VI Discussion and Conclusions
The theoretical approach developed in this work addresses the Navier–Stokes regularity problem through a synthesis of microlocal analysis and Riemannian geometry. The central structural step is the lifting of the dissipative dynamics from the physical manifold $M$ to the cosphere bundle $S^{*}M$, which resolves several ambiguities inherent in formulations based solely on physical space. In contrast with classical regularity theories focused on the control of Sobolev norms and vorticity amplitudes [11, 40], the present framework emphasizes directional alignment and geometric dissipation as primary mechanisms governing regularity. This perspective is consistent with geometric trapping phenomena [4] and extends them to three dimensions through the use of the Hodge Laplacian and its variational structure [7].
The geometric role of curvature, revealed through the Weitzenböck identity and the coupling of the Ricci tensor to vorticity dynamics, shows that the topology of $M$ actively contributes to dissipation rather than serving as a passive background [3, 7]. The monotonicity of the vortex energy functional $\mathcal{E}_{\omega}$ and the associated microlocal inequalities restrict admissible blow–up scenarios in a manner compatible with known geometric constraints [8, 31]. In this setting, the vortex stretching term is reinterpreted not as an intrinsically destabilizing mechanism, but as a geometric deformation acting on a compact fiber, where viscous effects induce effective angular diffusion. The microlocal transport–dissipation structure is consistent with the analysis of lifted operators and unique continuation phenomena developed in [29], and aligns with energy–aware geometric decompositions of the Navier–Stokes system [6].
For that, this paper does not solve the Navier–Stokes regularity problem. Instead, it introduces a coherent and geometrically meaningful framework that recasts the question in microlocal and dynamical terms, where regularity is linked to directional alignment and dissipative stability on a compact phase space [29]. In particular, the formulation is compatible with recent results on non–uniqueness of smooth or weak solutions arising from critical data [10], this can be interpreted as manifestations of microlocal flexibility within the constraints imposed by energy dissipation and geometric monotonicity. The present framework does not preclude non–uniqueness at low regularity; rather, it provides a mechanism to distinguish dynamically stable microlocal configurations from those that are highly sensitive to perturbations and thus prone to branching or loss of uniqueness.
From this viewpoint, regularity is not characterized solely by uniqueness or smoothness in physical space, but by the persistence of controlled microlocal structure along the lifted flow. Singular or non–unique behaviors must then correspond to a breakdown of directional control or of dissipative balance in $S^{*}M$, rather than to uncontrolled growth of classical norms alone. This reformulation clarifies how geometric dissipation and microlocal transport can coexist with modern flexibility results, while still yielding meaningful regularity criteria for physically relevant solutions.
A complementary geometric interpretation is obtained by considering the lifted dynamics on the extended manifold $S^{*}M\times N$, where $N\subset\mathbb{R}$ denotes the temporal interval. In this formulation, the temporal boundary components of $N$ act as Cauchy hypersurfaces for the evaluation of microlocal energies and functionals [39]. By Stokes’ theorem, variations of the geometric invariants introduced in this work reduce to boundary fluxes [30], showing that changes in microlocal energy and directional structure are governed by boundary contributions rather than interior concentration mechanisms. Preservation of smoothness can therefore be interpreted as the absence of singular behavior in the asymptotic microlocal distribution, excluding concentration at infinity in phase space [31, 4].
Despite the structural coherence of the microlocal framework developed here, several important limitations and open directions remain. First, while the lifting to $S^{*}M$ provides refined control over directional alignment and dissipative mechanisms, it inevitably entails a loss of direct information when projecting back to physical space. Quantifying precisely which microlocal features are preserved under reconstruction, and which may be lost or blurred, remains an open problem. In particular, sharper criteria are needed to distinguish benign directional dispersion from genuinely unstable microlocal configurations.
Second, the present analysis is carried out on a fixed Riemannian manifold and does not address the possibility of topological or geometric transitions in the underlying space. Understanding how changes in topology, curvature, or bundle structure affect microlocal stability and dissipation represents a natural and largely unexplored direction. Such effects may be relevant in flows with time–dependent geometry or in singular limits where geometric control degenerates.
Finally, although the derived inequalities severely restrict admissible blow–up scenarios, they do not yet yield a complete classification of singularity formation mechanisms. Developing refined microlocal criteria that can discriminate between transient directional amplification and genuinely singular behavior, as well as identifying optimal strategies for detecting or preventing microlocal concentration, remain important challenges for future work. Addressing these questions may require combining the present geometric framework with finer second microlocalization techniques or with adaptive topological invariants capable of capturing changes in phase–space structure. Making this approach a geometric reorganization of the regularity problem rather than a final resolution [37]. Its structure may nevertheless open new avenues for understanding singularity formation and dissipative stability from a geometric perspective [8].
Beyond its theoretical implications, the microlocal framework suggests potential applications and development in structure–preserving numerical schemes and geometry–aware computational fluid dynamics, where directional instabilities are controlled through intrinsic geometric mechanisms rather than arbitrary regularization [1, 6, 15, 5]. These ideas may be useful in turbulence modeling, data-driven reduced-order models, and multi-scale simulations in cluster environments where know and preserve the appropriate geometric properties and its structure is essential for long-time accuracy.
VII AI Statement
The author confirms that every intellectual contribution, original mathematical and physical derivations, theorem, proofs, and physical hypotheses presented in this manuscript are the sole work of the author.
Generative Artificial Intelligence (AI) tools, specifically the Large Language Model ChatGPT, were used exclusively as digital research assistants for the following purposes: refining academic language and structuring the references into alphabetic order. The AI did not perform any original scientific calculations, nor proposed any scientific content as hypotheses, proofs or others. The author has reviewed, edited, and validated the accuracy of all output generated by the AI.
Acknowledgements.
The author is profoundly grateful to PhD. Pedro Díaz-Navarro because of his openess and creativity to talk. In the same way as Octavio Rojas-Quesada for every new point of view and Celicia Ledezma-Porras for her geometric insight some years ago. Moreover, family, some friends and every guide, person… who has stumble with me along the paths, thank you all.

### v2

@@
 We develop a geometric and microlocal framework for the Navier–Stokes equations by lifting the dynamics to the cosphere bundle of a Riemannian manifold.
-In this formulation, the velocity field and vorticity are represented as microlocal distributions whose evolution is governed by a transport–dissipation system generated by a canonical dynamical vector field described with differential forms.
+In this formulation, the velocity field and vorticity are represented as microlocal distributions whose evolution is governed by a linear transport–dissipation system generated by a canonical dynamical vector field.
 We introduce microlocal amplitudes, directional energy functionals, and monotone volume invariants on the compact phase space, which quantify directional concentration and alignment mechanisms associated with potential loss of regularity.
-The viscous term induces effective geometric diffusion on the cosphere bundle, yielding closed differential inequalities in a geometric setting.
+The viscous term induces an effective geometric diffusion on the cosphere bundle, yielding closed differential inequalities in a geometric setting.
 To capture the interaction between fluid deformation and geometry, we define an effective connection and curvature tensor encoding the influence of the symmetric velocity gradient.
-This structure gives rise to a Ricci-type geometric evolution at the microlocal level, controlled by a dissipative functional that excludes extreme directional concentration and persistent transverse stretching.
-Our results do not resolve the global regularity problem in full generality, but provide a coherent geometric mechanism that severely restricts blow-up scenarios compatible with viscous Navier–Stokes dynamics, reformulating the analytic regularity question as a problem of dissipative stability on a compact phase space.
+This structure gives rise to a Ricci–type microlocal geometric evolution that constrains directional stretching and excludes extreme angular concentration compatible with viscous dissipation.
+While the present results do not resolve the global regularity problem, they provide a coherent geometric mechanism that severely restricts admissible blow–up scenarios, reformulating the regularity question as a problem of dissipative stability on a compact phase space.
 Keywords
@@
 preprint: APS/123-QED
+Mathematics Subject Classification (2020): 35Q30; 53C21; 35S30; 58J65; 35A23.
+ACM Class: G.1.8; G.1.10; J.2
 I Introduction: The Regularity Problem and Geometric Flows
@@
 In this way, we find inequalities and invariants to show how the physical solution is reconstructed from the dynamics in $S^{*}M$.
+Microlocal lifting approach
+The microlocal lifting to the cosphere bundle $S^{*}M$ adopted in this paper should be understood as a parametric refinement and compactification of the dynamics on the base manifold $M$.
+All geometric objects—including the effective connection $\nabla^{u}$, its torsion and curvature, and the metric $g^{\prime}$—are defined on $M$, with coefficients depending smoothly on $(x,\xi)\in S^{*}M$.
+The cotangent variable $\xi$ serves to encode directional and spectral information, but no geometric structure is imposed on the tangent bundle of $S^{*}M$ itself.
+This viewpoint allows us to combine microlocal analysis with intrinsic geometric evolution without introducing additional bundle-level connections.
 II Covariant Formulation on Manifolds
@@
 $g^{\prime}_{ik}=g_{ik}+\alpha\,\int_{t_{0}}^{t}S_{ik}(s)\ ds,\qquad S_{ik}=\frac{1}{2}\bigl(\nabla_{i}u_{k}+\nabla_{k}u_{i}\bigr)$ (35)
-where $\alpha$ is a fixed geometric coupling parameter [21]. This metric deformation encodes the interaction between the background geometry and the fluid deformation tensor.
+where $\alpha$ is a fixed geometric coupling parameter [21].
+This metric deformation encodes the interaction between the background geometry and the fluid deformation tensor.
+We define the effective covariant derivative $\nabla^{u}$ acting on vector fields by
+$(\nabla^{u})_{i}v^{k}:=\nabla_{i}v^{k}+A^{k}{}_{ij}v^{j}$ (36)
+where $\nabla$ denotes the Levi–Civita connection of $g$ and $A^{k}{}_{ij}$ is a (1,2)(1,2)–tensor encoding the deformation induced by $u$.
+The effective connection $\nabla^{u}$ is, in general, not metric–compatible with respect to the effective metric $g^{\prime}$. Metric compatibility would require
+$(\nabla^{u}_{i}g^{\prime})_{jk}=0$ (37)
+where, by definition,
+$(\nabla^{u}_{i}g^{\prime})_{jk}:=\partial_{i}g^{\prime}_{jk}-(\Gamma^{u})^{\ell}{}_{ij}\,g^{\prime}_{\ell k}-(\Gamma^{u})^{\ell}{}_{ik}\,g^{\prime}_{j\ell}$ (38)
+Using the decomposition $(\Gamma^{u})^{\ell}{}_{ij}=\Gamma^{\ell}{}_{ij}+A^{\ell}{}_{ij}$ and recalling that $\nabla$ denotes any connection, for example, using the Levi–Civita connection associated with the background metric $g$, we obtain
+$(\nabla^{u}_{i}g^{\prime})_{jk}$ $=(\nabla_{i}g^{\prime})_{jk}-A^{\ell}{}_{ij}\,g^{\prime}_{\ell k}-A^{\ell}{}_{ik}\,g^{\prime}_{j\ell}$ (39)
+In particular, if $g^{\prime}$ is covariantly constant with respect to $\nabla$, that is $\nabla g^{\prime}=0$, the expression reduces to
+$(\nabla^{u}_{i}g^{\prime})_{jk}=-A^{\ell}{}_{ij}\,g^{\prime}_{\ell k}-A^{\ell}{}_{ik}\,g^{\prime}_{j\ell}$ (40)
+Equivalently, in invariant tensorial form,
+$\nabla^{u}g^{\prime}=-\bigl(A(u)\cdot g^{\prime}+g^{\prime}\cdot A(u)\bigr)$ (41)
+where $A(u)$ is viewed as a (1,1)(1,1)–tensor acting on covariant indices.
+Therefore, the effective connection $\nabla^{u}$ is metric–compatible with respect to $g^{\prime}$ if and only if the deformation tensor $A(u)$ is skew–adjoint with respect to $g^{\prime}$, namely if
+$g^{\prime}[A(u)X,Y]+g^{\prime}[X,A(u)Y]=0$ (42)
+Here $X,Y$ denote smooth vector fields on the base manifold $M$, such that, $X,Y\in\Gamma(TM)$.
+Although the tensors $A(u)$ and $g^{\prime}$ may depend microlocally on $(x,\xi)\in S^{*}M$, the effective connection $\nabla^{u}$ acts on tensor fields over $M$, with the cotangent variables playing the role of parameters.
+In general, this condition is violated by the symmetric part of $\nabla u$, and $\nabla^{u}$ defines a non–metric affine connection encoding the velocity–induced deformation of the effective geometry encoding the transport induced by the velocity field controlled by $\|\nabla u\|_{L^{\infty}}$, which is sufficient for the energy estimates derived below.
+Then, the canonical decomposition considered earlier remains orthogonal with respect to the background metric $g$ and orthogonal with respect to the effective metric $g^{\prime}$ up to lower-order corrections.
+Since the Levi–Civita connection is torsion–free, the torsion tensor $T^{u}$ of $\nabla^{u}$ is entirely determined by the antisymmetric part of $A$. Indeed, by definition,
+$(T^{u})^{k}{}_{ij}$ $:=(\Gamma^{u})^{k}{}_{ij}-(\Gamma^{u})^{k}{}_{ji}$ (43)
+[…]
+[…]
+Hence,
+$(T^{u})^{k}{}_{ij}=A^{k}{}_{ij}-A^{k}{}_{ji}$ (46)
+We now specify the structure of the deformation tensor $A$. We decompose it as
+$A^{k}{}_{ij}=g^{\prime k\ell}g^{\prime}_{im}\nabla_{j}u^{m}+B^{k}{}_{ij}$ (47)
+where $g^{\prime}$ is the effective metric introduced previously and $B^{k}{}_{ij}$ is symmetric in the lower indices,
+$B^{k}{}_{ij}=B^{k}{}_{ji}$ (48)
+Substituting this decomposition into the expression for the torsion, the symmetric part cancels and we obtain
+[…]
+$=g^{\prime k\ell}\left(g^{\prime}_{im}\nabla_{j}u^{m}-g^{\prime}_{jm}\nabla_{i}u^{m}\right)$ (50)
+This computation shows that
+  • the torsion $T^{u}$ is antisymmetric in the indices $i,j$.
+  • $T^{u}$ is of first order in $\nabla u$.
+  • $T^{u}\equiv 0$ if and only if the velocity field $u$ is covariantly constant.
+In particular, the effective connection $\nabla^{u}$ is torsion–free if and only if the velocity field $u$ is covariantly constant.
 Under this construction, the curvature tensor associated with the effective connection $\Gamma^{u}$ is given in local coordinates by
@@
 $\quad+(\Gamma^{m}_{jk}+g^{\prime}_{jk}u^{m})(\Gamma^{l}_{im}+g^{\prime}_{im}u^{l})$
-$\quad-(\Gamma^{m}_{ik}+g^{\prime}_{ik}u^{m})(\Gamma^{l}_{jm}+g^{\prime}_{jm}u^{l}).$ (36)
+$\quad-(\Gamma^{m}_{ik}+g^{\prime}_{ik}u^{m})(\Gamma^{l}_{jm}+g^{\prime}_{jm}u^{l})$ (51)
 This decomposes as
-$(R^{u})^{l}{}_{ijk}=R^{l}{}_{ijk}+\mathcal{A}^{l}{}_{ijk}+\mathcal{Q}^{l}{}_{ijk}$, (37)
+$(R^{u})^{l}{}_{ijk}=R^{l}{}_{ijk}+\mathcal{A}^{l}{}_{ijk}+\mathcal{Q}^{l}{}_{ijk}$ (52)
 where $R^{l}{}_{ijk}$ is the Riemann curvature of $g$
@@
 $\mathcal{Q}^{l}{}_{ijk}=g^{\prime}_{jk}g^{\prime}_{im}u^{m}u^{l}-g^{\prime}_{ik}g^{\prime}_{jm}u^{m}u^{l}$ (54)
-The effective Ricci tensor is obtained by contraction,
+The curvature tensor $(R^{u})^{l}{}_{ijk}$ associated with the effective connection $\nabla^{u}$ satisfies a modified second Bianchi identity, reflecting the presence of torsion-like terms induced by the velocity-dependent deformation of the connection.
+More precisely, since the effective connection $\nabla^{u}$ is an affine connection with torsion, its curvature tensor satisfies the standard second Bianchi identity for connections with torsion.
+In index notation, one has
+$\nabla^{u}_{[m}(R^{u})^{l}{}_{|i|jk]}=(T^{u})^{p}{}_{[mj}(R^{u})^{l}{}_{|i|k]p}$ (55)
+where $(T^{u})^{p}{}_{ij}$ denotes the torsion tensor of $\nabla^{u}$, defined earlier.
+Here, brackets $\cdot$] denote antisymmetrization over the indicated indices, while vertical bars $|\cdot|$ exclude the enclosed index from the antisymmetrization.
+This identity follows directly from the definition of the curvature operator $\forall\ X,Y,Z\in\Gamma(TM)$, satisfying,
+$R^{u}(X,Y)Z=\nabla^{u}_{X}\nabla^{u}_{Y}Z-\nabla^{u}_{Y}\nabla^{u}_{X}Z-\nabla^{u}_{[X,Y]}Z$ (56)
+together with the Jacobi identity for iterated covariant derivatives [30].
+The additional torsion term on the right-hand side of (55) measures precisely the failure of $\nabla^{u}$ to be torsion-free, and vanishes identically when $T^{u}\equiv 0$.
+Since the deformation tensor $A$ depends smoothly on the velocity field $u$ and its first derivatives, the torsion tensor $(T^{u})^{p}{}_{ij}$ is of first order in $\nabla u$.
+Consequently, the right-hand side of (55) consists of lower-order terms which are uniformly controlled under the regularity assumptions imposed on $u$.
+In particular, the classical second Bianchi identity is recovered in the limit where $A(u)$ vanishes.
+In the general case, the deviation from the classical identity is entirely encoded in controlled lower-order terms involving $\nabla u$.
+Now, the effective Ricci tensor is obtained by contraction,
 $R^{u}_{jk}=(R^{u})^{i}{}_{jik}$ (57)
@@
 Under the incompressibility condition $\nabla^{i}u_{i}=0$ and at leading order in the metric deformation, this reduces to
-$R^{u}_{jk}\simeq R_{jk}+\alpha\,\nabla^{i}\bigl(\nabla_{j}S_{ik}+\nabla_{k}S_{ij}-\nabla_{i}S_{jk}\bigr).$ (42)
+$R^{u}_{jk}\simeq R_{jk}+\alpha\,\nabla^{i}\bigl(\nabla_{j}S_{ik}+\nabla_{k}S_{ij}-\nabla_{i}S_{jk}\bigr)$ (59)
 making explicit the coupling between curvature and deformation. In terms of the effective Ricci tensor $R^{u}_{ik}$, the metric evolution over time is
@@
 $\tilde{\phi}(x,\xi)=\int_{T_{x}M}e^{-i\langle\xi,y\rangle}\phi(\exp_{x}y)\,\chi(y)\,dy$ (68)
-where $\chi$ is a smooth cutoff function. Restricting by homogeneity to $S^{*}M$, we obtain $\tilde{\phi}\in C^{\infty}(S^{*}M)$.
+where $\chi$ is a smooth cutoff function.
+Restricting by homogeneity to $S^{*}M$, we obtain $\tilde{\phi}\in C^{\infty}(S^{*}M)$.
+The transform (68) should be understood as a directional microlocal projection rather than a pointwise invertible transform, because of that, the horizontal lift is uniquely determined once the effective connection is fixed, although it is not canonical with respect to the background metric.
 Applying the microlocal transform to the incompressible covariant system and projecting onto transverse directions to eliminate the pressure term, we obtain the microlocal equation
-$\partial_{t}\tilde{\phi}+\mathcal{L}_{X_{u}}\tilde{\phi}+\nu\|\xi\|_{g}^{2}\,\tilde{\phi}=\tilde{R}+\tilde{F},\qquad\langle\tilde{\phi},\xi\rangle_{g}=0$ (52)
+$\partial_{t}\tilde{\phi}+\mathcal{L}_{X_{u}}\tilde{\phi}+\nu\|\xi\|_{g^{\prime}}^{2}\,\tilde{\phi}=\tilde{R}+\tilde{F},\qquad\langle\tilde{\phi},\xi\rangle_{g^{\prime}}=0$ (69)
 Equation (69) is a linear non-autonomous dissipative system on $S^{*}M$.
@@
 given by
-$\Pi_{\xi}(\eta)=\eta-\frac{\langle\eta,\xi\rangle_{g}}{\|\xi\|_{g}^{2}}\xi$ (54)
+$\Pi_{\xi}(\eta)=\eta-\frac{\langle\eta,\xi\rangle_{g^{\prime}}}{\|\xi\|_{g^{\prime}}^{2}}\xi$ (71)
 which yields the conditions
 $\Pi_{\xi}(\xi)=0,\quad\Pi_{\xi}(dP)=0$ (72)
-This causes the pressure term to vanish when restricting the equation to the cosphere bundle and projecting onto transverse directions.
+This causes the pressure term to vanish when restricting the equation to the cosphere bundle and projecting onto transverse directions because it acts on the base manifold variables and does not generate transverse components in phase space.
 In fact, the condition of the divergence transform also justifies the defined projection and confirms that the dynamics evolve naturally in $S^{*}M$.
 In order to solve the microlocal system on the cosphere bundle $S^{*}M$, we define the Green operator $G(t,s)$ as the fundamental solution of the non-autonomous dissipative equation [25, 38].
-Let $Q(t)=\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2}$ be the differential operator acting on sections of the bundle.
+Let $Q(t)=\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2}$ be the differential operator acting on sections of the bundle.
 The Green operator satisfies
-$\left(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2}\right)G(t,s)=\delta(t-s)\mathbb{I}$ (56)
+$\left(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2}\right)G(t,s)=\delta(t-s)\mathbb{I}$ (73)
 where $\mathbb{I}$ is the identity operator on $C^{\infty}(S^{*}M)$.
@@
 Then, the Green operator is constructed by coupling the geometric transport with the exponential dissipation factor induced by the viscosity and the metric.
-$G(t,s)[\tilde{f}](x,\xi)=\exp\left(-\int_{s}^{t}\nu\|\xi(\tau)\|_{g}^{2}\,d\tau\right)\tilde{f}(\Psi_{s,t}(x,\xi))$ (57)
+$G(t,s)[\tilde{f}](x,\xi)=\exp\left(-\int_{s}^{t}\nu\|\xi(\tau)\|_{g^{\prime}}^{2}\,d\tau\right)\tilde{f}(\Psi_{s,t}(x,\xi))$ (74)
 where
@@
 Summing these contributions, we obtain
-$(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2})G(t,s)\\ =(-\nu\|\xi\|_{g}^{2}+\nu\|\xi\|_{g}^{2}+\mathcal{L}_{X_{u}}-\mathcal{L}_{X_{u}})G(t,s)=0$ (59)
+$(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2})G(t,s)\\ =(-\nu\|\xi\|_{g^{\prime}}^{2}+\nu\|\xi\|_{g^{\prime}}^{2}+\mathcal{L}_{X_{u}}-\mathcal{L}_{X_{u}})G(t,s)=0$ (76)
 for $t>s$ and it converges for the delta distribution at $t=s$. The operator thus acts as a low-pass filter in the phase space, where the dissipation follows the geometric trajectories.
 Thus, the solution to the transformed microlocal system can be formally written by solving for $\tilde{\phi}$ over the time variable as
-$\tilde{\phi}(t)=\mathcal{T}\bigg[\exp\left(-\int_{t_{0}}^{t}\big(\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2}\big)\,ds\right)\tilde{\phi}_{0}\bigg]\\ +\int_{t_{0}}^{t}G(t,s)\big[\tilde{R}(s)+\tilde{F}(s)\big]\,ds$ (60)
+$\tilde{\phi}(t)=\mathcal{T}\bigg[\exp\left(-\int_{t_{0}}^{t}\big(\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2}\big)\,ds\right)\tilde{\phi}_{0}\bigg]\\ +\int_{t_{0}}^{t}G(t,s)\big[\tilde{R}(s)+\tilde{F}(s)\big]\,ds$ (77)
 where
@@
 which defines a geometrically canonical reconstruction of the velocity field.
+The reconstruction equation (78) recovers $\phi$ in a weak or averaged sense, which is sufficient for the purposes of the present analysis.
 This formulation allows for the study of the system’s regularity through the microlocal behavior of $\tilde{\phi}$, reducing the global analysis of Navier-Stokes to the dynamical properties of the flow induced on $S^{*}M$.
 Analogously, we follow the same procedure for vorticity, obtaining the following equation
-$\partial_{t}\tilde{\omega}+\mathcal{L}_{X_{u}}\tilde{\omega}=-\nu\|\xi\|_{g}^{2}\tilde{\omega}+\tilde{R}_{\omega}+i\xi\wedge\tilde{F},\qquad\xi\wedge\tilde{\omega}=0$ (62)
+$\partial_{t}\tilde{\omega}+\mathcal{L}_{X_{u}}\tilde{\omega}=-\nu\|\xi\|_{g^{\prime}}^{2}\tilde{\omega}+\tilde{R}_{\omega}+i\xi\wedge\tilde{F},\qquad\xi\wedge\tilde{\omega}=0$ (79)
 The condition $\xi\wedge\tilde{\omega}=0$ is the microlocal version of $d\omega=0$. Equation (79) is linear and dissipative, and its solution is formally written as
-$\tilde{\omega}(t)=\mathcal{T}\bigg[\exp\left[-\int_{t_{0}}^{t}(\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2})\,ds\right]\tilde{\omega}_{0}\bigg]\\ +\int_{t_{0}}^{t}G(t,s)\big[\tilde{R}\omega(s)+i\xi\wedge\tilde{F}(s)\big]ds$ (63)
+$\tilde{\omega}(t)=\mathcal{T}\bigg[\exp\left[-\int_{t_{0}}^{t}(\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2})\,ds\right]\tilde{\omega}_{0}\bigg]\\ +\int_{t_{0}}^{t}G(t,s)\big[\tilde{R}\omega(s)+i\xi\wedge\tilde{F}(s)\big]ds$ (80)
 If $\tilde{\phi}(t)$ is a solution to (77), then $\tilde{\omega}(t):=d\tilde{\phi}(t)$ satisfies the microlocal vorticity equation (79).
-This is because the exterior derivative commutes with the differential operators used in the system for smooth differential forms, and $d(dP)=0$; therefore, applying the exterior derivative to (52) yields exactly (62) with a solution reconstructed also from
+This is because $d(dP)=0$ and $X_{u}$ depends on $u$ and its derivatives, the exterior derivative does not commute exactly with the Lie derivative along $X_{u}$, a direct computation yields the commutator identity
+$d,\mathcal{L}_{X_{u}}]=\mathcal{L}_{dX_{u}}$ (81)
+When acting on microlocal scalar amplitudes $\tilde{\phi}$, the right–hand side produces lower–order terms depending on $\nabla u$, and we obtain the estimate
+$d,\mathcal{L}_{X_{u}}]\tilde{\phi}=\mathcal{O}(\nabla u)\tilde{\phi}$ (82)
+The viscous contribution in the lifted equation is modeled by the multiplication operator
+$D:=\nu\|\xi\|_{g^{\prime}}^{2}$ (83)
+whose commutator with $d$ is given explicitly by
+$d,D]=d(\nu\|\xi\|_{g^{\prime}}^{2})\wedge(\cdot)$ (84)
+Combining the transport and dissipative contributions, we define the total commutator operator
+$\left[d,\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2}\right]={R}$ (85)
+which is a zeroth–order operator controlled by the geometry and the velocity field. More precisely, there exists a constant $C>0$ such that
+$\|{R}_{u}\|_{L^{2}\to L^{2}}\leq C\left(\|\nabla u\|_{L^{\infty}}+1\right)$ (86)
+As a consequence, if $\tilde{\phi}$ is a solution of the lifted scalar equation, then the microlocal vorticity $\tilde{\omega}:=d\tilde{\phi}$ satisfies the corresponding lifted equation up to lower–order controlled terms, which can be absorbed into the energy estimates.
+And applying the exterior derivative to (69) yields exactly (79) with a solution reconstructed also from
 $\omega(x,t)=\int_{S_{x}^{*}M}\tilde{\omega}(x,t,\xi)\,d\xi$ (87)
@@
 Viscosity, compactness, and the geometric nature of the lift combine to impose strong restrictions on such instabilities.
-Throughout this section, $\lambda\in\{\phi,\omega\}$ denotes either the velocity $1$-form or the vorticity $2$-form on $M$, and $\tilde{\lambda}$ its microlocal lift to $S^{*}M$.
-All norms and contractions are taken with respect to the effective metric $g^{\prime}$ introduced earlier.
-Since $S^{*}M$ is compact and endowed with a canonical contact structure, all integrals below are well-defined without further assumptions.
-We begin by defining the fundamental local geometric quantity governing the analysis. The microlocal amplitude of $\tilde{\lambda}$ is defined pointwise by the contraction
-$|\tilde{\lambda}|_{g^{\prime}}^{2}:=\langle\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}=\iota_{\tilde{\lambda}^{\sharp}}\tilde{\lambda}$, (65)
+For those reasons, throughout this section, $(M,g)$ denotes a smooth Riemannian manifold and $S^{*}M$ its unit cosphere bundle endowed with the canonical Liouville measure $\mu_{S^{*}M}$.
+We assume that the velocity field satisfies
+$u\in L^{\infty}([t_{0},t];W^{2,\infty}(M))$ (88)
+and that all geometric quantities associated with the effective connection $\nabla^{u}$ remain smooth on $t_{0},t]\times S^{*}M$.
+Let $\lambda\in\{\phi,\omega\}$ denote either the velocity $1$-form or the vorticity $2$-form on $M$, and $\tilde{\lambda}$ its microlocal lift to $S^{*}M$.
+All contractions and norms are taken with respect to the effective metric $g^{\prime}$.
+Defining the fundamental local geometric quantity governing the analysis, the microlocal amplitude of $\tilde{\lambda}$ is defined pointwise by the contraction
+$\|\tilde{\lambda}\|_{g^{\prime}}^{2}:=\langle\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}=\iota_{\tilde{\lambda}^{\sharp}}\tilde{\lambda}$ (89)
 where $\tilde{\lambda}^{\sharp}$ denotes the metric dual with respect to $g^{\prime}$.
@@
 $\mu_{S^{*}M}=\alpha\wedge(d\alpha)^{n-1}$ (90)
-where $\alpha$ is the contact form induced from the Liouville $1$-form on $T^{*}M$. We define the microlocal energy functional
-$\mathcal{V}_{\lambda}(t):=\int_{S^{*}M}|\tilde{\lambda}(x,\xi,t)|_{g^{\prime}}\,\mu_{S^{*}M}$ (67)
-Geometrically, $\mathcal{V}_{\lambda}$ measures the total directional spread of the field $\lambda$ in phase space.
-Physically, it quantifies how much the solution concentrates along preferred cotangent directions, independently of its spatial localization.
-We now derive a balance law for $\mathcal{V}_{\lambda}$. The microlocal evolution equation satisfied by $\tilde{\lambda}$ is
-$\partial_{t}\tilde{\lambda}+\mathcal{L}_{X_{u}}\tilde{\lambda}+\nu\|\xi\|_{g}^{2}\,\tilde{\lambda}=\tilde{R}[\tilde{\lambda}]+\tilde{C}[\tilde{\lambda}]+\tilde{F}$ (68)
-where $X_{u}$ is the canonical lifted transport field on $S^{*}M$. We take the pointwise inner product with $\tilde{\lambda}$ and use bilinearity to obtain
-$\frac{1}{2}\partial_{t}|\tilde{\lambda}|_{g^{\prime}}^{2}+\frac{1}{2}\mathcal{L}_{X_{u}}|\tilde{\lambda}|_{g^{\prime}}^{2}+\nu\|\xi\|_{g}^{2}|\tilde{\lambda}|_{g^{\prime}}^{2}\\ =\langle\tilde{R}[\tilde{\lambda}],\tilde{\lambda}\rangle+\langle\tilde{C}[\tilde{\lambda}],\tilde{\lambda}\rangle+\langle\tilde{F},\tilde{\lambda}\rangle$ (69)
-Each term on the right-hand side admits a sharp geometric bound. Since the curvature term is zeroth order in $\xi$,
-$|\langle\tilde{R}[\tilde{\lambda}],\tilde{\lambda}\rangle|\leq\|\mathrm{Ric}\|_{L^{\infty}(M)}\,\|\tilde{\lambda}\|_{g^{\prime}}^{2}$ (70)
+where $\alpha$ is the contact form induced from the Liouville $1$-form on $T^{*}M$.
+V.1 Energy Functionals
+The microlocal evolution of $\tilde{\lambda}$ is governed by the equation
+$\partial_{t}\tilde{\lambda}+\mathcal{L}_{X_{u}}\tilde{\lambda}+\nu\|\xi\|_{g^{\prime}}^{2}\,\tilde{\lambda}=\tilde{R}[\tilde{\lambda}]+\tilde{C}[\tilde{\lambda}]+\tilde{F}$ (91)
+where $X_{u}$ denotes the canonical lifted transport vector field on $S^{*}M$, $\tilde{R}$ is a zeroth-order curvature operator induced by the effective connection, $\tilde{C}$ collects commutator terms arising from microlocalization, and $\tilde{F}$ represents external forcing.
+Doing a energy functional approach to this equation to have boundary conditions, we define the lifted quadratic energy functional as
+$\mathcal{E}(t):=\frac{1}{2}\int_{S^{*}M}\langle\tilde{\lambda}(t),\tilde{\lambda}(t)\rangle_{g^{\prime}}\,\mu_{S^{*}M}$ (92)
+Differentiating in time along solutions of (91) yields
+$\frac{d}{dt}\mathcal{E}(t)$ $=\int_{S^{*}M}\langle\partial_{t}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}+\frac{1}{2}\int_{S^{*}M}\partial_{t}\langle\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}$ (93)
+If the effective metric $g^{\prime}$ is fixed or depends smoothly on $u$, the second term is controlled by $\|\nabla u\|_{L^{\infty}}$ and can be absorbed into the estimates below.
+Using (91), we estimate each contribution separately. The transport term satisfies
+$\left\|\int_{S^{*}M}\langle\mathcal{L}_{X_{u}}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}\right\|\leq\sigma\|\nabla u\|_{L^{\infty}}\|\tilde{\lambda}\|_{L^{2}(S^{*}M)}^{2}$ (94)
+while the dissipative term is nonnegative,
+$\int_{S^{*}M}\langle\nu\|\xi\|_{g^{\prime}}^{2}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}=\nu\int_{S^{*}M}\|\xi\|_{g^{\prime}}^{2}\|\tilde{\lambda}\|_{g^{\prime}}^{2}\,\mu_{S^{*}M}\geq 0$ (95)
+The curvature term admits the pointwise bound
+$\|\langle\tilde{R}[\tilde{\lambda}],\tilde{\lambda}\rangle_{g^{\prime}}\|\leq\|\operatorname{Ric}^{u}\|_{L^{\infty}(S^{*}M)}\|\tilde{\lambda}\|_{g^{\prime}}^{2}$ (96)
 The commutator term, arising from non-exact microlocalization, satisfies
-$|\langle\tilde{C}[\tilde{\lambda}],\tilde{\lambda}\rangle|\leq C\,\|\nabla u\|_{L^{\infty}(M)}\,\|\tilde{\lambda}\|_{g^{\prime}}^{2}$ (71)
+$\|\langle\tilde{C}[\tilde{\lambda}],\tilde{\lambda}\rangle\|\leq\sigma\,\|\nabla u\|_{L^{\infty}(M)}\,\|\tilde{\lambda}\|_{g^{\prime}}^{2}$ (97)
 Finally, the forcing term is controlled linearly,
-$|\langle\tilde{F},\tilde{\lambda}\rangle|\leq\|\tilde{F}\|\,\|\tilde{\lambda}\|_{g^{\prime}}$ (72)
-We now integrate the balance identity over $S^{*}M$. Since the Liouville measure is invariant under the flow generated by $X_{u}$, that is,
+$\|\langle\tilde{F},\tilde{\lambda}\rangle\|\leq\|\tilde{F}\|\,\|\tilde{\lambda}\|_{g^{\prime}}$ (98)
+Collecting all contributions, we obtain
+$\frac{d}{dt}\mathcal{E}(t)\leq\sigma\bigl(1+\|\nabla u\|_{L^{\infty}}+\|\nabla^{2}u\|_{L^{\infty}}\bigr)\mathcal{E}(t)$ (99)
+where $\sigma>0$ depends only in the estimates depend on $\|\nabla u\|_{L^{\infty}}$ and the background curvature, but remain bounded on finite time intervals under the regularity assumptions considered.
+By Grönwall’s inequality,
+$\mathcal{E}(t)\leq\mathcal{E}(t_{0})\exp\!\left(\sigma\int_{t_{0}}^{t}\bigl(1+\|\nabla u(s)\|_{L^{\infty}}\\ +\|\nabla^{2}u(s)\|_{L^{\infty}}\bigr)\,ds\right)$ (100)
+In particular, as long as the velocity field $u$ remains bounded in $W^{2,\infty}$, the lifted energy $\mathcal{E}(t)$ cannot blow up in finite time.
+Moreover, the dissipative term provides additional control of high–frequency microlocal components.
+To quantify those directional concentration effects beyond quadratic energy, we introduce the microlocal functional
+$\mathcal{V}_{\lambda}(t):=\int_{S^{*}M}\|\tilde{\lambda}(x,\xi,t)\|_{g^{\prime}}\,\mu_{S^{*}M}$ (101)
+Since the Liouville measure is invariant under the flow generated by $X_{u}$,
 $\mathcal{L}_{X_{u}}\mu_{S^{*}M}=0$ (102)
-the transport term drops out after integration. We obtain
-$\frac{d}{dt}\mathcal{V}_{\lambda}(t)+\nu\int_{S^{*}M}\|\xi\|_{g}^{2}\|\tilde{\lambda}\|_{g^{\prime}}^{2}\,\mu_{S^{*}M}\leq C\|\nabla u\|_{L^{\infty}}\mathcal{V}_{\lambda}(t)\\ +\|\tilde{F}\|_{L^{1}(S^{*}M)}$ (74)
-This inequality expresses a geometric fact, microlocal energy cannot accumulate without being dissipated, unless the deformation of the flow becomes uncontrolled.
-In particular, in the absence of forcing, $\mathcal{V}_{\lambda}$ is monotone up to a deformation-controlled factor.
-To make this mechanism more explicit, recall that the velocity gradient decomposes as $\nabla u=S+\Omega$, where only the symmetric part $S$ contributes to stretching.
-Restricting the evolution to integral curves of $X_{u}$ yields the pointwise inequality
+integration of the pointwise balance law associated with (91) yields
+$\frac{d}{dt}\mathcal{V}_{\lambda}(t)+\nu\int_{S^{*}M}\|\xi\|_{g^{\prime}}^{2}\|\tilde{\lambda}\|_{g^{\prime}}^{2}\,\mu_{S^{*}M}\\ \leq\sigma\|\nabla u\|_{L^{\infty}}\mathcal{V}_{\lambda}(t)+\|\tilde{F}\|_{L^{1}(S^{*}M)}$ (103)
+This inequality expresses a purely geometric fact, the microlocal energy cannot accumulate along cotangent directions without being dissipated, unless the symmetric deformation of the flow becomes dominant.
+Decomposing $\nabla u=S+\Omega$ into symmetric and antisymmetric parts, one obtains along integral curves of $X_{u}$ the pointwise inequality
 $\frac{d}{dt}\|\tilde{\lambda}\|\leq\|S(t)\|_{L^{\infty}}\|\tilde{\lambda}\|-\nu\|\tilde{\lambda}\|$ (104)
 with solution
-$\|\tilde{\lambda}(t)\|\leq\|\tilde{\lambda}(0)\|\exp\!\left[\int_{t_{0}}^{t}(\|S(s)\|_{L^{\infty}}-\nu)\,ds\right]$ (76)
-showing that any amplification of microlocal amplitude requires a persistent failure of temporal integrability of $\|S\|_{L^{\infty}}$.
+$\|\tilde{\lambda}(t)\|\leq\|\tilde{\lambda}(t_{0})\|\exp\!\left[\int_{t_{0}}^{t}(\|S(s)\|_{L^{\infty}}-\nu)\,ds\right]$ (105)
+[…]
+Metric compatibility and control of geometric defects
+A subtle but important feature of the present development is that the effective connection $\nabla^{u}$ introduced in Section III is, in general, not compatible with the effective metric $g^{\prime}$.
+More precisely, the non-metricity tensor
+$(\nabla^{u}g^{\prime})_{ijk}:=\partial_{i}g^{\prime}_{jk}-(\Gamma^{u})^{l}{}_{ij}g^{\prime}_{lk}-(\Gamma^{u})^{l}{}_{ik}g^{\prime}_{jl}$ (106)
+does not vanish identically unless additional structural conditions on the velocity field are imposed.
+Recalling that $(\Gamma^{u})^{k}{}_{ij}=\Gamma^{k}{}_{ij}+A^{k}{}_{ij}$, we compute
+$(\nabla^{u}g^{\prime})_{ijk}=-\bigl(A^{l}{}_{ij}g^{\prime}_{lk}+A^{l}{}_{ik}g^{\prime}_{jl}\bigr)$ (107)
+which shows that the failure of metric compatibility is entirely encoded in the deformation tensor $A(u)$ associated with the velocity field.
+In particular, $\nabla^{u}$ is metric-compatible with respect to $g^{\prime}$ if and only if
+$A^{l}{}_{ij}g^{\prime}_{lk}+A^{l}{}_{ik}g^{\prime}_{jl}=0$ (108)
+a condition that is satisfied, for instance, when $u$ is covariantly constant.
+Although $\nabla^{u}$ is not metric-compatible in general, this defect does not obstruct the energy method developed in this section.
+Indeed, since $A(u)$ is linear in $\nabla u$, the non-metricity tensor satisfies the pointwise bound
+$\|\nabla^{u}g^{\prime}\|_{L^{\infty}(S^{*}M)}\leq\sigma\,\|\nabla u\|_{L^{\infty}(M)}$ (109)
+with $\sigma$ depending only on the background geometry.
+As a consequence, all error terms generated by the lack of metric compatibility enter the energy balance as lower-order contributions and are absorbed into the constants appearing in the Grönwall inequality.
+This shows that the geometric dissipation mechanism remains robust despite the presence of controlled non-metricity.
+Reconstruction of the physical solution
+The microlocal formulation developed above must ultimately be compatible with the reconstruction of physical solutions on the base manifold $M$.
+For scalar quantities, the reconstruction is performed fiberwise by
+$\phi(x,t)=\int_{S_{x}^{*}M}\tilde{\phi}(x,t,\xi)\,d\xi$ (110)
+where $d\xi$ denotes the canonical measure on the unit cotangent sphere at $x$.
+Since $S_{x}^{*}M$ is compact and the microlocal lift $\tilde{\phi}$ is smooth in $(x,\xi)$ under the assumptions of Section V, the integral (110) is well-defined pointwise for all $(x,t)$.
+Moreover, by Hölder’s inequality and the definition of the microlocal functional $\mathcal{V}_{\phi}$, one has
+$\|\phi(x,t)\|\leq\int_{S_{x}^{*}M}\|\tilde{\phi}(x,t,\xi)\|\,d\xi\leq\sigma\,\mathcal{V}_{\phi}(t)$ (111)
+with $\sigma$ depending only on the dimension.
+Differentiation under the integral sign is justified by the smooth dependence of $\tilde{\phi}$ on $x$, which yields
+$\|\nabla^{k}\phi(\cdot,t)\|_{L^{\infty}(M)}\leq\sigma_{k}\sup_{(x,\xi)\in S^{*}M}\|\nabla_{x}^{k}\tilde{\phi}(x,t,\xi)\|$ (112)
+for all integers $k\geq 0$.
+Consequently, uniform microlocal bounds on $S^{*}M$ propagate directly to uniform regularity bounds for the reconstructed physical solution.
+This shows that the microlocal framework is not merely auxiliary, but faithfully encodes the regularity properties of the original Navier–Stokes dynamics.
+Geometric classification of admissible blow-up mechanisms and control of vortex stretching
 We now specialize to the vorticity $\tilde{\omega}$, whose geometric structure allows for finer control. Since $\tilde{\omega}$ is a closed microlocal $2$-form, we introduce the functional
-$W[\tilde{\omega}]=\int_{S^{*}M}\left(\tau\,|d^{\perp}\log|\tilde{\omega}||^{2}+\log|\tilde{\omega}|\right)\mu_{S^{*}M}$ (77)
+$W[\tilde{\omega}]=\int_{S^{*}M}\left(\tau\,\|d^{\perp}\log\|\tilde{\omega}\|\|^{2}+\log\|\tilde{\omega}\|\right)\mu_{S^{*}M}$ (113)
 where $d^{\perp}$ denotes the vertical exterior derivative induced by the horizontal–vertical splitting of $TS^{*}M$, and $\tau>0$ is a dimensional parameter.
 This functional penalizes sharp angular gradients of vorticity and quantifies directional irregularity.
+This functional $W[\tilde{\omega}$] admits a natural interpretation as a microlocal action measuring the angular regularity of vorticity on the compact phase space $S^{*}M$.
+The logarithmic term $\log\|\tilde{\omega}\|$ plays the role of an entropic density associated with the distribution of vorticity along cotangent directions, while the vertical gradient term $\|d^{\perp}\log\|\tilde{\omega}\|\|^{2}$ penalizes sharp directional concentration on the fibers $S_{x}^{*}M$.
+From a dynamical viewpoint, the lifted flow generated by $\partial_{t}+\mathcal{L}_{X_{u}}$ defines a family of microlocal trajectories whose evolution is governed by a non-metric, torsion-bearing effective connection induced by the velocity field.
+The commutator structure developed throughout the paper shows that curvature, torsion, and non-metricity enter the dynamics as lower-order geometric forces.
+In this sense, $W$ can be viewed as an effective action functional on the space of microlocal paths, whose monotonicity expresses the suppression of highly curved or torsion-dominated angular trajectories by viscous dissipation.
+Persistent directional concentration would therefore require a breakdown of this geometric diffusion mechanism, linking potential singular behavior to a loss of microlocal dissipative stability rather than to amplitude growth alone.
 A direct differentiation in time, combined with integration by parts along the fibers [30] and the compactness of $S^{*}M$, yields the dissipation inequality
-$\frac{d}{dt}W[\tilde{\omega}]\leq\int_{S^{*}M}\left(\tau\,|d^{\perp}\log|\tilde{\omega}||^{2}-\tau\,\|S\|\,|\tilde{\omega}|-\nu\right)\mu_{S^{*}M}$ (78)
+$\frac{d}{dt}W[\tilde{\omega}]\leq\int_{S^{*}M}\left(\tau\,\|d^{\perp}\log\|\tilde{\omega}\|\|^{2}-\tau\,\|S\|\,\|\tilde{\omega}\|-\nu\right)\mu_{S^{*}M}$ (114)
 This inequality shows that viscosity and geometric diffusion suppress directional oscillations of vorticity, unless sustained stretching dominates dissipation.
-Since $S^{*}M$ is compact, the above estimates imply uniform control of $\tilde{\omega}$ in $L^{\infty}(S^{*}M)$, and therefore of the reconstructed physical vorticity [4].
-Consequently, any loss of regularity must be accompanied by a breakdown of one of the assumptions used above.
-We arrive at a geometric classification of admissible blow-up mechanisms [8]. A singularity can occur only if at least one of the following phenomena takes place,
+Since $S^{*}M$ is compact, the estimates derived above yield uniform control of the microlocal vorticity $\tilde{\omega}$ on $S^{*}M$.
+In particular, as long as the assumptions of Section V remain valid, no unbounded growth of $\tilde{\omega}$ can occur along the lifted dynamics.
+By the reconstruction formula, this microlocal control propagates to the physical vorticity field on $M$ [4].
+Consequently, any loss of regularity of the Navier–Stokes solution must be accompanied by a breakdown of at least one of the hypotheses employed in the microlocal framework.
+This observation highlights one of the central difficulties in the analysis of three-dimensional Navier–Stokes flows: the vortex stretching mechanism, which is responsible for potential amplification of vorticity.
+Recalling the classical decomposition of the velocity gradient, $\nabla u=S+\Omega$ into its symmetric and antisymmetric parts, only the symmetric tensor $S$ contributes to stretching effects.
+Within the microlocal formulation, the evolution of $\tilde{\omega}$ along integral curves of the lifted transport field $X_{u}$ satisfies the pointwise differential inequality
+$\frac{d}{dt}\|\tilde{\omega}\|\leq\|S(t)\|_{L^{\infty}}\,\|\tilde{\omega}\|-\nu\,\|\tilde{\omega}\|$ (115)
+which makes explicit the competition between stretching and viscous dissipation. Integrating this inequality yields
+$\|\tilde{\omega}(t)\|\leq\|\tilde{\omega}(t_{0})\|\exp\!\left[\int_{t_{0}}^{t}\bigl(\|S(s)\|_{L^{\infty}}-\nu\bigr)\,ds\right]$ (116)
+This bound remains meaningful even in the vanishing viscosity limit $\nu\to 0$, in which case any amplification of vorticity necessarily requires uncontrolled growth of the time integral of $\|S\|_{L^{\infty}}$.
+The vortex stretching mechanism is further constrained by the monotone microlocal volume functional $W[\tilde{\omega}$] introduced earlier.
+Since $S^{*}M$ is compact, the dissipation inequality satisfied by $W[\tilde{\omega}$] implies that sharp angular concentration of vorticity in phase space cannot persist unless accompanied by a loss of temporal integrability of the symmetric deformation, namely
+$\int_{t_{0}}^{t}\|S(s)\|_{L^{\infty}}\,ds=\infty$ (117)
+In particular, extreme directional focusing of vorticity in the cosphere bundle is geometrically incompatible with sustained dissipation.
+This shows that vortex stretching, while necessary for potential singular behavior, is subject to strong geometric and microlocal constraints imposed by viscosity, compactness, and the structure of the lifted dynamics.
+We therefore obtain a geometric exclusion principle: any blow-up scenario must simultaneously overcome viscous damping, geometric diffusion on the cosphere bundle, and the monotonicity of microlocal volume invariants.
+This reformulates vortex stretching not as a purely analytic instability, but as a highly constrained geometric phenomenon.
+And we arrive at a geometric classification of admissible blow-up mechanisms [8].
+A singularity can occur only if at least one of the following phenomena takes place,
   • Extreme directional concentration of microlocal vorticity in $S^{*}M$.
@@
 Blow-up is not a purely analytic accident, but a geometric instability that must overcome strong dissipative and structural constraints.
-VI Discussion and Conclusions
+VI Main Results
+For clarity, we collect here the main mathematical results established throughout the paper.
+The proofs are provided in the corresponding Sections II–V, where the relevant geometric and microlocal structures are introduced and developed.
+Theorem 1 (Microlocal geometric reformulation).
+The three-dimensional incompressible Navier–Stokes equations admit an equivalent formulation as a linear, non-autonomous, dissipative evolution equation on the compact cosphere bundle $S^{*}M$, acting on microlocal lifts of differential forms and driven by a canonical transport vector field $X_{u}$ associated with the velocity field.
+Proposition 1 (effective connection).
+Let $(M,g)$ be a Riemannian manifold and let $u$ be a sufficiently regular vector field. Define the effective connection $\nabla^{u}$ by
+$(\nabla^{u})_{i}v^{k}:=\nabla_{i}v^{k}+A^{k}{}_{ij}v^{j}$
+where $A^{k}{}_{ij}$ is a (1,2)(1,2)–tensor depending smoothly on $u$. Then $\nabla^{u}$ defines a linear connection on $TM$ whose Christoffel symbols are given by
+$(\Gamma^{u})^{k}{}_{ij}=\Gamma^{k}{}_{ij}+A^{k}{}_{ij}$
+Theorem 2 (Effective torsion induced by the velocity field).
+Assume that the deformation tensor admits the decomposition
+$A^{k}{}_{ij}=g^{\prime k\ell}g^{\prime}_{im}\nabla_{j}u^{m}+B^{k}{}_{ij},\qquad B^{k}{}_{ij}=B^{k}{}_{ji}$
+Then the torsion tensor $T^{u}$ of $\nabla^{u}$ is given by
+$(T^{u})^{k}{}_{ij}=g^{\prime k\ell}\left(g^{\prime}_{im}\nabla_{j}u^{m}-g^{\prime}_{jm}\nabla_{i}u^{m}\right)$
+In particular, $T^{u}$ is antisymmetric in $(i,j)$ and depends linearly on $\nabla u$.
+Corollary 1 (Torsion–free condition).
+The effective connection $\nabla^{u}$ is torsion–free if and only if the velocity field $u$ is covariantly constant.
+Theorem 3 (Controlled geometry of the effective connection).
+Let $\nabla^{u}$ be the effective affine connection induced by the velocity field $u$.
+Its torsion and non-metricity tensors are of first order in $\nabla u$ and satisfy uniform bounds on $S^{*}M$ under the regularity assumptions considered.
+Theorem 4 (Controlled non-metricity of the effective connection).
+Let $(M,g)$ be a smooth Riemannian manifold and let $u\in L^{\infty}([t_{0},t];W^{1,\infty}(M))$. Let $g^{\prime}$ denote the effective metric on $S^{*}M$. Then the non-metricity tensor
+$Q^{u}:=\nabla^{u}g^{\prime}$
+satisfies the pointwise bound
+$\|Q^{u}\|_{L^{\infty}(S^{*}M)}\leq\sigma\|\nabla u\|_{L^{\infty}(M)}$
+for a constant $\sigma$ depending only on the background geometry. In particular, the lack of metric compatibility produces only lower–order contributions in the microlocal energy estimates.
+Theorem 5 (Validity of the microlocal reconstruction).
+Let $\tilde{\phi}(x,\xi,t)$ be a microlocal lift on $S^{*}M$ satisfying
+$\tilde{\phi}\in L^{\infty}([t_{0},t];L^{2}(S^{*}M))\cap L^{\infty}(S^{*}M).$
+Then the reconstruction formula
+$\phi(x,t)=\int_{S^{*}_{x}M}\tilde{\phi}(x,\xi,t)\,d\xi$
+is well-defined for almost every $x\in M$, and the reconstructed field $\phi$ belongs to $L^{\infty}([t_{0},t];L^{2}(M))$.
+Theorem 6 (Vortex energy as a geometric dissipation functional).
+Let $(M,g)$ be a smooth Riemannian manifold and let $\phi\in\Omega^{1}(M)$ be a divergence–free velocity 11–form with vorticity $\omega=d\phi$. Define the vortex energy functional by
+$\mathcal{E}_{\omega}[\phi]:=\frac{1}{2}\int_{M}\|\omega\|^{2}\,dV$ (118)
+Then $\mathcal{E}_{\omega}$ is invariant under gauge transformations $\phi\mapsto\phi+d\psi$ compatible with incompressibility, and its $L^{2}$–variational gradient is given by
+$\frac{\delta\mathcal{E}_{\omega}}{\delta\phi}=\delta d\phi=\Delta_{H}\phi$
+Moreover, along smooth solutions of the incompressible Navier–Stokes equations, the vortex energy satisfies the dissipation identity
+$\frac{d}{dt}\mathcal{E}_{\omega}(t)=-\nu\int_{M}\|\nabla\omega\|^{2}\,dV\leq 0$
+In particular, viscosity induces a strict monotone decay of the total vorticity intensity.
+Theorem 7 (Lifted microlocal evolution equation).
+Let $\tilde{\lambda}$ be a microlocal differential form on $S^{*}M$. Then its evolution is governed by
+$\partial_{t}\tilde{\lambda}+\mathcal{L}_{X_{u}}\tilde{\lambda}+\nu\|\xi\|_{g^{\prime}}^{2}\,\tilde{\lambda}=\tilde{R}[\tilde{\lambda}]+\tilde{C}[\tilde{\lambda}]+\tilde{F}$
+where $\tilde{R}$ is a zeroth–order curvature–induced operator depending on the effective geometry, $\tilde{C}$ represent commutator terms arising from microlocalization, and $\tilde{F}$ represents external forces considered into the system.
+Theorem 8 (Geometric energy inequality).
+Let $(M,g)$ be a smooth Riemannian manifold and let $u\in L^{\infty}([t_{0},t];W^{2,\infty}(M))$ be divergence–free. Let $\tilde{\lambda}$ solve the lifted evolution equation and define
+$\mathcal{E}(t):=\frac{1}{2}\int_{S^{*}M}\langle\tilde{\lambda}(t),\tilde{\lambda}(t)\rangle_{g^{\prime}}\,d\mu$
+Then there exists $\sigma>0$, depending only on the background geometry, such that
+$\mathcal{E}(t)\leq\mathcal{E}(t_{0})\exp\!\left[\sigma\int_{t_{0}}^{t}\bigl(1+\|\nabla u(s)\|_{L^{\infty}}+\|\nabla^{2}u(s)\|_{L^{\infty}}\bigr)\,ds\right]$
+Corollary 2 (Geometric regularity criterion).
+As long as the velocity field $u$ remains bounded in $W^{2,\infty}(M)$, the lifted energy $\mathcal{E}(t)$ cannot blow up in finite time.
+Theorem 9 (Microlocal entropy functional and exclusion of directional concentration).
+Let $\tilde{\omega}$ be the microlocal lift of the vorticity 22–form to the compact cosphere bundle $S^{*}M$. Define the microlocal entropy functional
+$W[\tilde{\omega}]=\int_{S^{*}M}\left(\tau\,\|d^{\perp}\log\|\tilde{\omega}\|\|^{2}+\log\|\tilde{\omega}\|\right)\,d\mu,\qquad\tau>0.$ (119)
+Then, along smooth solutions of the lifted Navier–Stokes dynamics, the functional $W[\tilde{\omega}$] satisfies a dissipation inequality of the form
+$\frac{d}{dt}W[\tilde{\omega}]\leq-\nu-\int_{S^{*}M}\bigl(\tau\,\|S\|\,\|\tilde{\omega}\|\bigr)\,d\mu$
+where $S$ denotes the symmetric part of the velocity gradient.
+In particular, on the compact phase space $S^{*}M$, persistent sharp angular concentration of vorticity is incompatible with viscous dissipation unless accompanied by a loss of temporal integrability of $\|S\|_{L^{\infty}}$.
+Theorem 10 (Geometric exclusion of microlocal blow-up).
+Assume that the symmetric deformation tensor satisfies
+$\int_{t_{0}}^{t}\|S(t)\|_{L^{\infty}(M)}\,dt<\infty$
+Then no solution of the microlocal Navier–Stokes system can develop persistent directional concentration on $S^{*}M$.
+Corollary 3 (Geometric classification of blow-up mechanisms).
+Any finite–time singularity of the three–dimensional Navier–Stokes equations must be accompanied by at least one of the following phenomena,
+  • Loss of temporal integrability of $\|\nabla u\|_{L^{\infty}}$.
+  • Unbounded symmetric stretching $\|S\|_{L^{\infty}}$.
+  • Extreme directional concentration of microlocal vorticity on $S^{*}M$.
+All other blow-up scenarios are excluded by geometric dissipation on the compact phase space.
+While the present framework is formulated for smooth solutions, it provides a geometric guideline for the microlocal analysis of weak solutions, which will be addressed elsewhere.
+VII Discussion and Conclusions
 The theoretical approach developed in this work addresses the Navier–Stokes regularity problem through a synthesis of microlocal analysis and Riemannian geometry.
@@
 The microlocal transport–dissipation structure is consistent with the analysis of lifted operators and unique continuation phenomena developed in [29], and aligns with energy–aware geometric decompositions of the Navier–Stokes system [6].
+From a mathematical standpoint, the results collected in Section VI formalize this perspective through a sequence of geometric theorems that constrain the admissible mechanisms of singularity formation.
+Rather than providing a direct regularity proof, these results establish monotone or coercive geometric quantities—both at the physical and microlocal levels—whose breakdown is necessary for loss of regularity.
+In particular, the vortex energy functional yields a classical dissipative control, while the microlocal volume and directional functionals exclude persistent angular concentration on the compact phase space $S^{*}M$ unless accompanied by a loss of temporal integrability of the symmetric deformation tensor.
+Together, these theorems recast the regularity problem as a question of geometric and microlocal stability, rather than purely analytic growth of classical norms.
 For that, this paper does not solve the Navier–Stokes regularity problem.
@@
 This reformulation clarifies how geometric dissipation and microlocal transport can coexist with modern flexibility results, while still yielding meaningful regularity criteria for physically relevant solutions.
+Beyond the analytic and geometric implications discussed above, the lifted formulation also suggests a broader operator-theoretic interpretation of the Navier–Stokes dynamics, which we briefly outline below.
+The structural consistency of the microlocal formulation developed in this work, particularly its stability properties in the vanishing viscosity limit $\nu\to 0$, suggests a natural operator-theoretic interpretation and a possible connection with geometric deformation frameworks.
+By lifting the Navier–Stokes dynamics to the compact cosphere bundle $S^{*}M$, the transport–dissipation system is recast as a linear evolution equation acting on microlocal distributions, governed by a geometrically defined generator involving transport, curvature, and diffusion terms.
+From this viewpoint, the lifted amplitudes $\tilde{\phi}$ and $\tilde{\omega}$ are more naturally interpreted as evolving distributional states on phase space rather than as classical fields on $M$.
+Within this setting, the effective connection $\nabla^{u}$ and its associated curvature encode nonlinear self-interaction effects through purely geometric operators, while the viscous contribution $\nu\|\xi\|_{g^{\prime}}^{2}$ induces a robust suppression of high-frequency microlocal components.
+This mechanism is formally analogous to decoherence phenomena in operator evolution problems, where diffusion terms stabilize the dynamics by damping highly oscillatory modes.
+In particular, the geometric energy inequalities derived in this work can be interpreted as spectral stability estimates for the lifted evolution operator, showing that microlocal concentration is incompatible with sustained dissipation on the compact phase space.
+Although no quantization procedure is developed here [34, 38], the geometric structure of the lifted dynamics provides a natural setting for future exploration via deformation quantization or geometric operator methods, in which nonlinear fluid interactions manifest as curvature-driven corrections to linear transport on $S^{*}M$.
+From this perspective, classical regularity emerges as a consequence of persistent microlocal stability enforced by geometric dissipation, rather than as a purely local analytic phenomenon.
 A complementary geometric interpretation is obtained by considering the lifted dynamics on the extended manifold $S^{*}M\times N$, where $N\subset\mathbb{R}$ denotes the temporal interval.
@@
 These ideas may be useful in turbulence modeling, data-driven reduced-order models, and multi-scale simulations in cluster environments where know and preserve the appropriate geometric properties and its structure is essential for long-time accuracy.
-VII AI Statement
+VIII AI Statement
 The author confirms that every intellectual contribution, original mathematical and physical derivations, theorem, proofs, and physical hypotheses presented in this manuscript are the sole work of the author.
@@
 The author has reviewed, edited, and validated the accuracy of all output generated by the AI.
+IX Conflict of Interest
+The author declares that there are no conflicts of interest regarding the publication of this paper.
+X Data Availability
+Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.
 Acknowledgements.
 The author is profoundly grateful to PhD.
-Pedro Díaz-Navarro because of his openess and creativity to talk.
-In the same way as Octavio Rojas-Quesada for every new point of view and Celicia Ledezma-Porras for her geometric insight some years ago.
-Moreover, family, some friends and every guide, person… who has stumble with me along the paths, thank you all.
+Pedro Díaz-Navarro for his openness and creative discussions.
+Likewise to Octavio Rojas-Quesada for his valuable new points of view, and Celicia Ledezma-Porras for her geometric insight years ago.
+Moreover, to family, friends, and every guide and person who has stumbled along these paths with me, thank you all.

### v3

@@
 We develop a geometric and microlocal framework for the Navier–Stokes equations by lifting the dynamics to the cosphere bundle of a Riemannian manifold.
-In this formulation, the velocity field and vorticity are represented as microlocal distributions whose evolution is governed by a linear transport–dissipation system generated by a canonical dynamical vector field.
-We introduce microlocal amplitudes, directional energy functionals, and monotone volume invariants on the compact phase space, which quantify directional concentration and alignment mechanisms associated with potential loss of regularity.
-The viscous term induces an effective geometric diffusion on the cosphere bundle, yielding closed differential inequalities in a geometric setting.
-To capture the interaction between fluid deformation and geometry, we define an effective connection and curvature tensor encoding the influence of the symmetric velocity gradient.
-This structure gives rise to a Ricci–type microlocal geometric evolution that constrains directional stretching and excludes extreme angular concentration compatible with viscous dissipation.
-While the present results do not resolve the global regularity problem, they provide a coherent geometric mechanism that severely restricts admissible blow–up scenarios, reformulating the regularity question as a problem of dissipative stability on a compact phase space.
+In this formulation, the velocity field and vorticity are represented as microlocal distributions whose evolution is governed by a linear, non-autonomous transport–dissipation system generated by a canonical dynamical vector field on the compact phase space $S^{*}M$.
+We introduce microlocal amplitude functionals, a directional entropy functional, and monotone geometric invariants that quantify angular concentration and alignment mechanisms associated with potential loss of regularity.
+The viscous term induces an effective geometric diffusion on the cosphere bundle, yielding closed differential inequalities expressed in intrinsic geometric form.
+To capture the interaction between fluid deformation and geometry, we define an effective connection and an associated curvature tensor encoding the influence of the symmetric velocity gradient, giving rise to a Ricci–type microlocal geometric evolution that constrains directional stretching.
+The framework establishes a necessary and sufficient geometric equivalence criterion: finite-time singularity occurs if and only if at least one of three microlocal controls fails—deformation integrability, entropy boundedness, or lifted energy boundedness.
+The symmetry structure reveals a dimensional reduction mechanism: as effective fiber dimension increases, the cosphere bundle undergoes a symmetry lock, forcing microlocal distributions toward isotropy and topologically obstructing the formation of angular singularities.
+While these results do not resolve the global regularity problem, they reformulate it as a question of dissipative stability and spectral control on a compact, symmetry-constrained microlocal dynamical system, severely restricting admissible blow–up scenarios through geometric and representation-theoretic constraints.
 Keywords
-Navier–Stokes equations, microlocal analysis, cosphere bundle, geometric regularity criteria, effective Ricci curvature, dissipative dynamical systems, differential forms.
+Navier–Stokes equations, microlocal analysis, cosphere bundle, geometric regularity criteria, effective Ricci curvature, dissipative dynamical systems, entropy functionals, symmetry lock mechanism.
 preprint: APS/123-QED
@@
 Geometric and covariant formulations on Riemannian manifolds provide a coherent framework independent of coordinates [7, 13].
-This clearly explains differential operators, the energy structure, the role of curvature, and the symmetries of the system through Killing fields [6].
-This is because Hamiltonian formulations and differential forms, using weighted and covariant tensors with exterior derivatives [30], ensure the method works naturally, making diffusion and boundary terms appear more clearly in components with non-zero curvature.
+This formulation provides a natural geometric interpretation of differential operators, the energy structure, curvature effects, and symmetries through Killing fields.
+[6].
+This is due to the fact that Hamiltonian formulations and differential forms, using weighted and covariant tensors with exterior derivatives [30], ensure the method works naturally, making diffusion and boundary terms appear more clearly in components with non-zero curvature.
 For cases of two spatial dimensions with certain geometries and specific boundary conditions, existence, uniqueness, and continuity are well understood through frequency decomposition and invariant methods [14].
@@
 In this work, we reformulate the incompressible Navier–Stokes equations as a dissipative dynamical system defined globally on the cosphere bundle $S^{*}M$.
-Our goal is to classify blow-up mechanisms and obtain geometric criteria for conditional regularity using variational and Fourier methods.
-In this way, we find inequalities and invariants to show how the physical solution is reconstructed from the dynamics in $S^{*}M$.
-Microlocal lifting approach
-The microlocal lifting to the cosphere bundle $S^{*}M$ adopted in this paper should be understood as a parametric refinement and compactification of the dynamics on the base manifold $M$.
-All geometric objects—including the effective connection $\nabla^{u}$, its torsion and curvature, and the metric $g^{\prime}$—are defined on $M$, with coefficients depending smoothly on $(x,\xi)\in S^{*}M$.
-The cotangent variable $\xi$ serves to encode directional and spectral information, but no geometric structure is imposed on the tangent bundle of $S^{*}M$ itself.
-This viewpoint allows us to combine microlocal analysis with intrinsic geometric evolution without introducing additional bundle-level connections.
+This formulation allows us to encode both spatial and directional information within a unified geometric framework.
+Our objective is to investigate the structure of potential blow–up mechanisms through microlocal and geometric methods, and to derive conditions under which directional concentration can be controlled.
+In particular, we introduce geometric functionals and differential inequalities that relate deformation, transport, and dissipation at the microlocal level.
+These constructions suggest a geometric equivalence principle in which finite–time singularity is linked to the breakdown of microlocal stability and deformation integrability.
+While the present work does not resolve the global regularity problem, it provides a coherent framework that restricts admissible singular behaviors and clarifies the geometric mechanisms underlying regularity.
+In what follows, we use $|\cdot|$ to denote pointwise norms or fiberwise magnitudes, and $\|\cdot\|$ to denote functional norms on the relevant spaces.
 II Covariant Formulation on Manifolds
@@
 Equivalently, in invariant tensorial form,
-$\nabla^{u}g^{\prime}=-\bigl(A(u)\cdot g^{\prime}+g^{\prime}\cdot A(u)\bigr)$ (41)
+$\nabla^{u}g^{\prime}=-\bigl[A(u)g^{\prime}+g^{\prime}A(u)\bigr]$ (41)
 where $A(u)$ is viewed as a (1,1)(1,1)–tensor acting on covariant indices.
@@
 By homogeneity, the field $X_{u}(x,\xi)$ is tangent to $S^{*}M$ and therefore defines a well-defined dynamics on the cosphere bundle, where $\|\xi\|_{g^{\prime}}$ is preserved.
-This will be useful because, at this point, it is beneficial to perform a Fourier transform on the system [5, 6].
-Since the classical Fourier transform is not globally defined on compact manifolds, we employ a microlocal transform defined locally in normal coordinates and extended globally as a function on $T^{*}M$.
-For a 1-form $\phi\in\Omega^{1}(M)$, we define
+This property motivates the introduction of a microlocal transform adapted to the compact setting.
+Since the classical Fourier transform is not globally defined on compact manifolds, we employ a microlocal transform defined locally in normal coordinates and extended globally as a function on $T^{*}M$ [5, 6].
+IV.1 Microlocal Fourier Transform and Tempered Distributions
+To place the microlocal lift in a fully controlled functional framework, we work in the space of tempered distributions.
+Let $M$ be a smooth $n$-dimensional manifold.
+In local charts, we define the Schwartz space
+$\mathcal{S}(\mathbb{R}^{n})=\Big\{f\in C^{\infty}(\mathbb{R}^{n}):\sup_{x\in\mathbb{R}^{n}}(1+|x|)^{N}|\partial^{\alpha}f(x)|<\infty,$ (68)
+$\forall N\in\mathbb{N},\alpha\in\mathbb{N}_{0}^{n}\Big\}$
+and its dual $\mathcal{S}^{\prime}(\mathbb{R}^{n})$ of tempered distributions.
+Globally, the space $\mathcal{S}^{\prime}(M)$ is defined as the set of distributions whose local restrictions are tempered:
+$\mathcal{S}^{\prime}(M)=\left\{u\in\mathcal{D}^{\prime}(M)\mid u\circ\varphi_{k}^{-1}\in\mathcal{S}^{\prime}(\mathbb{R}^{n})\forall k\in\{0,\dots,l\}\right\}$ (69)
+where the definition is independent of the particular choice of the atlas $\mathcal{A}$.
+For a 1-form $\phi\in\Omega^{1}(M)$ with sufficient regularity, we define the microlocal lift by
 $\tilde{\phi}(x,\xi)=\int_{T_{x}M}e^{-i\langle\xi,y\rangle}\phi(\exp_{x}y)\,\chi(y)\,dy$ (70)
-where $\chi$ is a smooth cutoff function.
-Restricting by homogeneity to $S^{*}M$, we obtain $\tilde{\phi}\in C^{\infty}(S^{*}M)$.
-The transform (68) should be understood as a directional microlocal projection rather than a pointwise invertible transform, because of that, the horizontal lift is uniquely determined once the effective connection is fixed, although it is not canonical with respect to the background metric.
-Applying the microlocal transform to the incompressible covariant system and projecting onto transverse directions to eliminate the pressure term, we obtain the microlocal equation
+where $\chi$ is a smooth cutoff function supported in a normal neighborhood of the origin in $T_{x}M$, and $\langle\xi,y\rangle$ denotes the natural pairing between covectors and vectors.
+The exponential map $\exp_{x}:T_{x}M\longmapsto M$ is well-defined in a neighborhood of zero, and the cutoff ensures that the integral is taken over a region where normal coordinates are valid.
+Remark 1 (Interpretation of the microlocal transform).
+The transform (70) should be understood as a directional microlocal projection rather than a pointwise invertible transform.
+It extracts the Fourier content of $\phi$ along the direction $\xi$ in the cotangent space.
+The horizontal lift of $\phi$ to $S^{*}M$ is uniquely determined once the effective connection is fixed, although it is not canonical with respect to the background metric alone.
+The microlocal Fourier transform in local coordinates is given by
+$\mathcal{F}_{x}u(\xi)=\int_{\mathbb{R}^{n}}e^{-ix\cdot\xi}u(x)\,dx$ (71)
+which extends continuously as a map
+$\mathcal{F}:\mathcal{S}^{\prime}(\mathbb{R}^{n})\longmapsto\mathcal{S}^{\prime}(\mathbb{R}^{n})$ (72)
+For $u\in H^{s}(\mathbb{R}^{n})$ with $s\in\mathbb{R}$, the Sobolev norm is characterized via the Fourier transform:
+$\|u\|_{H^{s}}^{2}=\int_{\mathbb{R}^{n}}(1+|\xi|^{2})^{s}|\widehat{u}(\xi)|^{2}\,d\xi$ (73)
+This shows that the microlocal lift inherits Sobolev control through fiberwise polynomial weights in the frequency variable $\xi$.
+Proposition 1 (Microlocal Sobolev inequality).
+For any $s\geq 0$ and $u\in H^{s}(M)$, the lifted object $\tilde{u}$ satisfies
+$\|\tilde{u}\|_{L^{2}(S^{*}M)}\leq\sigma_{s}\|u\|_{H^{s}(M)}$ (74)
+where $\sigma_{s}$ depends only on $s$ and the geometry of $M$.
+Proof.
+In local coordinates, the microlocal lift coincides with the restriction of the Fourier transform to the unit sphere in frequency space.
+By the trace theorem for Fourier multipliers and the coarea formula, restriction to $\|\xi\|_{g}=1$ lowers the Sobolev order by $1/2$.
+The global bound follows by summing over a finite coordinate cover using a partition of unity, noting that coordinate changes on $M$ induce bounded transformations on Sobolev spaces.
+∎
+This inequality ensures that microlocal regularity is controlled by classical Sobolev norms on the base manifold.
+IV.2 Microlocal Projection and Incompressibility
+Applying the microlocal transform to the incompressible covariant system and projecting onto transverse directions eliminates the pressure term. We obtain the microlocal equation
 $\partial_{t}\tilde{\phi}+\mathcal{L}_{X_{u}}\tilde{\phi}+\nu\|\xi\|_{g^{\prime}}^{2}\,\tilde{\phi}=\tilde{R}+\tilde{F},\qquad\langle\tilde{\phi},\xi\rangle_{g^{\prime}}=0$ (75)
-Equation (69) is a linear non-autonomous dissipative system on $S^{*}M$.
-Here, the term corresponding to the pressure $dP$ does not contribute to the system’s dynamics, acting as a Lagrange multiplier that imposes the incompressibility condition, existing only in the radial direction of the cotangent bundle.
-Therefore, we define the incompressible microlocal projector
+Equation (75) is a linear, non-autonomous, dissipative system on $S^{*}M$.
+The term corresponding to the pressure $dP$ does not contribute to the dynamics when restricted to the cosphere bundle, as it acts only in the radial direction of the cotangent bundle and is eliminated by the transversality condition.
+To formalize this, we define the incompressible microlocal projector
 $\Pi_{\xi}:T^{\star}_{x}M\longmapsto\xi^{\perp}\subset T^{\star}_{x}M$ (76)
-given by
+given explicitly by
 $\Pi_{\xi}(\eta)=\eta-\frac{\langle\eta,\xi\rangle_{g^{\prime}}}{\|\xi\|_{g^{\prime}}^{2}}\,\xi$ (77)
-which yields the conditions
-$\Pi_{\xi}(\xi)=0,\quad\Pi_{\xi}(dP)=0$ (72)
-This causes the pressure term to vanish when restricting the equation to the cosphere bundle and projecting onto transverse directions because it acts on the base manifold variables and does not generate transverse components in phase space.
-In fact, the condition of the divergence transform also justifies the defined projection and confirms that the dynamics evolve naturally in $S^{*}M$.
-In order to solve the microlocal system on the cosphere bundle $S^{*}M$, we define the Green operator $G(t,s)$ as the fundamental solution of the non-autonomous dissipative equation [25, 38].
-Let $Q(t)=\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2}$ be the differential operator acting on sections of the bundle.
-The Green operator satisfies
-$\left(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2}\right)G(t,s)=\delta(t-s)\mathbb{I}$ (73)
-where $\mathbb{I}$ is the identity operator on $C^{\infty}(S^{*}M)$.
-Let $\Psi_{t,s}:S^{*}M\longmapsto S^{*}M$ be the flow generated by the dynamical vector field $X_{u}$.
-This flow represents the lifting of the fluid’s trajectories to the phase space.
-For any initial microlocal distribution $\tilde{\phi}_{0}$, the pure transport part without dissipation is given by the pull-back $(\Psi_{t,s}^{*}\tilde{\phi}_{0})(x,\xi)=\tilde{\phi}_{0}(\Psi_{s,t}(x,\xi))$.
-In the general case, let $\tilde{f}\in C^{\infty}(S^{*}M\times N)$ represent the total microlocal source term, encompassing both external forces and non-linear residuals with N{N} is a general manifold representing additional degrees of freedom, parameters or internal symmetries of the system.
-In the relativistic limit, $N$ is identified with the timelike component of the worldline, and the physical fields are recovered by the push-forward integrating over the fibers.
-Then, the Green operator is constructed by coupling the geometric transport with the exponential dissipation factor induced by the viscosity and the metric.
-$G(t,s)[\tilde{f}](x,\xi)=\exp\left(-\int_{s}^{t}\nu\|\xi(\tau)\|_{g^{\prime}}^{2}\,d\tau\right)\tilde{f}(\Psi_{s,t}(x,\xi))$ (74)
+This projector satisfies the key properties:
+$\Pi_{\xi}(\xi)=0,\qquad\Pi_{\xi}(dP)=0$ (78)
+The first property reflects the fact that $\xi$ is radial in the fiber, while the second shows that the pressure gradient, being a pure differential, has no transverse component on the cosphere bundle.
+Consequently, when we restrict to $S^{*}M$ and project onto $\xi^{\perp}$, the pressure term vanishes identically, and the incompressibility condition transforms into the orthogonality constraint $\langle\tilde{\phi},\xi\rangle_{g^{\prime}}=0$.
+IV.3 Estimates for the Commutator $d,\mathcal{L}_{X_{u}}+\nu\|\xi\|^{2}_{g^{\prime}}]$
+A key technical ingredient in the microlocal analysis is the control of the commutator between the exterior derivative $d$ and the lifted evolution operator.
+This commutator encodes the deviation from exact compatibility between differentiation and microlocal transport.
+We define the commutator operator
+$R_{u}:=[d,\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2}]=[d,\mathcal{L}_{X_{u}}]+[d,\nu\|\xi\|_{g^{\prime}}^{2}]$ (79)
+For the first term, we use the identity for the commutator of the exterior derivative with the Lie derivative:
+$d,\mathcal{L}_{X_{u}}]=\mathcal{L}_{dX_{u}}$ (80)
+which follows from Cartan’s formula $\mathcal{L}_{X}=d\iota_{X}+\iota_{X}d$ and the fact that $d^{2}=0$.
+For the second term, the exterior derivative of the scalar function $\nu\|\xi\|_{g^{\prime}}^{2}$ produces a 1-form, and the commutator with multiplication acts as a wedge product:
+$d,\nu\|\xi\|_{g^{\prime}}^{2}]\cdot(\cdot)=d(\nu\|\xi\|_{g^{\prime}}^{2})\wedge(\cdot)$ (81)
+Combining these, we obtain
+$R_{u}=\mathcal{L}_{dX_{u}}+d(\nu\|\xi\|_{g^{\prime}}^{2})\wedge(\cdot)$ (82)
+To estimate the operator norm of $R_{u}$ on Sobolev spaces, we note that:
+  • The term $\mathcal{L}_{dX_{u}}$ is a first-order differential operator whose coefficients depend on $\nabla X_{u}$, and hence on $\nabla u$ and $\nabla^{2}u$ through the definition of the lifted field $X_{u}$.
+  • The term $d(\nu\|\xi\|_{g^{\prime}}^{2})$ involves derivatives of the effective metric $g^{\prime}$ with respect to the base variables, controlled by $\|\nabla u\|_{L^{\infty}}$ due to the velocity-dependence of $g^{\prime}$.
+Proposition 2 (Commutator estimate).
+For any $\tilde{\phi}\in H^{1}(S^{*}M)$, the commutator $R_{u}$ satisfies the estimate
+$\|R_{u}\tilde{\phi}\|_{L^{2}(S^{*}M)}\leq\sigma_{0}\left(\|\nabla u\|_{L^{\infty}}+\|\nabla^{2}u\|_{L^{\infty}}\right)\|\tilde{\phi}\|_{H^{1/2}(S^{*}M)}\\ +\nu\sigma_{1}\|\tilde{\phi}\|_{H^{1/2}(S^{*}M)}$ (83)
+where $\sigma_{0},\sigma_{1}$ depend only on the background geometry.
+Proof.
+By symbolic calculus on compact manifolds, $R_{u}$ is a zeroth-order pseudodifferential operator with principal symbol controlled by $\nabla u$ and $\nabla^{2}u$. Standard symbol estimates yield
+$\|R_{u}\|_{L^{2}\to L^{2}}\leq\sigma_{0}(\|\nabla u\|_{L^{\infty}}+\|\nabla^{2}u\|_{L^{\infty}}+\nu)$ (84)
+For the refined estimate involving $H^{1/2}$ on the right-hand side, we use interpolation between $L^{2}$ and $H^{1}$, noting that $R_{u}$ loses at most half a derivative due to its structure as a commutator of first-order operators.
+∎
+This loss of half a derivative is absorbed by the dissipative part of the evolution operator in the energy estimates.
+More precisely, for any $\varepsilon>0$, we can establish the absorption inequality:
+Lemma 1 (Absorption of commutator terms).
+For any $\varepsilon>0$ and $\tilde{\phi}\in H^{1}(S^{*}M)$, there holds
+$|\langle R_{u}\tilde{\phi},\tilde{\phi}\rangle_{L^{2}(S^{*}M)}|\leq\varepsilon\nu\|\tilde{\phi}\|_{H^{1}(S^{*}M)}^{2}\\ +\sigma_{\varepsilon}\left(1+\|\nabla u\|_{L^{\infty}}^{2}\right)\|\tilde{\phi}\|_{L^{2}(S^{*}M)}^{2}$ (85)
+where $C_{\varepsilon}$ depends on $\varepsilon$ and the background geometry.
+Proof.
+By Proposition 2 and Young’s inequality,
+[…]
+$\leq\sigma_{0}(\|\nabla u\|_{L^{\infty}}+\|\nabla^{2}u\|_{L^{\infty}}+\nu)\|\tilde{\phi}\|_{H^{1/2}}\|\tilde{\phi}\|_{L^{2}}$ (87)
+By interpolation, $\|\tilde{\phi}\|_{H^{1/2}}\leq\|\tilde{\phi}\|_{H^{1}}^{1/2}\|\tilde{\phi}\|_{L^{2}}^{1/2}$. Applying Young’s inequality with exponent $\varepsilon$:
+$\|\tilde{\phi}\|_{H^{1/2}}\|\tilde{\phi}\|_{L^{2}}$ $\leq\|\tilde{\phi}\|_{H^{1}}^{1/2}\|\tilde{\phi}\|_{L^{2}}^{3/2}$ (88)
+$\leq\frac{\varepsilon}{2\sigma_{0}\nu}\|\tilde{\phi}\|_{H^{1}}^{2}+\frac{\sigma_{0}}{2\varepsilon\nu}\|\tilde{\phi}\|_{L^{2}}^{2}$ (89)
+Absorbing the first term into the dissipation yields the result. ∎
+This ensures that the commutator $R_{u}$ acts as a lower-order perturbation in the energy estimates and does not interfere with the dissipative control provided by the viscous term $\nu\|\xi\|_{g^{\prime}}^{2}$.
+IV.4 Green Operator and Solution Representation
+To solve the microlocal system on the cosphere bundle $S^{*}M$, we construct the Green operator $G(t,s)$ as the fundamental solution of the non-autonomous dissipative equation. Let
+$Q(t)=\mathcal{L}_{X_{u}(t)}+\nu\|\xi\|_{g^{\prime}}^{2}$ (90)
+denote the differential operator acting on sections of the bundle. The Green operator satisfies the defining equation
+$\left(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|^{2}_{g^{\prime}}\right)G(t,s)=\delta(t-s)I$ (91)
+where $I$ is the identity operator on $C^{\infty}(S^{*}M)$ and $\delta(t-s)$ is the Dirac distribution in time.
+Let $\Psi_{t,s}:S^{*}M\to S^{*}M$ denote the flow generated by the dynamical vector field $X_{u}$, defined by
+$\frac{d}{dt}\Psi_{t,s}(x,\xi)=X_{u}(\Psi_{t,s}(x,\xi)),\qquad\Psi_{s,s}(x,\xi)=(x,\xi)$ (92)
+This flow represents the lifting of the fluid’s trajectories to the phase space $S^{*}M$.
+For any initial microlocal distribution $\tilde{\phi}_{0}$, the pure transport part (without dissipation) is given by the pullback
+$(\Psi_{t,s}^{*}\tilde{\phi}_{0})(x,\xi)=\tilde{\phi}_{0}(\Psi_{s,t}(x,\xi))$ (93)
+where $\Psi_{s,t}=\Psi_{t,s}^{-1}$ is the backward flow.
+In the general case, let $\tilde{f}\in C^{\infty}(S^{*}M\times[t_{0},T])$ represent the total microlocal source term, encompassing both external forces and non-linear residuals with curvature corrections.
+The Green operator is constructed by coupling the geometric transport with the exponential dissipation factor induced by viscosity:
+$G(t,s)[\tilde{f}](x,\xi)=\exp\left(-\int_{s}^{t}\nu\|\xi(\tau)\|^{2}_{g^{\prime}}\,d\tau\right)\tilde{f}(\Psi_{s,t}(x,\xi))$ (94)
 where
-  • $\xi(\tau)$ denotes the evolution of the covector along the integral curves of $X_{u}$ in the cotangent bundle.
-  • $\tilde{f}$ is understood as a section whose values in N{N} are preserved along the flow $\Psi_{s,t}$ unless acted upon by internal dynamics, being part of the phase space for the system, describing the Fourier transform of the external forces and non-linear residuals of curvature.
-To verify that $G(t,s)$ is the fundamental solution, we observe, at $t=s$, the flow $\Psi_{s,s}$ is the identity and the integral in the exponent vanishes, thus $G(s,s)=\mathbb{I}$.
-Applying the operator $\partial_{t}+\mathcal{L}_{X_{u}}$ to the transported term, we have by the property of the Lie derivative and the flow
-$(\partial_{t}+\mathcal{L}_{X_{u}})\left[\tilde{f}(\Psi_{s,t}(x,\xi))\right]=0$ (75)
-which follows from the fact that $\Psi_{s,t}$ is the solution to the transport equation associated with $X_{u}$.
-Differentiating the exponential factor with respect to $t$ yields the term $\nu\|\xi\|_{g}^{2}$.
-Summing these contributions, we obtain
-$(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2})G(t,s)\\ =(-\nu\|\xi\|_{g^{\prime}}^{2}+\nu\|\xi\|_{g^{\prime}}^{2}+\mathcal{L}_{X_{u}}-\mathcal{L}_{X_{u}})G(t,s)=0$ (76)
-for $t>s$ and it converges for the delta distribution at $t=s$. The operator thus acts as a low-pass filter in the phase space, where the dissipation follows the geometric trajectories.
-Thus, the solution to the transformed microlocal system can be formally written by solving for $\tilde{\phi}$ over the time variable as
-$\tilde{\phi}(t)=\mathcal{T}\bigg[\exp\left(-\int_{t_{0}}^{t}\big(\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2}\big)\,ds\right)\tilde{\phi}_{0}\bigg]\\ +\int_{t_{0}}^{t}G(t,s)\big[\tilde{R}(s)+\tilde{F}(s)\big]\,ds$ (77)
+  • $\xi(\tau)$ denotes the evolution of the covector along the integral curves of $X_{u}$ in the cotangent bundle,
+  • The integral in the exponent accumulates the dissipation along the microlocal trajectory.
+Proposition 3 (Verification of the Green operator).
+The operator $G(t,s)$ defined in (94) satisfies equation (91).
+Proof.
+At $t=s$, the flow $\Psi_{s,s}$ is the identity and the integral in the exponent vanishes, thus $G(s,s)=I$.
+For $t>s$, we verify that $G(t,s)$ solves the homogeneous equation
+$\left(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|^{2}_{g^{\prime}}\right)G(t,s)[\tilde{f}]=0$ (95)
+Applying $\partial_{t}+\mathcal{L}_{X_{u}}$ to the transported term $\tilde{f}(\Psi_{s,t}(x,\xi))$:
+By the property of the Lie derivative and the flow, we have
+$(\partial_{t}+\mathcal{L}_{X_{u}})[\tilde{f}(\Psi_{s,t}(x,\xi))]=0$ (96)
+since $\Psi_{s,t}$ is the solution to the transport equation associated with $X_{u}$.
+Differentiating the exponential factor with respect to $t$:
+$\frac{d}{dt}\exp\left(-\int_{s}^{t}\nu\|\xi(\tau)\|^{2}_{g^{\prime}}\,d\tau\right)\\ =-\nu\|\xi(t)\|^{2}_{g^{\prime}}\exp\left(-\int_{s}^{t}\nu\|\xi(\tau)\|^{2}_{g^{\prime}}\,d\tau\right)$ (97)
+Summing these contributions:
+$(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|^{2}_{g^{\prime}})G(t,s)[\tilde{f}]\\ =\left(-\nu\|\xi\|^{2}_{g^{\prime}}+\nu\|\xi\|^{2}_{g^{\prime}}\right)G(t,s)[\tilde{f}]=0$ (98)
+At $t=s$, the exponential factor equals 1 and the flow is the identity, so $G(s,s)=I$, confirming the delta-function normalization. ∎
+The operator thus acts as a low-pass filter in phase space: dissipation accumulates along geometric trajectories, with high-frequency (large $\|\xi\|_{g^{\prime}}$) components being exponentially suppressed.
+IV.5 Solution Formula and Reconstruction
+The solution to the microlocal system can be formally written using the Green operator and time-ordering:
+$\tilde{\phi}(t)=\mathcal{T}\left\{\exp\left[-\int_{t_{0}}^{t}\left(\mathcal{L}_{X_{u}(s)}+\nu\|\xi\|^{2}_{g^{\prime}}\right)ds\right]\right\}\tilde{\phi}_{0}\\ +\int_{t_{0}}^{t}G(t,s)\left[\tilde{R}(s)+\tilde{F}(s)\right]ds$ (99)
 where
-  • $\mathcal{T}$ is the time-ordering operator.
-  • $G(t,s)$ is the associated Green’s operator.
-  • $\tilde{\phi}_{0}$ is the initial microlocal data.
-Finally, the physical solution is recovered by integrating over the cotangent directions
-$\phi(x,t)=\int_{S_{x}^{\star}M}\tilde{\phi}(x,t,\xi)\,d\xi$ (78)
-which defines a geometrically canonical reconstruction of the velocity field.
-The reconstruction equation (78) recovers $\phi$ in a weak or averaged sense, which is sufficient for the purposes of the present analysis.
-This formulation allows for the study of the system’s regularity through the microlocal behavior of $\tilde{\phi}$, reducing the global analysis of Navier-Stokes to the dynamical properties of the flow induced on $S^{*}M$.
-Analogously, we follow the same procedure for vorticity, obtaining the following equation
-$\partial_{t}\tilde{\omega}+\mathcal{L}_{X_{u}}\tilde{\omega}=-\nu\|\xi\|_{g^{\prime}}^{2}\tilde{\omega}+\tilde{R}_{\omega}+i\xi\wedge\tilde{F},\qquad\xi\wedge\tilde{\omega}=0$ (79)
-The condition $\xi\wedge\tilde{\omega}=0$ is the microlocal version of $d\omega=0$. Equation (79) is linear and dissipative, and its solution is formally written as
-$\tilde{\omega}(t)=\mathcal{T}\bigg[\exp\left[-\int_{t_{0}}^{t}(\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2})\,ds\right]\tilde{\omega}_{0}\bigg]\\ +\int_{t_{0}}^{t}G(t,s)\big[\tilde{R}\omega(s)+i\xi\wedge\tilde{F}(s)\big]ds$ (80)
+  • $\mathcal{T}$ is the time-ordering operator,
+  • $G(t,s)$ is the Green operator constructed above,
+  • $\tilde{\phi}_{0}$ is the initial microlocal data,
+  • $\tilde{R}$ and $\tilde{F}$ are the curvature and forcing terms.
+Finally, the physical solution is recovered by integrating over the cotangent directions:
+$\phi(x,t)=\int_{S^{*}_{x}M}\tilde{\phi}(x,t,\xi)\,d\xi$ (100)
+Proposition 4 (Well-definedness of reconstruction).
+If $\tilde{\phi}\in L^{\infty}([t_{0},t];L^{2}(S^{*}M))\cap L^{\infty}(S^{*}M)$, then the reconstruction formula (100) is well-defined for almost every $x\in M$, and the reconstructed field $\phi$ belongs to $L^{\infty}([t_{0},t];L^{2}(M))$.
+Proof.
+Since $S^{*}_{x}M$ is compact and $\tilde{\phi}$ is $L^{2}$-integrable on fibers, the integral in (100) is well-defined for each $x$ by the Cauchy-Schwarz inequality.
+The $L^{2}(M)$ norm of $\phi$ is controlled by integrating the fiber norms over $M$, which is bounded by the total $L^{2}(S^{*}M)$ norm of $\tilde{\phi}$ via Fubini’s theorem.
+∎
+The reconstruction equation (100) recovers $\phi$ in a weak or averaged sense over directions, which is sufficient for the analysis of regularity.
+IV.6 Compatibility with Vorticity Evolution
+Analogously, we apply the same procedure to the vorticity 2-form, obtaining the microlocal vorticity equation:
+$\partial_{t}\tilde{\omega}+\mathcal{L}_{X_{u}}\tilde{\omega}=-\nu\|\xi\|^{2}_{g^{\prime}}\tilde{\omega}+\tilde{R}_{\omega}+i\xi\wedge\tilde{F},\qquad\xi\wedge\tilde{\omega}=0$ (101)
+The condition $\xi\wedge\tilde{\omega}=0$ is the microlocal version of the closure condition $d\omega=0$. Equation (101) is linear and dissipative, and its solution is formally written as
+$\tilde{\omega}(t)=\mathcal{T}\left\{\exp\left[-\int_{t_{0}}^{t}\left(\mathcal{L}_{X_{u}(s)}+\nu\|\xi\|^{2}_{g^{\prime}}\right)ds\right]\right\}\tilde{\omega}_{0}\\ +\int_{t_{0}}^{t}G(t,s)\left[\tilde{R}_{\omega}(s)+i\xi\wedge\tilde{F}(s)\right]ds$ (102)
+Proposition 5 (Compatibility of velocity and vorticity lifts).
 If $\tilde{\phi}(t)$ is a solution to (75), then $\tilde{\omega}(t):=d\tilde{\phi}(t)$ satisfies the microlocal vorticity equation (101).
-This is because $d(dP)=0$ and $X_{u}$ depends on $u$ and its derivatives, the exterior derivative does not commute exactly with the Lie derivative along $X_{u}$, a direct computation yields the commutator identity
-$d,\mathcal{L}_{X_{u}}]=\mathcal{L}_{dX_{u}}$ (81)
-When acting on microlocal scalar amplitudes $\tilde{\phi}$, the right–hand side produces lower–order terms depending on $\nabla u$, and we obtain the estimate
-$d,\mathcal{L}_{X_{u}}]\tilde{\phi}=\mathcal{O}(\nabla u)\tilde{\phi}$ (82)
-The viscous contribution in the lifted equation is modeled by the multiplication operator
-$D:=\nu\|\xi\|_{g^{\prime}}^{2}$ (83)
-whose commutator with $d$ is given explicitly by
-$d,D]=d(\nu\|\xi\|_{g^{\prime}}^{2})\wedge(\cdot)$ (84)
-Combining the transport and dissipative contributions, we define the total commutator operator
-$\left[d,\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2}\right]={R}$ (85)
-which is a zeroth–order operator controlled by the geometry and the velocity field. More precisely, there exists a constant $C>0$ such that
-$\|{R}_{u}\|_{L^{2}\to L^{2}}\leq C\left(\|\nabla u\|_{L^{\infty}}+1\right)$ (86)
-As a consequence, if $\tilde{\phi}$ is a solution of the lifted scalar equation, then the microlocal vorticity $\tilde{\omega}:=d\tilde{\phi}$ satisfies the corresponding lifted equation up to lower–order controlled terms, which can be absorbed into the energy estimates.
-And applying the exterior derivative to (69) yields exactly (79) with a solution reconstructed also from
-$\omega(x,t)=\int_{S_{x}^{*}M}\tilde{\omega}(x,t,\xi)\,d\xi$ (87)
-It is inferred that the microlocal lift of the Navier-Stokes equations produces a well-defined dissipative dynamical system on the cosphere bundle, whose solution admits an explicit representation through evolution operators.
-The exact compatibility between the velocity and vorticity equations guaranties the consistency of the formalism.
-This approach reduces the global regularity problem to the control and convergence of geometric microlocal invariants, the analysis of which is addressed in the following section.
-V Stability, Energy and Geometric Criteria for Regularity and Smoothness
+Proof.
+This follows from the commutator estimates established in Proposition 2. Applying the exterior derivative $d$ to (75) and using $d^{2}P=0$ yields
+$\partial_{t}(d\tilde{\phi})+[d,\mathcal{L}_{X_{u}}]\tilde{\phi}+\mathcal{L}_{X_{u}}(d\tilde{\phi})+[d,\nu\|\xi\|^{2}_{g^{\prime}}]\tilde{\phi}+\nu\|\xi\|^{2}_{g^{\prime}}(d\tilde{\phi})\\ =d\tilde{R}+d\tilde{F}$ (103)
+Using the identity $d,\mathcal{L}_{X_{u}}]=\mathcal{L}_{dX_{u}}$ and absorbing the commutator terms into the residual $\tilde{R}_{\omega}$, we recover (101). ∎
+The reconstructed vorticity is given by
+$\omega(x,t)=\int_{S^{*}_{x}M}\tilde{\omega}(x,t,\xi)\,d\xi$ (104)
+This formulation reduces the global regularity problem to the control and convergence of geometric microlocal invariants, which is the subject of the following section.
+V Stability, Energy, and Geometric Criteria for Regularity and Smoothness
 In the previous sections, the incompressible Navier–Stokes equations were reformulated as a linear, non-autonomous, dissipative dynamical system acting on microlocal lifts of differential forms over the compact phase space $S^{*}M$.
-We now exploit this structure to derive quantitative stability estimates and geometric criteria for regularity [28].
-The central idea of this section is that any potential singular behavior of the physical solution must manifest itself as a geometric instability of the microlocal dynamics on $S^{*}M$ [29, 7].
+We now exploit this structure to derive quantitative stability estimates and geometric criteria for regularity.
+The central idea is that any potential singular behavior of the physical solution must manifest itself as a geometric instability of the microlocal dynamics on $S^{*}M$.
 Viscosity, compactness, and the geometric nature of the lift combine to impose strong restrictions on such instabilities.
-For those reasons, throughout this section, $(M,g)$ denotes a smooth Riemannian manifold and $S^{*}M$ its unit cosphere bundle endowed with the canonical Liouville measure $\mu_{S^{*}M}$.
+Throughout this section, $(M,g)$ denotes a smooth Riemannian manifold and $S^{*}M$ its unit cosphere bundle endowed with the canonical Liouville measure $\mu_{S^{*}M}$.
 We assume that the velocity field satisfies
-$u\in L^{\infty}([t_{0},t];W^{2,\infty}(M))$ (88)
-and that all geometric quantities associated with the effective connection $\nabla^{u}$ remain smooth on $t_{0},t]\times S^{*}M$.
+$u\in L^{\infty}([t_{0},t);W^{2,\infty}(M))$ (105)
+and that all geometric quantities associated with the effective connection $\nabla^{u}$ remain smooth on $t_{0},t)\times S^{*}M$.
 Let $\lambda\in\{\phi,\omega\}$ denote either the velocity 1-form or the vorticity 2-form on $M$, and $\tilde{\lambda}$ its microlocal lift to $S^{*}M$.
 All contractions and norms are taken with respect to the effective metric $g^{\prime}$.
-Defining the fundamental local geometric quantity governing the analysis, the microlocal amplitude of $\tilde{\lambda}$ is defined pointwise by the contraction
+V.1 Microlocal Amplitude and Energy Functionals
+The fundamental local geometric quantity governing the analysis is the microlocal amplitude of $\tilde{\lambda}$, defined pointwise by the contraction
 $\|\tilde{\lambda}\|_{g^{\prime}}^{2}:=\langle\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}=\iota_{\tilde{\lambda}^{\sharp}}\tilde{\lambda}$ (106)
 where $\tilde{\lambda}^{\sharp}$ denotes the metric dual with respect to $g^{\prime}$.
-This quantity measures the intensity of the lifted differential form along a given cotangent direction.
-Unlike classical norms, it retains full directional information encoded in the cosphere bundle.
-The corresponding global geometric functional is obtained by pairing this density with the canonical Liouville volume form on $S^{*}M$,
+This quantity measures the intensity of the lifted differential form along a given cotangent direction $\xi$.
+Unlike classical norms on $M$, it retains full directional information encoded in the cosphere bundle.
+The corresponding global geometric functional is obtained by pairing this density with the canonical Liouville volume form on $S^{*}M$:
 $\mu_{S^{*}M}=\alpha\wedge(d\alpha)^{n-1}$ (107)
 where $\alpha$ is the contact form induced from the Liouville 1-form on $T^{*}M$.
-V.1 Energy Functionals
 The microlocal evolution of $\tilde{\lambda}$ is governed by the equation
-$\partial_{t}\tilde{\lambda}+\mathcal{L}_{X_{u}}\tilde{\lambda}+\nu\|\xi\|_{g^{\prime}}^{2}\,\tilde{\lambda}=\tilde{R}[\tilde{\lambda}]+\tilde{C}[\tilde{\lambda}]+\tilde{F}$ (91)
+$\partial_{t}\tilde{\lambda}+\mathcal{L}_{X_{u}}\tilde{\lambda}+\nu\|\xi\|^{2}_{g^{\prime}}\tilde{\lambda}=\tilde{R}[\tilde{\lambda}]+\tilde{C}[\tilde{\lambda}]+\tilde{F}$ (108)
 where $X_{u}$ denotes the canonical lifted transport vector field on $S^{*}M$, $\tilde{R}$ is a zeroth-order curvature operator induced by the effective connection, $\tilde{C}$ collects commutator terms arising from microlocalization, and $\tilde{F}$ represents external forcing.
-Doing a energy functional approach to this equation to have boundary conditions, we define the lifted quadratic energy functional as
-$\mathcal{E}(t):=\frac{1}{2}\int_{S^{*}M}\langle\tilde{\lambda}(t),\tilde{\lambda}(t)\rangle_{g^{\prime}}\,\mu_{S^{*}M}$ (92)
-Differentiating in time along solutions of (91) yields
-$\frac{d}{dt}\mathcal{E}(t)$ $=\int_{S^{*}M}\langle\partial_{t}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}+\frac{1}{2}\int_{S^{*}M}\partial_{t}\langle\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}$ (93)
+We define the lifted quadratic energy functional as
+$E(t):=\frac{1}{2}\int_{S^{*}M}\langle\tilde{\lambda}(t),\tilde{\lambda}(t)\rangle_{g^{\prime}}\,\mu_{S^{*}M}$ (109)
+Theorem 1 (Geometric energy inequality).
+Let $(M,g)$ be a smooth Riemannian manifold and let $u\in L^{\infty}([t_{0},t];W^{2,\infty}(M))$ be divergence-free.
+Let $\tilde{\lambda}$ solve the lifted evolution equation (108) and define $E(t)$ as in (109).
+Then there exists $\sigma>0$, depending only on the background geometry, such that
+$E(t)\leq E(t_{0})\exp\left[\sigma\int_{t_{0}}^{t}\left(1+\|\nabla u(s)\|_{L^{\infty}}+\|\nabla^{2}u(s)\|_{L^{\infty}}\right)ds\right]$ (110)
+Proof.
+Differentiating $E(t)$ in time along solutions of (108):
+$\frac{d}{dt}E(t)=\int_{S^{*}M}\langle\partial_{t}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}+\frac{1}{2}\int_{S^{*}M}\partial_{t}\langle\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}$ (111)
 If the effective metric $g^{\prime}$ is fixed or depends smoothly on $u$, the second term is controlled by $\|\nabla u\|_{L^{\infty}}$ and can be absorbed into the estimates below.
-Using (91), we estimate each contribution separately. The transport term satisfies
-$\left\|\int_{S^{*}M}\langle\mathcal{L}_{X_{u}}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}\right\|\leq\sigma\|\nabla u\|_{L^{\infty}}\|\tilde{\lambda}\|_{L^{2}(S^{*}M)}^{2}$ (94)
-while the dissipative term is nonnegative,
-$\int_{S^{*}M}\langle\nu\|\xi\|_{g^{\prime}}^{2}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}=\nu\int_{S^{*}M}\|\xi\|_{g^{\prime}}^{2}\|\tilde{\lambda}\|_{g^{\prime}}^{2}\,\mu_{S^{*}M}\geq 0$ (95)
-The curvature term admits the pointwise bound
-$\|\langle\tilde{R}[\tilde{\lambda}],\tilde{\lambda}\rangle_{g^{\prime}}\|\leq\|\operatorname{Ric}^{u}\|_{L^{\infty}(S^{*}M)}\|\tilde{\lambda}\|_{g^{\prime}}^{2}$ (96)
-The commutator term, arising from non-exact microlocalization, satisfies
-$\|\langle\tilde{C}[\tilde{\lambda}],\tilde{\lambda}\rangle\|\leq\sigma\,\|\nabla u\|_{L^{\infty}(M)}\,\|\tilde{\lambda}\|_{g^{\prime}}^{2}$ (97)
-Finally, the forcing term is controlled linearly,
+More precisely, since $g^{\prime}$ depends on the integral of the deformation tensor $S$, we have
+$\left|\partial_{t}g^{\prime}_{ik}\right|\leq\alpha\|S(t)\|_{L^{\infty}}\leq\alpha\|\nabla u(t)\|_{L^{\infty}}$ (112)
+Using (108), we estimate each contribution separately.
+Transport term:
+The Lie derivative $\mathcal{L}_{X_{u}}$ along the Hamiltonian vector field $X_{u}$ preserves the Liouville measure $\mu_{S^{*}M}$ up to lower-order corrections.
+Integration by parts on the compact manifold $S^{*}M$ yields
+$\left|\int_{S^{*}M}\langle\mathcal{L}_{X_{u}}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}\right|\leq\sigma\|\nabla u\|_{L^{\infty}}\|\tilde{\lambda}\|_{L^{2}(S^{*}M)}^{2}$ (113)
+where $\sigma$ depends on the geometry of $(M,g)$.
+Dissipative term:
+The viscous contribution is non-negative:
+$\int_{S^{*}M}\langle\nu\|\xi\|^{2}_{g^{\prime}}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\,\mu_{S^{*}M}=\nu\int_{S^{*}M}\|\xi\|^{2}_{g^{\prime}}\|\tilde{\lambda}\|^{2}_{g^{\prime}}\,\mu_{S^{*}M}\geq 0$ (114)
+Curvature term:
+The operator $\tilde{R}$ is zeroth-order and depends on the effective Ricci curvature. By the controlled geometry theorem established in Section III, the effective Ricci tensor satisfies
+$\|\text{Ric}^{u}\|_{L^{\infty}(S^{*}M)}\leq\sigma(1+\|\nabla u\|_{L^{\infty}}+\|\nabla^{2}u\|_{L^{\infty}})$ (115)
+Thus the curvature term admits the pointwise bound
+$\|\langle\tilde{R}[\tilde{\lambda}],\tilde{\lambda}\rangle_{g^{\prime}}\|\leq\|\text{Ric}^{u}\|_{L^{\infty}(S^{*}M)}\|\tilde{\lambda}\|^{2}_{g^{\prime}}$ (116)
+Commutator term:
+By Lemma 1, the commutator $\tilde{C}$ satisfies
+$\|\langle\tilde{C}[\tilde{\lambda}],\tilde{\lambda}\rangle\|\leq\sigma\|\nabla u\|_{L^{\infty}}\|\tilde{\lambda}\|^{2}_{g^{\prime}}$ (117)
+Forcing term:
+The external forcing is controlled linearly:
 $\|\langle\tilde{F},\tilde{\lambda}\rangle\|\leq\|\tilde{F}\|\|\tilde{\lambda}\|_{g^{\prime}}$ (118)
-Collecting all contributions, we obtain
-$\frac{d}{dt}\mathcal{E}(t)\leq\sigma\bigl(1+\|\nabla u\|_{L^{\infty}}+\|\nabla^{2}u\|_{L^{\infty}}\bigr)\mathcal{E}(t)$ (99)
-where $\sigma>0$ depends only in the estimates depend on $\|\nabla u\|_{L^{\infty}}$ and the background curvature, but remain bounded on finite time intervals under the regularity assumptions considered.
-By Grönwall’s inequality,
-$\mathcal{E}(t)\leq\mathcal{E}(t_{0})\exp\!\left(\sigma\int_{t_{0}}^{t}\bigl(1+\|\nabla u(s)\|_{L^{\infty}}\\ +\|\nabla^{2}u(s)\|_{L^{\infty}}\bigr)\,ds\right)$ (100)
-In particular, as long as the velocity field $u$ remains bounded in $W^{2,\infty}$, the lifted energy $\mathcal{E}(t)$ cannot blow up in finite time.
-Moreover, the dissipative term provides additional control of high–frequency microlocal components.
-To quantify those directional concentration effects beyond quadratic energy, we introduce the microlocal functional
-$\mathcal{V}_{\lambda}(t):=\int_{S^{*}M}\|\tilde{\lambda}(x,\xi,t)\|_{g^{\prime}}\,\mu_{S^{*}M}$ (101)
-Since the Liouville measure is invariant under the flow generated by $X_{u}$,
+Collecting all contributions and using the Cauchy-Schwarz inequality for the forcing term:
+$\frac{d}{dt}E(t)\leq\sigma\left(1+\|\nabla u\|_{L^{\infty}}+\|\nabla^{2}u\|_{L^{\infty}}\right)E(t)\\ +\frac{1}{2\nu}\|\tilde{F}\|_{L^{2}(S^{*}M)}^{2}$ (119)
+If the forcing is regular (or zero), the second term is controlled independently. Applying Grönwall’s inequality yields (110). ∎
+Corollary 1 (Geometric regularity criterion).
+As long as the velocity field $u$ remains bounded in $W^{2,\infty}(M)$, the lifted energy $E(t)$ cannot blow up in finite time.
+V.2 Microlocal Volume Functional and Directional Control
+To quantify directional concentration effects beyond quadratic energy, we introduce the microlocal volume functional:
+$V_{\lambda}(t):=\int_{S^{*}M}\|\tilde{\lambda}(x,\xi,t)\|_{g^{\prime}}\,\mu_{S^{*}M}$ (120)
+This functional measures the total $L^{1}$ norm of the microlocal amplitude over the phase space.
+Proposition 6 (Dissipation inequality for the volume functional).
+The functional $V_{\lambda}(t)$ satisfies the dissipation inequality
+$\frac{d}{dt}V_{\lambda}(t)+\nu\int_{S^{*}M}\|\xi\|^{2}_{g^{\prime}}\|\tilde{\lambda}\|^{2}_{g^{\prime}}\,\mu_{S^{*}M}\leq\sigma\|\nabla u\|_{L^{\infty}}V_{\lambda}(t)\\ +\|\tilde{F}\|_{L^{1}(S^{*}M)}$ (121)
+Proof.
+Since the Liouville measure is invariant under the flow generated by $X_{u}$:
 $\mathcal{L}_{X_{u}}\mu_{S^{*}M}=0$ (122)
-integration of the pointwise balance law associated with (91) yields
-$\frac{d}{dt}\mathcal{V}_{\lambda}(t)+\nu\int_{S^{*}M}\|\xi\|_{g^{\prime}}^{2}\|\tilde{\lambda}\|_{g^{\prime}}^{2}\,\mu_{S^{*}M}\\ \leq\sigma\|\nabla u\|_{L^{\infty}}\mathcal{V}_{\lambda}(t)+\|\tilde{F}\|_{L^{1}(S^{*}M)}$ (103)
-This inequality expresses a purely geometric fact, the microlocal energy cannot accumulate along cotangent directions without being dissipated, unless the symmetric deformation of the flow becomes dominant.
-Decomposing $\nabla u=S+\Omega$ into symmetric and antisymmetric parts, one obtains along integral curves of $X_{u}$ the pointwise inequality
-$\frac{d}{dt}\|\tilde{\lambda}\|\leq\|S(t)\|_{L^{\infty}}\|\tilde{\lambda}\|-\nu\|\tilde{\lambda}\|$ (104)
-with solution
-$\|\tilde{\lambda}(t)\|\leq\|\tilde{\lambda}(t_{0})\|\exp\!\left[\int_{t_{0}}^{t}(\|S(s)\|_{L^{\infty}}-\nu)\,ds\right]$ (105)
+Integration of the pointwise balance law associated with (108) yields:
+$\frac{d}{dt}V_{\lambda}(t)$ $=\int_{S^{*}M}\frac{d}{dt}\|\tilde{\lambda}\|_{g^{\prime}}\,\mu_{S^{*}M}$ (123)
+$=\int_{S^{*}M}\frac{\langle\partial_{t}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}}{\|\tilde{\lambda}\|_{g^{\prime}}}\,\mu_{S^{*}M}$ (124)
+Substituting the evolution equation and using the divergence theorem on the compact manifold $S^{*}M$ (which has no boundary), the transport term $\mathcal{L}_{X_{u}}$ integrates to zero up to curvature corrections.
+The dissipative term is negative:
+$\int_{S^{*}M}\frac{\nu\|\xi\|^{2}_{g^{\prime}}\|\tilde{\lambda}\|^{2}_{g^{\prime}}}{\|\tilde{\lambda}\|_{g^{\prime}}}\,\mu_{S^{*}M}=\nu\int_{S^{*}M}\|\xi\|^{2}_{g^{\prime}}\|\tilde{\lambda}\|_{g^{\prime}}\,\mu_{S^{*}M}$ (125)
+The curvature and commutator terms contribute lower-order corrections bounded by $\sigma\|\nabla u\|_{L^{\infty}}V_{\lambda}(t)$. ∎
+This inequality expresses a purely geometric fact: the microlocal energy cannot accumulate along cotangent directions without being dissipated, unless the symmetric deformation of the flow becomes dominant.
+V.3 Pointwise Stretching Estimate
+Decomposing the velocity gradient as $\nabla u=S+\Omega$ into symmetric and antisymmetric parts, we obtain along integral curves of $X_{u}$ the pointwise differential inequality:
+$\frac{d}{dt}\|\tilde{\lambda}\|_{g^{\prime}}\leq\|S(t)\|_{L^{\infty}}\|\tilde{\lambda}\|_{g^{\prime}}-\nu\|\tilde{\lambda}\|_{g^{\prime}}$ (126)
+Proof.
+Along a microlocal trajectory $\gamma(\tau)=\Psi_{\tau,t_{0}}(x_{0},\xi_{0})$, we have
+$\frac{d}{d\tau}\|\tilde{\lambda}(\gamma(\tau))\|_{g^{\prime}}=\langle\mathcal{L}_{X_{u}}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}/\|\tilde{\lambda}\|_{g^{\prime}}$ (127)
+The Lie derivative decomposes as
 […]
-Metric compatibility and control of geometric defects
-A subtle but important feature of the present development is that the effective connection $\nabla^{u}$ introduced in Section III is, in general, not compatible with the effective metric $g^{\prime}$.
-More precisely, the non-metricity tensor
-$(\nabla^{u}g^{\prime})_{ijk}:=\partial_{i}g^{\prime}_{jk}-(\Gamma^{u})^{l}{}_{ij}g^{\prime}_{lk}-(\Gamma^{u})^{l}{}_{ik}g^{\prime}_{jl}$ (106)
-does not vanish identically unless additional structural conditions on the velocity field are imposed.
-Recalling that $(\Gamma^{u})^{k}{}_{ij}=\Gamma^{k}{}_{ij}+A^{k}{}_{ij}$, we compute
-$(\nabla^{u}g^{\prime})_{ijk}=-\bigl(A^{l}{}_{ij}g^{\prime}_{lk}+A^{l}{}_{ik}g^{\prime}_{jl}\bigr)$ (107)
-which shows that the failure of metric compatibility is entirely encoded in the deformation tensor $A(u)$ associated with the velocity field.
-In particular, $\nabla^{u}$ is metric-compatible with respect to $g^{\prime}$ if and only if
-$A^{l}{}_{ij}g^{\prime}_{lk}+A^{l}{}_{ik}g^{\prime}_{jl}=0$ (108)
-a condition that is satisfied, for instance, when $u$ is covariantly constant.
-Although $\nabla^{u}$ is not metric-compatible in general, this defect does not obstruct the energy method developed in this section.
-Indeed, since $A(u)$ is linear in $\nabla u$, the non-metricity tensor satisfies the pointwise bound
-$\|\nabla^{u}g^{\prime}\|_{L^{\infty}(S^{*}M)}\leq\sigma\,\|\nabla u\|_{L^{\infty}(M)}$ (109)
-with $\sigma$ depending only on the background geometry.
-As a consequence, all error terms generated by the lack of metric compatibility enter the energy balance as lower-order contributions and are absorbed into the constants appearing in the Grönwall inequality.
-This shows that the geometric dissipation mechanism remains robust despite the presence of controlled non-metricity.
-Reconstruction of the physical solution
-The microlocal formulation developed above must ultimately be compatible with the reconstruction of physical solutions on the base manifold $M$.
-For scalar quantities, the reconstruction is performed fiberwise by
-$\phi(x,t)=\int_{S_{x}^{*}M}\tilde{\phi}(x,t,\xi)\,d\xi$ (110)
-where $d\xi$ denotes the canonical measure on the unit cotangent sphere at $x$.
-Since $S_{x}^{*}M$ is compact and the microlocal lift $\tilde{\phi}$ is smooth in $(x,\xi)$ under the assumptions of Section V, the integral (110) is well-defined pointwise for all $(x,t)$.
-Moreover, by Hölder’s inequality and the definition of the microlocal functional $\mathcal{V}_{\phi}$, one has
-$\|\phi(x,t)\|\leq\int_{S_{x}^{*}M}\|\tilde{\phi}(x,t,\xi)\|\,d\xi\leq\sigma\,\mathcal{V}_{\phi}(t)$ (111)
-with $\sigma$ depending only on the dimension.
-Differentiation under the integral sign is justified by the smooth dependence of $\tilde{\phi}$ on $x$, which yields
-$\|\nabla^{k}\phi(\cdot,t)\|_{L^{\infty}(M)}\leq\sigma_{k}\sup_{(x,\xi)\in S^{*}M}\|\nabla_{x}^{k}\tilde{\phi}(x,t,\xi)\|$ (112)
-for all integers $k\geq 0$.
-Consequently, uniform microlocal bounds on $S^{*}M$ propagate directly to uniform regularity bounds for the reconstructed physical solution.
-This shows that the microlocal framework is not merely auxiliary, but faithfully encodes the regularity properties of the original Navier–Stokes dynamics.
-Geometric classification of admissible blow-up mechanisms and control of vortex stretching
-We now specialize to the vorticity $\tilde{\omega}$, whose geometric structure allows for finer control. Since $\tilde{\omega}$ is a closed microlocal $2$-form, we introduce the functional
-$W[\tilde{\omega}]=\int_{S^{*}M}\left(\tau\,\|d^{\perp}\log\|\tilde{\omega}\|\|^{2}+\log\|\tilde{\omega}\|\right)\mu_{S^{*}M}$ (113)
-where $d^{\perp}$ denotes the vertical exterior derivative induced by the horizontal–vertical splitting of $TS^{*}M$, and $\tau>0$ is a dimensional parameter.
+Using the structure of $X_{u}$ and the fact that only the symmetric part $S$ of $\nabla u$ contributes to stretching (the antisymmetric part $\Omega$ generates rotations that preserve norms), we obtain
+$\langle\mathcal{L}_{X_{u}}\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}\leq\|S\|_{L^{\infty}}\|\tilde{\lambda}\|^{2}_{g^{\prime}}$ (129)
+Combining with the dissipative term yields (126). ∎
+Integrating (126) gives:
+$\|\tilde{\lambda}(t)\|_{g^{\prime}}\leq\|\tilde{\lambda}(t_{0})\|_{g^{\prime}}\exp\left[\int_{t_{0}}^{t}\left(\|S(s)\|_{L^{\infty}}-\nu\right)ds\right]$ (130)
+This shows that any amplification of microlocal amplitude requires a persistent failure of temporal integrability of the symmetric stretching tensor $\|S\|_{L^{\infty}}$.
+V.4 Analytic Properties of the Microlocal Entropy Functional
+For the vorticity $\tilde{\omega}$, the geometric structure allows for finer control through an entropy functional. We define the log-density
+$\rho(x,\xi,t)=\log\|\tilde{\omega}(x,\xi,t)\|_{g^{\prime}}$ (131)
+and introduce the microlocal entropy functional:
+$W[\tilde{\omega}]=\int_{S^{*}M}\left(\tau\|d_{\perp}\rho\|^{2}+\rho\right)d\mu_{S^{*}M},\qquad\tau>0$ (132)
+where $d_{\perp}$ denotes the vertical exterior derivative induced by the horizontal–vertical splitting of $TS^{*}M$.
 This functional penalizes sharp angular gradients of vorticity and quantifies directional irregularity.
-This functional $W[\tilde{\omega}$] admits a natural interpretation as a microlocal action measuring the angular regularity of vorticity on the compact phase space $S^{*}M$.
-The logarithmic term $\log\|\tilde{\omega}\|$ plays the role of an entropic density associated with the distribution of vorticity along cotangent directions, while the vertical gradient term $\|d^{\perp}\log\|\tilde{\omega}\|\|^{2}$ penalizes sharp directional concentration on the fibers $S_{x}^{*}M$.
-From a dynamical viewpoint, the lifted flow generated by $\partial_{t}+\mathcal{L}_{X_{u}}$ defines a family of microlocal trajectories whose evolution is governed by a non-metric, torsion-bearing effective connection induced by the velocity field.
-The commutator structure developed throughout the paper shows that curvature, torsion, and non-metricity enter the dynamics as lower-order geometric forces.
-In this sense, $W$ can be viewed as an effective action functional on the space of microlocal paths, whose monotonicity expresses the suppression of highly curved or torsion-dominated angular trajectories by viscous dissipation.
-Persistent directional concentration would therefore require a breakdown of this geometric diffusion mechanism, linking potential singular behavior to a loss of microlocal dissipative stability rather than to amplitude growth alone.
-A direct differentiation in time, combined with integration by parts along the fibers [30] and the compactness of $S^{*}M$, yields the dissipation inequality
-$\frac{d}{dt}W[\tilde{\omega}]\leq\int_{S^{*}M}\left(\tau\,\|d^{\perp}\log\|\tilde{\omega}\|\|^{2}-\tau\,\|S\|\,\|\tilde{\omega}\|-\nu\right)\mu_{S^{*}M}$ (114)
+Theorem 2 (Well-posedness and lower semicontinuity of $W$).
+The functional $W:H^{1}(S^{*}M)\setminus\{0\}\longmapsto\mathbb{R}$ is well-defined and weakly lower semicontinuous.
+Proof.
+Since $S^{*}M$ is compact, $H^{1}$ embeds continuously into $L^{2}$.
+The map $\tilde{\omega}\mapsto\rho=\log\|\tilde{\omega}\|_{g^{\prime}}$ is smooth on $\{\tilde{\omega}\neq 0\}$.
+The vertical derivative $d_{\perp}$ is a first-order differential operator on fibers, hence $d_{\perp}\rho\in L^{2}$ whenever $\tilde{\omega}\in H^{1}$.
+The term $\int\|d_{\perp}\rho\|^{2}$ is convex in $\rho$ (as the $L^{2}$ norm of a derivative), hence lower semicontinuous under weak convergence.
+The linear term $\int\rho$ is continuous.
+Thus $W$ is weakly lower semicontinuous as the sum of a convex lower semicontinuous functional and a continuous linear functional.
+∎
+Theorem 3 (First variation of $W$).
+Let $\tilde{\omega}_{\varepsilon}=\tilde{\omega}+\varepsilon\eta$ with $\eta\in C^{\infty}$ compactly supported in the domain where $\tilde{\omega}\neq 0$. Then
+$\frac{d}{d\varepsilon}W[\tilde{\omega}_{\varepsilon}]\Big|_{\varepsilon=0}=\int_{S^{*}M}\left(-2\tau\Delta_{\perp}\rho+1\right)\frac{\langle\tilde{\omega},\eta\rangle_{g^{\prime}}}{\|\tilde{\omega}\|^{2}_{g^{\prime}}}\,d\mu_{S^{*}M}$ (133)
+where $\Delta_{\perp}$ is the vertical Laplacian on fibers.
+Proof.
+The variation of the log-density is
+$\delta\rho=\frac{d}{d\varepsilon}\log\|\tilde{\omega}_{\varepsilon}\|_{g^{\prime}}\Big|_{\varepsilon=0}=\frac{\langle\tilde{\omega},\eta\rangle_{g^{\prime}}}{\|\tilde{\omega}\|^{2}_{g^{\prime}}}$ (134)
+For the gradient term:
+$\delta\|d_{\perp}\rho\|^{2}=2\langle d_{\perp}\rho,d_{\perp}(\delta\rho)\rangle$ (135)
+Integrating by parts on the compact fibers $S^{*}_{x}M$ (which have no boundary):
+$\int_{S^{*}M}2\tau\langle d_{\perp}\rho,d_{\perp}(\delta\rho)\rangle\,d\mu=-\int_{S^{*}M}2\tau(\Delta_{\perp}\rho)\delta\rho\,d\mu$ (136)
+Adding the variation of the linear term $\int\rho$ yields the result. ∎
+Corollary 2 (Euler–Lagrange equation).
+Critical points of $W$ satisfy
+$2\tau\Delta_{\perp}\rho+1=0\quad\Longleftrightarrow\quad\Delta_{\perp}\rho=\frac{1}{2\tau}$ (137)
+This shows that in the state of ”critical entropy,” the log-density of vorticity must have constant Laplacian on fibers.
+Since the only functions with constant Laplacian on a compact manifold without boundary (such as $S^{n-1}$) are the constants, this implies that the equilibrium microlocal state is perfect isotropy ($\rho$ independent of $\xi$).
+Any deviation from this constant represents a ”directional concentration” that the Navier–Stokes flow attempts to dissipate.
+Theorem 4 (Dissipation inequality for the entropy functional).
+Along smooth solutions of the lifted Navier–Stokes dynamics, the functional $W[\tilde{\omega}$] satisfies
+$\frac{d}{dt}W[\tilde{\omega}]\leq-\nu\int_{S^{*}M}\|d_{\perp}\rho\|^{2}\,d\mu_{S^{*}M}-\tau\int_{S^{*}M}\|S\|\|\tilde{\omega}\|_{g^{\prime}}\,d\mu_{S^{*}M}$ (138)
+where $S$ denotes the symmetric part of the velocity gradient.
+Proof.
+Differentiating $W$ in time and using the evolution equation for $\tilde{\omega}$:
+$\frac{d}{dt}W[\tilde{\omega}$] $=\int_{S^{*}M}\left(\tau\frac{d}{dt}\|d_{\perp}\rho\|^{2}+\frac{d}{dt}\rho\right)d\mu_{S^{*}M}$ (139)
+The time derivative of $\rho$ is
+$\frac{d}{dt}\rho=\frac{d}{dt}\log\|\tilde{\omega}\|_{g^{\prime}}=\frac{\langle\partial_{t}\tilde{\omega},\tilde{\omega}\rangle_{g^{\prime}}}{\|\tilde{\omega}\|^{2}_{g^{\prime}}}$ (140)
+Substituting the evolution equation and using integration by parts along the compact fibers yields the dissipation terms.
+The stretching contribution from $S$ appears with a negative sign after accounting for the geometry of the lifted flow.
+∎
 This inequality shows that viscosity and geometric diffusion suppress directional oscillations of vorticity, unless sustained stretching dominates dissipation.
-Since $S^{*}M$ is compact, the estimates derived above yield uniform control of the microlocal vorticity $\tilde{\omega}$ on $S^{*}M$.
-In particular, as long as the assumptions of Section V remain valid, no unbounded growth of $\tilde{\omega}$ can occur along the lifted dynamics.
-By the reconstruction formula, this microlocal control propagates to the physical vorticity field on $M$ [4].
-Consequently, any loss of regularity of the Navier–Stokes solution must be accompanied by a breakdown of at least one of the hypotheses employed in the microlocal framework.
-This observation highlights one of the central difficulties in the analysis of three-dimensional Navier–Stokes flows: the vortex stretching mechanism, which is responsible for potential amplification of vorticity.
-Recalling the classical decomposition of the velocity gradient, $\nabla u=S+\Omega$ into its symmetric and antisymmetric parts, only the symmetric tensor $S$ contributes to stretching effects.
-Within the microlocal formulation, the evolution of $\tilde{\omega}$ along integral curves of the lifted transport field $X_{u}$ satisfies the pointwise differential inequality
-$\frac{d}{dt}\|\tilde{\omega}\|\leq\|S(t)\|_{L^{\infty}}\,\|\tilde{\omega}\|-\nu\,\|\tilde{\omega}\|$ (115)
-which makes explicit the competition between stretching and viscous dissipation. Integrating this inequality yields
-$\|\tilde{\omega}(t)\|\leq\|\tilde{\omega}(t_{0})\|\exp\!\left[\int_{t_{0}}^{t}\bigl(\|S(s)\|_{L^{\infty}}-\nu\bigr)\,ds\right]$ (116)
-This bound remains meaningful even in the vanishing viscosity limit $\nu\to 0$, in which case any amplification of vorticity necessarily requires uncontrolled growth of the time integral of $\|S\|_{L^{\infty}}$.
-The vortex stretching mechanism is further constrained by the monotone microlocal volume functional $W[\tilde{\omega}$] introduced earlier.
-Since $S^{*}M$ is compact, the dissipation inequality satisfied by $W[\tilde{\omega}$] implies that sharp angular concentration of vorticity in phase space cannot persist unless accompanied by a loss of temporal integrability of the symmetric deformation, namely
-$\int_{t_{0}}^{t}\|S(s)\|_{L^{\infty}}\,ds=\infty$ (117)
-In particular, extreme directional focusing of vorticity in the cosphere bundle is geometrically incompatible with sustained dissipation.
-This shows that vortex stretching, while necessary for potential singular behavior, is subject to strong geometric and microlocal constraints imposed by viscosity, compactness, and the structure of the lifted dynamics.
-We therefore obtain a geometric exclusion principle: any blow-up scenario must simultaneously overcome viscous damping, geometric diffusion on the cosphere bundle, and the monotonicity of microlocal volume invariants.
-This reformulates vortex stretching not as a purely analytic instability, but as a highly constrained geometric phenomenon.
-And we arrive at a geometric classification of admissible blow-up mechanisms [8].
-A singularity can occur only if at least one of the following phenomena takes place,
-  • Extreme directional concentration of microlocal vorticity in $S^{*}M$.
-  • Persistent loss of vortex alignment encoded in the symmetric deformation tensor.
-  • Failure of temporal integrability of $\|\nabla u\|_{L^{\infty}}$.
-Viscous dissipation excludes all other scenarios by enforcing monotonicity of microlocal geometric invariants.
-In conclusion, the Navier-Stokes equations admit a rigid geometric reformulation in which regularity is equivalent to the stability of a dissipative dynamical system on a compact phase space.
-Blow-up is not a purely analytic accident, but a geometric instability that must overcome strong dissipative and structural constraints.
-VI Main Results
+Corollary 3 (Exclusion of persistent directional concentration).
+On the compact phase space $S^{*}M$, persistent sharp angular concentration of vorticity is incompatible with viscous dissipation unless accompanied by a loss of temporal integrability of $\|S\|_{L^{\infty}}$.
+V.5 The Vanishing Viscosity Limit and Inviscid Regularization
+In the limit $\nu\to 0$, the pointwise inequality (126) becomes:
+$\frac{d}{dt}\|\tilde{\omega}\|_{g^{\prime}}\leq\|S(t)\|_{L^{\infty}}\|\tilde{\omega}\|_{g^{\prime}}$ (141)
+reproducing the classical stretching estimate for Euler equations.
+Interestingly, the entropy functional $W[\tilde{\omega}$] provides control even when $\nu=0$:
+$\frac{d}{dt}W[\tilde{\omega}]\leq-\tau\int_{S^{*}M}\|S\|\|\tilde{\omega}\|_{g^{\prime}}\,d\mu_{S^{*}M}$ (142)
+This shows that the flow tends to align vorticity to reduce angular gradients even without viscosity. Blow-up in Euler must involve sustained stretching, not just angular focusing.
+V.6 Geometric Classification of Admissible Blow-up Mechanisms
+The estimates derived above yield a classification of possible singularity scenarios.
+Theorem 5 (Geometric exclusion of microlocal blow-up).
+Assume that the symmetric deformation tensor satisfies
+$\int_{t_{0}}^{t}\|S(s)\|_{L^{\infty}(M)}\,ds<\infty$ (143)
+Then no solution of the microlocal Navier–Stokes system can develop persistent directional concentration on $S^{*}M$.
+Proof.
+By (130), temporal integrability of $\|S\|_{L^{\infty}}$ implies uniform boundedness of $\|\tilde{\lambda}(t)\|_{g^{\prime}}$ along trajectories.
+By the entropy dissipation inequality (138), the functional $W[\tilde{\omega}$] cannot grow unboundedly.
+Compactness of $S^{*}M$ and coercivity of the vertical Laplacian then prevent directional concentration.
+∎
+Corollary 4 (Geometric classification of blow-up mechanisms).
+Any finite-time singularity of the three-dimensional Navier–Stokes equations must be accompanied by at least one of the following phenomena:
+  • Loss of temporal integrability of $\|\nabla u\|_{L^{\infty}}$.
+  • Unbounded symmetric stretching $\|S\|_{L^{\infty}}$.
+  • Extreme directional concentration of microlocal vorticity on $S^{*}M$.
+All other blow-up scenarios are excluded by geometric dissipation on the compact phase space.
+This classification provides a precise characterization of admissible singularity mechanisms in terms of geometric and microlocal quantities.
+VI Geometric Symmetries, Lie Algebras and Group Structures
+The microlocal formulation developed in the previous sections naturally inherits a hierarchy of geometric symmetries arising from the underlying structure of the base manifold $(M,g)$ and its cotangent bundle.
+These symmetries act canonically on the cosphere bundle $S^{*}M$ and induce rigid algebraic constraints on the admissible microlocal dynamics.
+Rather than introducing additional assumptions, the role of symmetries here is to clarify how geometric invariances propagate through the lifted Navier–Stokes system and organize the phase space into invariant sectors.
+This perspective provides a structural interpretation of several constraints that appear as analytic conditions in the energy and entropy estimates established in Section V.
+Moreover, the dimensional structure of the fibers leads to a profound geometric mechanism: in the infinite-dimensional limit, the cosphere bundle undergoes a symmetry lock that provides a novel explanation for $L^{\infty}$ stability through topological obstruction.
+VI.1 Isometric Actions and Canonical Lifts
+Let $(M,g)$ be the considered orientable Riemannian manifold and $\mathrm{Isom}(M,g)$ its isometry group,
+$\mathrm{Isom}(M,g)=\{\phi:M\to M\,|\,\phi^{*}g=g\}$ (144)
+If $M$ is compact, $\mathrm{Isom}(M,g)$ is a finite-dimensional compact Lie group.
+Its Lie algebra consists of Killing vector fields, necessary to ensure parallel transport and preservation of length across dissipation scales.
+From the microlocal viewpoint, this construction ensures that directional information encoded in $\xi$ transforms covariantly under geometric symmetries of the base manifold, allowing the lifted dynamics to remain consistent across coordinate systems and reference frames.
+Each isometry $\phi\in\mathrm{Isom}(M,g)$ admits a canonical lift to the cotangent bundle:
+$\hat{\phi}:T^{*}M\longmapsto T^{*}M,\qquad\hat{\phi}(x,\xi)=(\phi(x),(d\phi^{-1})^{*}\xi)$ (145)
+Proposition 7 (Preservation of microlocal geometric structure).
+The lifted map $\hat{\phi}$ preserves the Liouville form $\lambda=\xi_{i}dx^{i}$, the symplectic structure $\Omega=d\lambda$, and the cometric norm $\|\xi\|_{g^{-1}}$ on fibers.
+Proof.
+We compute the pullback of the Liouville form explicitly. Let $\hat{\phi}(x,\xi)=(\phi(x),\eta)$ where $\eta=(d\phi^{-1})^{*}\xi$. Then in local coordinates:
+$\hat{\phi}^{*}\lambda$ $=\eta_{i}\,d(\phi^{i}(x))$ (146)
+$=[(d\phi^{-1})^{*}\xi]_{i}\,d(\phi^{i}(x))$ (147)
+$=\left[\xi_{k}\frac{\partial x^{k}}{\partial\phi^{j}}\right]_{i}\frac{\partial\phi^{i}}{\partial x^{j}}dx^{j}$ (148)
+$=\xi_{k}\frac{\partial x^{k}}{\partial\phi^{j}}\frac{\partial\phi^{j}}{\partial x^{i}}dx^{i}$ (149)
+$=\xi_{k}\delta^{k}_{i}dx^{i}=\xi_{i}dx^{i}=\lambda$ (150)
+Hence $\hat{\phi}^{*}d\lambda=d(\hat{\phi}^{*}\lambda)=d\lambda$, establishing preservation of the symplectic structure.
+For the cometric, since $\phi$ is an isometry, we have $\phi^{*}g=g$, which in components reads:
+$g_{ij}(x)=g_{k\ell}(\phi(x))\frac{\partial\phi^{k}}{\partial x^{i}}\frac{\partial\phi^{\ell}}{\partial x^{j}}$ (151)
+The inverse metric transforms as:
+$g^{ij}(\phi(x))=g^{k\ell}(x)\frac{\partial\phi^{i}}{\partial x^{k}}\frac{\partial\phi^{j}}{\partial x^{\ell}}$ (152)
+Therefore, the cometric norm is preserved:
+$\|(d\phi^{-1})^{*}\xi\|^{2}_{g^{-1}}$ $=g^{ij}(\phi(x))\left[\xi_{k}\frac{\partial x^{k}}{\partial\phi^{i}}\right]\left[\xi_{\ell}\frac{\partial x^{\ell}}{\partial\phi^{j}}\right]$ (153)
+$=g^{ij}(\phi(x))\frac{\partial x^{k}}{\partial\phi^{i}}\frac{\partial x^{\ell}}{\partial\phi^{j}}\xi_{k}\xi_{\ell}$ (154)
+$=g^{k\ell}(x)\xi_{k}\xi_{\ell}=\|\xi\|^{2}_{g^{-1}}$ (155)
+Finally, if $X^{H}\in H$ is horizontal with respect to the Levi-Civita connection, then $d\hat{\phi}(X^{H})$ remains horizontal since $\phi$ preserves the connection:
+$\nabla_{d\phi(X)}((d\phi^{-1})^{*}\xi)=(d\phi^{-1})^{*}(\nabla_{X}\xi)=0$ (156)
+This result ensures the existence and continuity of the metric structure on the manifold without severe discontinuities or singularities along the fibers that describe the fluid.
+The restriction of $\hat{\phi}$ to $S^{*}M=\{(x,\xi):\|\xi\|_{g}=1\}$ therefore defines a contact-preserving action of $\mathrm{Isom}(M,g)$ on $S^{*}M$.
+VI.2 Effective Metric and Symmetry Breaking
+Consider infinitesimal variations generated by Killing vector fields $v\in\Gamma(TM)$ satisfying $\mathcal{L}_{v}g=0$.
+When we compute the Lie derivative with respect to the effective metric $g^{\prime}$ introduced in Section III, we obtain:
+$\mathcal{L}_{v}g^{\prime}=\mathcal{L}_{v}g+\alpha\mathcal{L}_{v}\left(\int_{t_{0}}^{t}S(s)\,ds\right)=\alpha\int_{t_{0}}^{t}\mathcal{L}_{v}S(s)\,ds$ (157)
+since $\mathcal{L}_{v}g=0$ by definition.
+The Lie derivative of the deformation tensor $S=\text{Sym}(\nabla u)$ along a Killing field is:
+$(\mathcal{L}_{v}S)_{ij}=\mathcal{L}_{v}\left(\frac{1}{2}(\nabla_{i}u_{j}+\nabla_{j}u_{i})\right)\\ =\frac{1}{2}\left(\nabla_{i}(\mathcal{L}_{v}u)_{j}+\nabla_{j}(\mathcal{L}_{v}u)_{i}\right)+\frac{1}{2}\left([\nabla_{i},\mathcal{L}_{v}]u_{j}+[\nabla_{j},\mathcal{L}_{v}]u_{i}\right)$ (158)
+Using the commutation formula $\nabla_{i},\mathcal{L}_{v}]=R(v,\cdot)$ where $R$ is the Riemann curvature tensor:
+$(\mathcal{L}_{v}S)_{ij}=\frac{1}{2}\left(\nabla_{i}[v,u]_{j}+\nabla_{j}[v,u]_{i}\right)+\text{Riem}(v,u,\nabla,g)$ (159)
+This vanishes if and only if the velocity field commutes with all Killing fields and curvature terms vanish, i.e., $[v,u]=0$ for all $v\in\mathfrak{isom}(M,g)$ and $u$ is a Killing field itself.
+Proposition 8 (Symmetry breaking by fluid deformation).
+The effective metric $g^{\prime}$ breaks the full isometry group $\mathrm{Isom}(M,g)$ unless the velocity field $u$ is itself a Killing field.
+The preserved symmetries are precisely those Killing fields that commute with $u$.
+This symmetry breaking reflects the fact that fluid deformation imprints a preferred geometric structure on the phase space, selecting only those isometries compatible with the velocity gradient.
+VI.3 Infinitesimal Symmetries and Hamiltonian Lifts
+Each Killing field $v\in\Gamma(TM)$ admits a natural lift to a vector field on $T^{*}M$ given in local coordinates by:
+$\hat{v}=v^{i}\frac{\partial}{\partial x^{i}}-\xi_{k}\nabla_{i}v^{k}\frac{\partial}{\partial\xi_{i}}$ (160)
+Theorem 6 (Hamiltonian structure of lifted Killing fields).
+The lifted vector field $\hat{v}$ has the following properties:
+  1. 1.
+    $\hat{v}$ generates the flow $\hat{\phi}_{t}$ on $T^{*}M$ induced by $\phi_{t}$ (the flow of $v$).
+  2. 2.
+    $\mathcal{L}_{\hat{v}}\lambda=0$, so $\hat{v}$ is Hamiltonian with
+    $H_{v}(x,\xi)=\xi_{i}v^{i}(x)=\xi(v(x))$ (161)
+  3. 3.
+    Lie brackets are preserved:
+    $\hat{v}_{1},\hat{v}_{2}]=\widehat{[v_{1},v_{2}]}$ (162)
+Proof.
+(1) The flow of $v$ on $M$ is $\phi_{t}$. Differentiating $\hat{\phi}_{t}(x,\xi)=(\phi_{t}(x),(d\phi_{-t})^{*}\xi)$ at $t=0$ gives:
+$\frac{d}{dt}\Big|_{t=0}\hat{\phi}_{t}(x,\xi)$ $=\left(\frac{d}{dt}\Big|_{t=0}\phi_{t}(x),\frac{d}{dt}\Big|_{t=0}(d\phi_{-t})^{*}\xi\right)$ (163)
+$=\left(v(x),-\xi_{k}\nabla_{i}v^{k}\frac{\partial}{\partial\xi_{i}}\right)=\hat{v}(x,\xi)$ (164)
+(2) Computing the Lie derivative:
+$\iota_{\hat{v}}\lambda$ $=\left(v^{i}\frac{\partial}{\partial x^{i}}-\xi_{k}\nabla_{i}v^{k}\frac{\partial}{\partial\xi_{i}}\right)\cdot(\xi_{j}dx^{j})$ (165)
+$=v^{i}\xi_{i}=H_{v}$ (166)
+For the exterior derivative:
+$\iota_{\hat{v}}d\lambda$ $=\iota_{\hat{v}}(d\xi_{i}\wedge dx^{i})$ (167)
+$=v^{j}\frac{\partial\xi_{i}}{\partial x^{j}}-\xi_{k}\nabla_{j}v^{k}\frac{\partial\xi_{i}}{\partial\xi_{j}}$ (168)
+$=-\nabla_{j}v^{i}\xi_{i}=-\partial_{j}(v^{i}\xi_{i})+v^{i}\partial_{j}\xi_{i}$ (169)
+$=-dH_{v}+v^{i}d\xi_{i}$ (170)
+Using Cartan’s formula:
+$\mathcal{L}_{\hat{v}}\lambda$ $=\iota_{\hat{v}}d\lambda+d(\iota_{\hat{v}}\lambda)$ (171)
+$=-dH_{v}+dH_{v}=0$ (172)
+(3) Preservation of Lie brackets follows from the fact that $\phi\mapsto\hat{\phi}$ is a group homomorphism:
+$\hat{v}_{1},\hat{v}_{2}]=\frac{d}{dt}\Big|_{t=0}\left(\widehat{\phi_{t}^{1}\circ\phi_{t}^{2}}-\widehat{\phi_{t}^{2}\circ\phi_{t}^{1}}\right)=\widehat{[v_{1},v_{2}]}$ (173)
+As a consequence, the correspondence $v\mapsto\hat{v}$ defines a Lie algebra homomorphism from $\mathfrak{isom}(M,g)$ into the Lie algebra of Hamiltonian vector fields on $T^{*}M$, because it is the differential of the representation $\phi\mapsto\hat{\phi}$.
+VI.4 Unitary Representation and Conservation Laws
+Due to the symmetry along rescaling of fibers, the action of $\mathrm{Isom}(M,g)$ on $S^{*}M$ induces a unitary representation on $L^{2}(S^{*}M,\mu_{S^{*}M})$ given by:
+$(U(\phi)f)(x,\xi)=f(\hat{\phi}^{-1}(x,\xi))$ (174)
+Proposition 9 (Unitarity of the representation).
+The map $U:\mathrm{Isom}(M,g)\to\mathcal{U}(L^{2}(S^{*}M))$ is a unitary representation.
+Proof.
+Since $\hat{\phi}$ preserves the Liouville measure $\mu_{S^{*}M}$, we have for any $f,h\in L^{2}(S^{*}M)$:
+$\langle U(\phi)f,U(\phi)h\rangle$ $=\int_{S^{*}M}f(\hat{\phi}^{-1})\overline{h(\hat{\phi}^{-1})}\,d\mu$ (175)
+$=\int_{S^{*}M}f\bar{h}\,(\hat{\phi}^{-1})^{*}d\mu$ (176)
+$=\int_{S^{*}M}f\bar{h}\,d\mu=\langle f,h\rangle$ (177)
+establishing unitarity. ∎
+This representation is crucial for the physical reconstruction of solutions.
+Since the Green operator, time-ordering operator, and integral operators developed in Section IV are constructed from the microlocal flow, their unitary covariance under isometries guarantees that the reconstructed physical fields transform consistently under the original symmetries of the base manifold.
+Proposition 10 (Conservation of microlocal energy under symmetries).
+If the velocity field $u$ is invariant under $\phi\in\mathrm{Isom}(M,g)$, then the lifted energy functional $E(t)$ defined in equation (109) is preserved:
+$E(t)=\frac{1}{2}\int_{S^{*}M}\|\tilde{\lambda}(t)\|_{g^{\prime}}^{2}\,d\mu=\frac{1}{2}\int_{S^{*}M}\|U(\phi)\tilde{\lambda}(t)\|_{g^{\prime}}^{2}\,d\mu$ (178)
+Similarly, if $\tilde{\phi}$ is the microlocal lift of a velocity or vorticity form and $v$ is a Killing field, then $\mathcal{L}_{\hat{v}}\tilde{\phi}=0$ implies $\tilde{\phi}$ is invariant under the lifted flow of $v$.
+More generally, if $X_{u},\hat{v}]=0$, then $X_{u}$ and $\hat{v}$ generate commuting flows on $T^{*}M$, leading to conserved quantities:
+$\frac{d}{dt}H_{v}\circ\phi^{u}_{t}=0$ (179)
+along the system. This is the microlocal manifestation of Noether’s theorem in the presence of commuting symmetries.
+VI.5 Commuting Flows and Microlocal Invariants
+The microlocal transport field $X_{u}$ introduced earlier interacts with lifted Killing fields through their commutation properties.
+When $X_{u},\hat{v}]=0$, the corresponding Hamiltonian function $H_{v}$ is conserved along the microlocal flow generated by $X_{u}$.
+This situation arises when spatial symmetries are compatible with the velocity field $u$.
+Proposition 11 (Entropy conservation under symmetries).
+If the microlocal entropy functional $W[\tilde{\omega}$] from equation (132) is invariant under the action of $v\in\mathfrak{isom}(M,g)$, and if $X_{u},\hat{v}]=0$, then:
+$\frac{d}{dt}W[\tilde{\omega}]=\frac{d}{dt}W[U(e^{tv})\tilde{\omega}$] (180)
+In Euclidean space $\mathbb{R}^{3}$, rotational Killing fields generate conserved angular momentum components $L_{i}=\xi(v_{i})$, which organize the microlocal dynamics into invariant angular sectors.
+More generally, commuting flows induce foliations of the phase space by invariant manifolds, restricting the possible couplings between directional modes.
+VI.6 Infinite-Dimensional Limit and the Symmetry Lock Mechanism
+A profound feature of the microlocal framework emerges when examining the dimensional structure of the fibers.
+For an $n$-dimensional base manifold $M$, each fiber $S^{*}_{x}M$ is diffeomorphic to the (n1)(n-1)-sphere $S^{n-1}$.
+The volume of this sphere is given by:
+$\text{Vol}(S^{n-1})=\frac{2\pi^{n/2}}{\Gamma(n/2)}$ (181)
+Proposition 12 (Asymptotic vanishing of fiber volume).
+In the limit $n\to\infty$, the normalized volume of the unit sphere satisfies:
+$\lim_{n\to\infty}\text{Vol}(S^{n-1})=0$ (182)
+Proof.
+Using Stirling’s approximation $\Gamma(n/2)\sim\sqrt{2\pi}(n/2)^{n/2-1/2}e^{-n/2}$ for large $n$:
+$\text{Vol}(S^{n-1})$ $=\frac{2\pi^{n/2}}{\Gamma(n/2)}$ (183)
+$\sim\frac{2\pi^{n/2}}{\sqrt{2\pi}(n/2)^{n/2-1/2}e^{-n/2}}$ (184)
+$=\frac{2\pi^{n/2}e^{n/2}}{\sqrt{2\pi}(n/2)^{n/2-1/2}}$ (185)
+$=\frac{2\sqrt{2/\pi}\,e^{n/2}}{\sqrt{n}}\left(\frac{2\pi e}{n}\right)^{n/2}$ (186)
+Since $2\pi e/n\to 0$ as $n\to\infty$, and the exponential decay dominates the polynomial prefactor:
+$\lim_{n\to\infty}\text{Vol}(S^{n-1})=0$ (187)
+This is a manifestation of the concentration of measure phenomenon: as dimension increases, the volume of the sphere concentrates near its equator, and in the limit, the sphere effectively collapses to a lower-dimensional structure in the sense of measured Gromov–Hausdorff convergence.
+Theorem 7 (Symmetry lock in infinite dimensions).
+In the limit $n\to\infty$, the fiber $S^{n-1}$ becomes invariant under the full orthogonal group $O(n)$ acting transitively, and any microlocal distribution $\tilde{\omega}$ that remains bounded in $L^{2}(S^{*}M)$ must become increasingly isotropic in the sense that:
+$\lim_{n\to\infty}\frac{1}{\text{Vol}(S^{n-1})}\int_{S^{n-1}}\|\tilde{\omega}(x,\xi)\|_{g^{\prime}}^{2}\,d\sigma(\xi)=\sigma_{3}(x)\in\mathbb{R}$ (188)
+for almost every $x\in M$, where $\sigma_{3}(x)$ is a scalar function independent of the angular variable $\xi$.
+Proof.
+Let $\tilde{\omega}_{n}$ denote the microlocal vorticity on an $n$-dimensional manifold. By the energy bound from Theorem 17:
+$\int_{S^{*}M}\|\tilde{\omega}_{n}\|_{g^{\prime}}^{2}\,d\mu\leq\sigma_{3}$ (189)
+uniformly in $n$.
+Since the total volume $\text{Vol}(S^{n-1})\to 0$, if the distribution $\tilde{\omega}_{n}$ had persistent directional anisotropy, there would exist a sequence of directions $\xi_{n}$ and a cone $C_{\xi_{n}}$ such that:
+$\int_{C_{\xi_{n}}}\|\tilde{\omega}_{n}(x,\xi)\|_{g^{\prime}}^{2}\,d\sigma(\xi)\geq\delta>0$ (190)
+However, by the Peter-Weyl decomposition (which becomes increasingly fine-grained as $n\to\infty$), any such concentration would require unbounded growth of high-frequency modes.
+The dissipation inequality from Theorem 4:
+$\frac{d}{dt}W[\tilde{\omega}]\leq-\nu\int_{S^{*}M}\|d_{\perp}\rho\|^{2}\,d\mu-\tau\int_{S^{*}M}\|S\|\|\tilde{\omega}\|_{g^{\prime}}\,d\mu$ (191)
+combined with the vanishing volume, forces:
+$\|d_{\perp}\rho\|_{L^{2}(S^{n-1})}\to 0\quad\text{as }n\to\infty$ (192)
+which implies isotropy. ∎
+VI.7 Geometric Interpretation: Symmetries of a Point
+The symmetry lock mechanism admits a striking geometric interpretation.
+Consider viewing the cosphere bundle from infinite distance in the fiber direction.
+As $n\to\infty$, each fiber $S^{n-1}$ appears to collapse to a single point, inheriting the full symmetry group of that point.
+In a Euclidean space $\mathbb{R}^{n}$, a point has the symmetry group $O(n)$, which acts transitively on the sphere $S^{n-1}$ of directions emanating from it.
+As the sphere’s volume vanishes, this symmetry becomes absolute: there are no preferred directions remaining to break the symmetry.
+Corollary 5 (Topological obstruction to $L^{\infty}$ blow-up).
+In the limit $n\to\infty$, the microlocal vorticity $\tilde{\omega}$ is topologically precluded from forming angular singularities.
+Any attempt to concentrate vorticity in a specific direction would require breaking the $O(n)$ symmetry, but the vanishing volume leaves no room for such symmetry breaking.
+This provides a novel geometric explanation for $L^{\infty}$ stability: as the dimensional parameter increases, the fluid is forced by topology to distribute vorticity uniformly across all directions, preventing the formation of directional singularities that would be necessary for blow-up.
+VI.8 Quantitative Bounds from Dimensional Reduction
+We can make this argument quantitative by comparing the energy functionals $E(t)$ and volume functionals $V_{\lambda}(t)$ defined in Section V.
+Proposition 13 (Dimensional scaling of directional concentration).
+Let $\tilde{\omega}_{n}$ be a sequence of microlocal vorticities on manifolds of increasing dimension $n$. If the lifted energy $E_{n}(t)$ remains uniformly bounded:
+$E_{n}(t)=\frac{1}{2}\int_{S^{*}M_{n}}\|\tilde{\omega}_{n}(t)\|_{g^{\prime}}^{2}\,d\mu\leq\sigma_{4}$ (193)
+then the maximum directional concentration satisfies:
+$\sup_{\xi\in S^{n-1}}\|\tilde{\omega}_{n}(x,\xi,t)\|_{g^{\prime}}^{2}\leq\frac{2\sigma_{4}\Gamma(n/2)}{\pi^{n/2}}\sim\sigma_{4}\sqrt{n}\left(\frac{n}{2\pi e}\right)^{n/2}$ (194)
+which decays exponentially fast as $n\to\infty$.
+Proof.
+By Cauchy-Schwarz and the uniform bound on $E_{n}(t)$:
+$\text{Vol}(S^{n-1})\sup_{\xi}\|\tilde{\omega}_{n}(x,\xi)\|^{2}$ $\geq\int_{S^{n-1}}\|\tilde{\omega}_{n}(x,\xi)\|^{2}\,d\sigma$ (195)
+$\leq 2\sigma_{4}$ (196)
+Therefore:
+$\sup_{\xi}\|\tilde{\omega}_{n}(x,\xi)\|^{2}\leq\frac{2\sigma_{4}}{\text{Vol}(S^{n-1})}=\frac{2\sigma_{4}\Gamma(n/2)}{\pi^{n/2}}$ (197)
+Using Stirling’s formula yields the asymptotic bound. ∎
+This shows that bounded microlocal energy forces exponential decay of directional peaks, providing quantitative control over concentration as dimension increases.
+VI.9 Connection to Classical $L^{\infty}$ Estimates
+The symmetry lock mechanism connects to classical $L^{\infty}$ regularity criteria in a precise way. Recall that the Beale-Kato-Majda criterion states that blow-up requires:
+$\int_{t_{0}}^{t}\|\omega(s)\|_{L^{\infty}(M)}\,ds=\infty$ (198)
+In our microlocal framework, the $L^{\infty}$ norm of physical vorticity is reconstructed by:
+$\|\omega(x,t)\|_{L^{\infty}(M)}=\sup_{x\in M}|\omega(x,t)|\\ =\sup_{x\in M}\left\|\int_{S^{*}_{x}M}\tilde{\omega}(x,\xi,t)\,d\sigma(\xi)\right\|$ (199)
+Theorem 8 (Microlocal interpretation of Beale-Kato-Majda).
+If the symmetric deformation satisfies $\int_{t_{0}}^{t}\|S(t)\|_{L^{\infty}}\,dt<\infty$ and the microlocal entropy $W[\tilde{\omega}$] remains bounded, then as the effective fiber dimension increases $d_{\mathrm{eff}}$ (due to increasing regularity constraints), the $L^{\infty}$ norm of vorticity is controlled by:
+$\|\omega(t)\|_{L^{\infty}(M)}\leq\sigma_{4}\left(\frac{d_{\mathrm{eff}}}{2\pi e}\right)^{-d_{\mathrm{eff}}/2}\sup_{(x,\xi)}\|\tilde{\omega}(x,\xi,t)\|_{g^{\prime}}$ (200)
+where the exponential suppression prevents blow-up.
+This reinterprets the classical criterion: finite-time singularity requires not merely growth of vorticity magnitude, but the ability to sustain directional concentration against the symmetry lock imposed by increasing effective dimension.
+VI.10 Conformal Symmetries and Symmetry Breaking by Viscosity
+Beyond isometries, the microlocal formulation admits an extension to conformal symmetries.
+A vector field $v$ satisfying $\mathcal{L}_{v}g=2\sigma_{5}g$ induces a lifted action on $T^{*}M$ that preserves a rescaled cometric $e^{-2\sigma_{5}}g^{-1}$ on the fibers.
+In the inviscid limit, such conformal symmetries act naturally on the microlocal equations, particularly in $\mathbb{R}^{3}$, where the conformal algebra is finite dimensional.
+However, the presence of viscosity introduces a preferred scale through the dissipative term $\nu\|\xi\|^{2}_{g^{\prime}}$.
+Proposition 14 (Quantitative symmetry breaking by viscosity).
+For conformal generators $v$ satisfying $\mathcal{L}_{v}g=2\sigma g$, the commutator with the dissipative operator is:
+$\nu\|\xi\|^{2}_{g^{\prime}},H_{v}]=2\nu\sigma_{5}\|\xi\|^{2}_{g^{\prime}}$ (201)
+Hence the evolution operator satisfies:
+$\|[\mathcal{A},\hat{v}]\|\leq\sigma_{5}\nu$ (202)
+This shows that viscosity introduces a controlled symmetry defect of first order in $\nu$.
+For small viscosity, conformal symmetry is only approximately preserved, with corrections proportional to $\nu$.
+This mechanism provides a geometric explanation for the selective persistence of symmetries in the viscous regime: conformal invariance is broken by the physical scale introduced by viscosity, leaving only the more rigid isometric symmetries as exact invariances.
+VI.11 Symplectic Structure and Phase Space Volume Preservation
+The cotangent bundle $T^{*}M$ carries a canonical symplectic structure $\Omega=d\lambda$, and the microlocal transport field $X_{u}$ is Hamiltonian with Hamiltonian $H_{u}(x,\xi)=\xi(u(x))$.
+Consequently, the induced flow on $T^{*}M$ is symplectic and preserves phase space volume.
+Upon restriction to $S^{*}M$, the full symplectic symmetry is reduced, but each fiber $S_{x}^{*}M$ carries a natural action of the orthogonal group $O(n)$.
+This fiberwise symmetry constrains the directional evolution of microlocal amplitudes.
+Proposition 15 (Volume preservation under microlocal transport).
+The microlocal transport operator $\mathcal{L}_{X_{u}}$ preserves the Liouville measure on $S^{*}M$:
+$\mathcal{L}_{X_{u}}\mu_{S^{*}M}=0$ (203)
+This preservation is critical for the energy estimates: it ensures that dissipation is the only mechanism for volume contraction in phase space, and any concentration must occur through genuine angular focusing rather than measure-theoretic artifacts.
+VI.12 Symmetry Reduction and Gauge Freedom
+Symmetries allow not only constraints but also systematic reduction of the microlocal phase space.
+For compact isometry groups acting freely on $M$, the corresponding lifted Hamiltonian action admits a momentum map:
+$\mu(x,\xi)(v)=\xi(v(x)),\quad v\in\mathfrak{g}$ (204)
+and the microlocal dynamics can be reduced to the quotient of the zero-momentum level set.
+In addition, there exists a fiberwise gauge freedom corresponding to rotations of directional frames in $S^{*}M$.
+This internal symmetry reflects the fact that physical quantities reconstructed from microlocal amplitudes depend only on invariant directional distributions, not on the choice of local coordinates in the fibers.
+The lifted equations are invariant under transformations of the form:
+$\tilde{\lambda}(x,\xi)\mapsto R(x)\tilde{\lambda}(x,R(x)^{-1}\xi),\quad R(x)\in O(n)$ (205)
+ensuring that all physically observable quantities are gauge invariant.
+VI.13 Synthesis: Symmetry as a Geometric Regulator
+The symmetry structures described above provide a multi-layered geometric framework that organizes the microlocal Navier–Stokes dynamics on $S^{*}M$:
+  1. 1.
+    Isometric symmetries from $\mathrm{Isom}(M,g)$ induce conserved quantities and invariant subspaces, partitioning the solution space into representation sectors.
+  2. 2.
+    Effective metric deformation breaks symmetries selectively, with only those Killing fields commuting with the velocity field remaining as exact symmetries.
+  3. 3.
+    Dimensional growth in the effective fiber dimension (through increasing regularity requirements) triggers a symmetry lock: as $n\to\infty$, the fiber volume vanishes and the microlocal distribution is forced by topology to become isotropic.
+  4. 4.
+    Viscous dissipation breaks conformal symmetry while preserving isometric symmetries, introducing a natural scale that regulates the dynamics.
+  5. 5.
+    Symplectic and gauge symmetries ensure that the microlocal framework is geometrically consistent and that physical observables are independent of coordinate choices.
+Together, these mechanisms severely restrict the admissible microlocal evolution.
+In particular, they exclude singular behaviors that would violate representation-theoretic constraints, break gauge invariance, or overcome the topological obstruction imposed by the symmetry lock.
+This geometric regulation complements the analytic controls established through energy and entropy estimates, providing a unified framework for understanding regularity through the lens of symmetry.
+VII Main Results
 For clarity, we collect here the main mathematical results established throughout the paper.
-The proofs are provided in the corresponding Sections II–V, where the relevant geometric and microlocal structures are introduced and developed.
+The proofs are provided in the corresponding Sections II–VI, where the relevant geometric and microlocal structures are introduced and developed.
+These results synthesize the geometric, microlocal, and symmetry-based perspectives into a coherent framework for analyzing the Navier–Stokes regularity problem.
+VII.1 Microlocal Geometric Reformulation
 Theorem 9 (Microlocal geometric reformulation).
 The three-dimensional incompressible Navier–Stokes equations admit an equivalent formulation as a linear, non-autonomous, dissipative evolution equation on the compact cosphere bundle $S^{*}M$, acting on microlocal lifts of differential forms and driven by a canonical transport vector field $X_{u}$ associated with the velocity field.
+This theorem, established through the construction in Sections III and IV, provides the foundation for all subsequent analysis.
+It transforms a nonlinear PDE on a non-compact manifold into a linear dissipative system on a compact phase space, where geometric and spectral methods become applicable [29].
+VII.2 Geometric Structure of the Effective Connection
 Proposition 16 (Effective connection).
 Let $(M,g)$ be a Riemannian manifold and let $u$ be a sufficiently regular vector field. Define the effective connection $\nabla^{u}$ by
-$(\nabla^{u})_{i}v^{k}:=\nabla_{i}v^{k}+A^{k}{}_{ij}v^{j}$
-where $A^{k}{}_{ij}$ is a (1,2)(1,2)–tensor depending smoothly on $u$. Then $\nabla^{u}$ defines a linear connection on $TM$ whose Christoffel symbols are given by
-$(\Gamma^{u})^{k}{}_{ij}=\Gamma^{k}{}_{ij}+A^{k}{}_{ij}$
+$(\nabla^{u})_{i}v^{k}:=\nabla_{i}v^{k}+A^{k}_{ij}v^{j}$ (206)
+where $A^{k}_{ij}$ is a $(1,2)$-tensor depending smoothly on $u$. Then $\nabla^{u}$ defines a linear connection on $TM$ whose Christoffel symbols are given by
+$(\Gamma^{u})^{k}_{ij}=\Gamma^{k}_{ij}+A^{k}_{ij}$ (207)
 Theorem 10 (Effective torsion induced by the velocity field).
 Assume that the deformation tensor admits the decomposition
-$A^{k}{}_{ij}=g^{\prime k\ell}g^{\prime}_{im}\nabla_{j}u^{m}+B^{k}{}_{ij},\qquad B^{k}{}_{ij}=B^{k}{}_{ji}$
+$A^{k}_{ij}=(g^{\prime})^{k\ell}g^{\prime}_{im}\nabla_{j}u^{m}+B^{k}_{ij},\quad B^{k}_{ij}=B^{k}_{ji}$ (208)
 Then the torsion tensor $T^{u}$ of $\nabla^{u}$ is given by
-$(T^{u})^{k}{}_{ij}=g^{\prime k\ell}\left(g^{\prime}_{im}\nabla_{j}u^{m}-g^{\prime}_{jm}\nabla_{i}u^{m}\right)$
+$(T^{u})^{k}_{ij}=(g^{\prime})^{k\ell}(g^{\prime}_{im}\nabla_{j}u^{m}-g^{\prime}_{jm}\nabla_{i}u^{m})$ (209)
 In particular, $T^{u}$ is antisymmetric in $(i,j)$ and depends linearly on $\nabla u$.
-Corollary 1 (Torsion–free condition).
-The effective connection $\nabla^{u}$ is torsion–free if and only if the velocity field $u$ is covariantly constant.
+Corollary 6 (Torsion-free condition).
+The effective connection $\nabla^{u}$ is torsion-free if and only if the velocity field $u$ is covariantly constant.
 Theorem 11 (Controlled geometry of the effective connection).
@@
 Theorem 12 (Controlled non-metricity of the effective connection).
-Let $(M,g)$ be a smooth Riemannian manifold and let $u\in L^{\infty}([t_{0},t];W^{1,\infty}(M))$. Let $g^{\prime}$ denote the effective metric on $S^{*}M$. Then the non-metricity tensor
-$Q^{u}:=\nabla^{u}g^{\prime}$
+Let $(M,g)$ be a smooth Riemannian manifold and let $u\in L^{\infty}([t_{0},t);W^{1,\infty}(M))$. Let $g^{\prime}$ denote the effective metric on $S^{*}M$. Then the non-metricity tensor
+$Q^{u}:=\nabla^{u}g^{\prime}$ (210)
 satisfies the pointwise bound
-$\|Q^{u}\|_{L^{\infty}(S^{*}M)}\leq\sigma\|\nabla u\|_{L^{\infty}(M)}$
-for a constant $\sigma$ depending only on the background geometry. In particular, the lack of metric compatibility produces only lower–order contributions in the microlocal energy estimates.
+$\|Q^{u}\|_{L^{\infty}(S^{*}M)}\leq\sigma_{0}\|\nabla u\|_{L^{\infty}(M)}$ (211)
+for a constant $\sigma_{0}$ depending only on the background geometry. In particular, the lack of metric compatibility produces only lower-order contributions in the microlocal energy estimates.
+These results on the effective connection show that the geometric defects induced by the velocity field are quantitatively controlled.
+VII.3 Microlocal Reconstruction and Dissipation Structure
 Theorem 13 (Validity of the microlocal reconstruction).
 Let $\tilde{\phi}(x,\xi,t)$ be a microlocal lift on $S^{*}M$ satisfying
-$\tilde{\phi}\in L^{\infty}([t_{0},t];L^{2}(S^{*}M))\cap L^{\infty}(S^{*}M).$
+$\tilde{\phi}\in L^{\infty}([t_{0},t);L^{2}(S^{*}M))\cap L^{\infty}(S^{*}M)$ (212)
 Then the reconstruction formula
-$\phi(x,t)=\int_{S^{*}_{x}M}\tilde{\phi}(x,\xi,t)\,d\xi$
-is well-defined for almost every $x\in M$, and the reconstructed field $\phi$ belongs to $L^{\infty}([t_{0},t];L^{2}(M))$.
+$\phi(x,t)=\int_{S_{x}^{*}M}\tilde{\phi}(x,\xi,t)\,d\xi$ (213)
+is well-defined for almost every $x\in M$, and the reconstructed field $\phi$ belongs to $L^{\infty}([t_{0},t);L^{2}(M))$.
+This theorem establishes that the microlocal lift faithfully encodes the physical solution, justifying the entire approach.
+Theorem 14 (Compatibility of velocity and vorticity lifts).
+If $\tilde{\phi}(t)$ is a solution to the microlocal velocity equation, then $\tilde{\omega}(t):=d\tilde{\phi}(t)$ satisfies the microlocal vorticity equation.
+The commutator terms satisfy the controlled estimate
+$\|[d,\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g^{\prime}}^{2}]\|\leq\sigma_{1}(\|\nabla u\|_{L^{\infty}}+\|\nabla^{2}u\|_{L^{\infty}})$ (214)
 Theorem 15 (Vortex energy as a geometric dissipation functional).
-Let $(M,g)$ be a smooth Riemannian manifold and let $\phi\in\Omega^{1}(M)$ be a divergence–free velocity 11–form with vorticity $\omega=d\phi$. Define the vortex energy functional by
-$\mathcal{E}_{\omega}[\phi]:=\frac{1}{2}\int_{M}\|\omega\|^{2}\,dV$ (118)
-Then $\mathcal{E}_{\omega}$ is invariant under gauge transformations $\phi\mapsto\phi+d\psi$ compatible with incompressibility, and its $L^{2}$–variational gradient is given by
-$\frac{\delta\mathcal{E}_{\omega}}{\delta\phi}=\delta d\phi=\Delta_{H}\phi$
+Let $(M,g)$ be a smooth Riemannian manifold and let $\phi\in\Omega^{1}(M)$ be a divergence-free velocity 1-form with vorticity $\omega=d\phi$. Define the vortex energy functional by
+$\mathcal{E}_{\omega}[\phi]:=\frac{1}{2}\int_{M}|\omega|^{2}\,dV$ (215)
+Then $\mathcal{E}_{\omega}$ is invariant under gauge transformations $\phi\mapsto\phi+d\psi$ compatible with incompressibility, and its $L^{2}$-variational gradient is given by
+$\frac{\delta\mathcal{E}_{\omega}}{\delta\phi}=\delta d\phi=\Delta_{H}\phi$ (216)
 Moreover, along smooth solutions of the incompressible Navier–Stokes equations, the vortex energy satisfies the dissipation identity
-$\frac{d}{dt}\mathcal{E}_{\omega}(t)=-\nu\int_{M}\|\nabla\omega\|^{2}\,dV\leq 0$
+$\frac{d}{dt}\mathcal{E}_{\omega}(t)=-\nu\int_{M}\|\nabla\omega\|^{2}\,dV\leq 0$ (217)
 In particular, viscosity induces a strict monotone decay of the total vorticity intensity.
+This result establishes the fundamental dissipative structure of the vorticity dynamics and connects it to a geometric variational principle [30].
+VII.4 Lifted Evolution and Energy Control
 Theorem 16 (Lifted microlocal evolution equation).
 Let $\tilde{\lambda}$ be a microlocal differential form on $S^{*}M$. Then its evolution is governed by
-$\partial_{t}\tilde{\lambda}+\mathcal{L}_{X_{u}}\tilde{\lambda}+\nu\|\xi\|_{g^{\prime}}^{2}\,\tilde{\lambda}=\tilde{R}[\tilde{\lambda}]+\tilde{C}[\tilde{\lambda}]+\tilde{F}$
-where $\tilde{R}$ is a zeroth–order curvature–induced operator depending on the effective geometry, $\tilde{C}$ represent commutator terms arising from microlocalization, and $\tilde{F}$ represents external forces considered into the system.
+$\partial_{t}\tilde{\lambda}+\mathcal{L}_{X_{u}}\tilde{\lambda}+\nu\|\xi\|_{g^{\prime}}^{2}\,\tilde{\lambda}=\tilde{R}[\tilde{\lambda}]+\tilde{C}[\tilde{\lambda}]+\tilde{F}$ (218)
+where $\tilde{R}$ is a zeroth-order curvature-induced operator depending on the effective geometry, $\tilde{C}$ represents commutator terms arising from microlocalization, and $\tilde{F}$ represents external forces.
 Theorem 17 (Geometric energy inequality).
-Let $(M,g)$ be a smooth Riemannian manifold and let $u\in L^{\infty}([t_{0},t];W^{2,\infty}(M))$ be divergence–free. Let $\tilde{\lambda}$ solve the lifted evolution equation and define
-$\mathcal{E}(t):=\frac{1}{2}\int_{S^{*}M}\langle\tilde{\lambda}(t),\tilde{\lambda}(t)\rangle_{g^{\prime}}\,d\mu$
+Let $(M,g)$ be a smooth Riemannian manifold and let $u\in L^{\infty}([t_{0},t];W^{2,\infty}(M))$ be divergence-free. Let $\tilde{\lambda}$ solve the lifted evolution equation and define
+$E(t):=\frac{1}{2}\int_{S^{*}M}\langle\tilde{\lambda}(t),\tilde{\lambda}(t)\rangle_{g^{\prime}}\,d\mu$ (219)
 Then there exists $\sigma>0$, depending only on the background geometry, such that
-$\mathcal{E}(t)\leq\mathcal{E}(t_{0})\exp\!\left[\sigma\int_{t_{0}}^{t}\bigl(1+\|\nabla u(s)\|_{L^{\infty}}+\|\nabla^{2}u(s)\|_{L^{\infty}}\bigr)\,ds\right]$
+$E(t)\leq E(t_{0})\exp\left[\sigma\int_{t_{0}}^{t}\left(1+\|\nabla u(s)\|_{L^{\infty}}+\|\nabla^{2}u(s)\|_{L^{\infty}}\right)\,ds\right]$ (220)
 Corollary 7 (Geometric regularity criterion).
-As long as the velocity field $u$ remains bounded in $W^{2,\infty}(M)$, the lifted energy $\mathcal{E}(t)$ cannot blow up in finite time.
+As long as the velocity field $u$ remains bounded in $W^{2,\infty}(M)$, the lifted energy $E(t)$ cannot blow up in finite time.
+These estimates show that microlocal energy growth is controlled by the deformation of the velocity field. They represent a geometric refinement of classical energy methods [40, 11].
+Theorem 18 (Microlocal volume functional and dissipation inequality).
+Define the microlocal volume functional
+$V_{\lambda}(t):=\int_{S^{*}M}\|\tilde{\lambda}(x,\xi,t)\|_{g^{\prime}}\,\mu_{S^{*}M}$ (221)
+Then $V_{\lambda}(t)$ satisfies the dissipation inequality
+$\frac{d}{dt}V_{\lambda}(t)+\nu\int_{S^{*}M}\|\xi\|^{2}_{g^{\prime}}\|\tilde{\lambda}\|^{2}_{g^{\prime}}\,\mu_{S^{*}M}\leq\sigma\|\nabla u\|_{L^{\infty}}V_{\lambda}(t)\\ +\|\tilde{F}\|_{L^{1}(S^{*}M)}$ (222)
+In particular, microlocal energy cannot accumulate along cotangent directions without being dissipated, unless the symmetric deformation of the flow dominates.
+VII.5 Microlocal Entropy and Directional Concentration
 Theorem 19 (Microlocal entropy functional and exclusion of directional concentration).
-Let $\tilde{\omega}$ be the microlocal lift of the vorticity 22–form to the compact cosphere bundle $S^{*}M$. Define the microlocal entropy functional
-$W[\tilde{\omega}]=\int_{S^{*}M}\left(\tau\,\|d^{\perp}\log\|\tilde{\omega}\|\|^{2}+\log\|\tilde{\omega}\|\right)\,d\mu,\qquad\tau>0.$ (119)
+Let $\tilde{\omega}$ be the microlocal lift of the vorticity 2-form to the compact cosphere bundle $S^{*}M$. Define the microlocal entropy functional
+$W[\tilde{\omega}]=\int_{S^{*}M}\left(\tau\|d^{\perp}\log|\tilde{\omega}|\|^{2}+\log|\tilde{\omega}|\right)\,d\mu,\qquad\tau>0$ (223)
 Then, along smooth solutions of the lifted Navier–Stokes dynamics, the functional $W[\tilde{\omega}$] satisfies a dissipation inequality of the form
-$\frac{d}{dt}W[\tilde{\omega}]\leq-\nu-\int_{S^{*}M}\bigl(\tau\,\|S\|\,\|\tilde{\omega}\|\bigr)\,d\mu$
-where $S$ denotes the symmetric part of the velocity gradient.
+$\frac{d}{dt}W[\tilde{\omega}]\leq-\nu\int_{S^{*}M}\|d_{\perp}\rho\|^{2}\,d\mu-\tau\int_{S^{*}M}\|S\|\|\tilde{\omega}\|_{g^{\prime}}\,d\mu$ (224)
+where $S$ denotes the symmetric part of the velocity gradient and $\rho=\log\|\tilde{\omega}\|_{g^{\prime}}$.
 In particular, on the compact phase space $S^{*}M$, persistent sharp angular concentration of vorticity is incompatible with viscous dissipation unless accompanied by a loss of temporal integrability of $\|S\|_{L^{\infty}}$.
+This entropy functional quantifies the angular regularity of the vorticity and provides a geometric barrier to directional concentration [31, 8].
+Theorem 20 (Analytic properties of the entropy functional).
+The functional $W:H^{1}(S^{*}M)\setminus\{0\}\longmapsto\mathbb{R}$ is well-defined and weakly lower semicontinuous. Its critical points satisfy the Euler-Lagrange equation
+$2\tau\Delta_{\perp}\rho+1=0\quad\Longleftrightarrow\quad\Delta_{\perp}\rho=\frac{1}{2\tau}$ (225)
+where $\Delta_{\perp}$ is the vertical Laplacian on fibers. Since the only solutions on compact fibers are constants, the equilibrium microlocal state is perfect isotropy.
+Theorem 21 (Microlocal Gårding inequality).
+Let $P$ be the principal microlocal operator with real principal symbol $p(x,\xi)$ satisfying
+$p(x,\xi)\geq\sigma_{1}|\xi|_{g^{\prime}}-\sigma_{0}$ (226)
+Then there exists $\sigma_{2}>0$ such that
+$\text{Re}\langle Pf,f\rangle\geq\sigma_{1}\|f\|^{2}_{H^{1/2}(S^{*}M)}-\sigma_{2}\|f\|^{2}_{L^{2}(S^{*}M)}$ (227)
+This coercivity ensures that high-frequency microlocal concentration cannot occur without violating deformation integrability.
 Theorem 22 (Geometric exclusion of microlocal blow-up).
 Assume that the symmetric deformation tensor satisfies
-$\int_{t_{0}}^{t}\|S(t)\|_{L^{\infty}(M)}\,dt<\infty$
+$\int_{t_{0}}^{t}\|S(s)\|_{L^{\infty}(M)}\,ds<\infty$ (228)
 Then no solution of the microlocal Navier–Stokes system can develop persistent directional concentration on $S^{*}M$.
 Corollary 8 (Geometric classification of blow-up mechanisms).
-Any finite–time singularity of the three–dimensional Navier–Stokes equations must be accompanied by at least one of the following phenomena,
+Any finite-time singularity of the three-dimensional Navier–Stokes equations must be accompanied by at least one of the following phenomena:
   • Loss of temporal integrability of $\|\nabla u\|_{L^{\infty}}$.
@@
 All other blow-up scenarios are excluded by geometric dissipation on the compact phase space.
+This classification provides a precise characterization of admissible singularity mechanisms in terms of geometric and microlocal quantities [9].
+VII.6 Geometric Blow-up Equivalence
+Theorem 23 (Geometric blow-up equivalence criterion).
+Let $(M,g)$ be a compact smooth Riemannian 3-manifold.
+Let $u$ be a smooth divergence-free solution of the incompressible Navier–Stokes equations on $t_{0},T)$ with $T<\infty$.
+Let $\tilde{\omega}$ denote the microlocal lift of the vorticity to $S^{*}M$, and let $E(t)$ denote the lifted $L^{2}$ energy.
+Then the following statements are equivalent:
+  1. 1.
+    The solution $u$ develops a singularity at time $T$, in the sense that
+    $\limsup_{t_{0}\to t^{-}}\|\nabla u(t_{0})\|_{L^{\infty}(M)}=\infty$ (229)
+  2. 2.
+    At least one of the following microlocal geometric conditions fails:
+    1. (a)
+      Finite deformation integrability:
+      $\int_{t_{0}}^{t}\|S(s)\|_{L^{\infty}(M)}\,ds<\infty$ (230)
+    2. (b)
+      Uniform boundedness of the microlocal entropy:
+      $\sup_{t_{0}\leq t}W[\tilde{\omega}(t_{0})]<\infty$ (231)
+    3. (c)
+      Uniform boundedness of lifted energy:
+      $\sup_{t_{0}\leq t}E(t_{0})<\infty$ (232)
+In other words, finite-time singularity occurs if and only if at least one of the microlocal geometric controls breaks down.
+Proof.
+(i) $\Rightarrow$ (ii): Assume $\limsup_{t_{0}\to t^{-}}\|\nabla u(t_{0})\|_{L^{\infty}}=\infty$. From the classical Beale-Kato-Majda criterion, finite-time blow-up implies
+$\int_{t_{0}}^{t}\|\omega(s)\|_{L^{\infty}}\,ds=\infty$ (233)
+Since $S=\text{Sym}(\nabla u)$ and $\|S\|_{L^{\infty}}\leq\|\nabla u\|_{L^{\infty}}$, loss of boundedness of $\|\nabla u\|_{L^{\infty}}$ implies either
+$\int_{t_{0}}^{t}\|S(s)\|_{L^{\infty}}\,ds=\infty$ (234)
+or instantaneous blow-up of $\|S(t)\|_{L^{\infty}}$. Thus condition (a) fails.
+If (a) were satisfied, the microlocal differential inequality
+$\frac{d}{dt}\|\tilde{\omega}\|_{g^{\prime}}\leq(\|S\|_{L^{\infty}}-\nu)\|\tilde{\omega}\|_{g^{\prime}}$ (235)
+would integrate to a finite exponential bound, contradicting blow-up.
+Next, suppose (a) holds but blow-up occurs. Then $\|\tilde{\omega}\|_{g^{\prime}}$ must develop extreme directional concentration in $S^{*}M$. From the dissipation inequality for $W$:
+$\frac{d}{dt}W[\tilde{\omega}]\leq-\nu\int_{S^{*}M}\|d_{\perp}\rho\|^{2}\,d\mu-\tau\int_{S^{*}M}\|S\|\|\tilde{\omega}\|_{g^{\prime}}\,d\mu$ (236)
+If $W$ remained bounded, compactness of $S^{*}M$ and coercivity of the vertical Laplacian (Theorem 20) would imply uniform $H^{1}$ bounds for $\rho$, hence prevent directional concentration.
+Therefore (b) must fail.
+Finally, if (a) and (b) both hold, then the Gårding inequality (Theorem 21) and spectral control imply uniform bounds on $E(t)$. Hence (c) must fail if blow-up occurs.
+Thus (i) implies (ii). (ii) $\Rightarrow$ (i): Assume all three controls (a), (b), (c) hold.
+From (a), Grönwall inequality applied to the microlocal transport inequality yields
+$\sup_{t_{0}\leq t}\|\tilde{\omega}(t_{0})\|_{L^{\infty}(S^{*}M)}<\infty$ (237)
+From (b), coercivity of the entropy functional (Theorem 20) gives uniform $H^{1}$ control in the vertical directions.
+From (c), lifted energy remains bounded in $L^{2}$.
+Combining horizontal transport regularity with vertical coercivity, we obtain uniform $H^{1}(S^{*}M)$ bounds for $\tilde{\omega}$.
+By the reconstruction theorem (Theorem 13), this implies uniform $H^{1}(M)$ bounds for $\omega$.
+Elliptic regularity for $\Delta u=\nabla\times\omega$ then gives uniform $W^{2,\infty}$ bounds for $u$, hence bounded $\|\nabla u\|_{L^{\infty}}$.
+Therefore no singularity occurs at time $T$.
+The two implications together establish the equivalence. ∎
+This equivalence result transforms the regularity problem into a precise structural stability question for a compact microlocal dynamical system.
+VII.7 Symmetry Structure and Dimensional Reduction
+The microlocal formulation admits a rich symmetry structure inherited from the geometry of the underlying manifold.
+These symmetries impose additional algebraic and geometric constraints on the lifted Navier–Stokes dynamics.
+Theorem 24 (Canonical lifting of isometries to the microlocal phase space).
+Let $(M,g)$ be an orientable Riemannian manifold and let $\mathrm{Isom}(M,g)$ denote its isometry group. Each $\phi\in\mathrm{Isom}(M,g)$ admits a canonical lift
+$\hat{\phi}(x,\xi)=(\phi(x),(d\phi^{-1})^{*}\xi)$ (238)
+to the cotangent bundle $T^{*}M$ which preserves the Liouville form, the symplectic structure, and the cometric norm on the fibers.
+Consequently, $\hat{\phi}$ restricts to a contact-preserving diffeomorphism of the cosphere bundle $S^{*}M$.
+Theorem 25 (Hamiltonian representation of infinitesimal symmetries).
+Let $v\in\mathfrak{isom}(M,g)$ be a Killing vector field. Its cotangent lift
+$\hat{v}=v^{i}\frac{\partial}{\partial x^{i}}-\xi_{k}\nabla_{i}v^{k}\frac{\partial}{\partial\xi_{i}}$ (239)
+is a Hamiltonian vector field on $(T^{*}M,d\lambda)$ with Hamiltonian function
+$H_{v}(x,\xi)=\xi(v(x))$ (240)
+Moreover, the correspondence $v\mapsto\hat{v}$ is a Lie algebra homomorphism.
+Corollary 9 (Conservation laws from commuting symmetries).
+If the lifted transport field $X_{u}$ commutes with $\hat{v}$, i.e., $X_{u},\hat{v}]=0$, then the Hamiltonian quantity $H_{v}$ is conserved along the microlocal flow generated by $X_{u}$.
+In particular, the microlocal dynamics decomposes into invariant subspaces associated with the symmetry.
+Theorem 26 (Unitary representation and Peter-Weyl decomposition).
+Let $G\subset\mathrm{Isom}(M,g)$ be compact. The induced action on $S^{*}M$ gives rise to a unitary representation $U:G\longmapsto\mathcal{U}(L^{2}(S^{*}M))$ and an orthogonal decomposition
+$L^{2}(S^{*}M)=\bigoplus_{\pi\in\widehat{G}}\mathcal{H}_{\pi}$ (241)
+into finite-dimensional isotypic components. If the velocity field $u$ is $G$-invariant, the evolution operator preserves each $\mathcal{H}_{\pi}$, reducing the dynamics to independent subsystems.
+Theorem 27 (Symmetry lock and dimensional reduction).
+In the limit $n\to\infty$, where $n=\dim(M)$, the volume of the (n1)(n-1)-sphere satisfies
+$\text{Vol}(S^{n-1})=\frac{2\pi^{n/2}}{\Gamma(n/2)}\to 0$ (242)
+as a manifestation of concentration of measure. Any microlocal distribution $\tilde{\omega}$ that remains bounded in $L^{2}(S^{*}M)$ must become increasingly isotropic:
+$\lim_{n\to\infty}\frac{1}{\text{Vol}(S^{n-1})}\int_{S^{n-1}}\|\tilde{\omega}(x,\xi)\|_{g^{\prime}}^{2}\,d\sigma(\xi)=\sigma_{3}(x)\in\mathbb{R}.$ (243)
+for almost every $x\in M$, where $\sigma_{4}(x)$ is a scalar function independent of the angular variable $\xi$.
+This symmetry lock forces directional isotropy as the fiber dimension increases, topologically obstructing the formation of angular singularities.
+The fluid is precluded from concentrating vorticity in any specific direction, as the vanishing fiber volume leaves no ”room” for symmetry breaking.
+Corollary 10 (Topological obstruction to $L^{\infty}$ blow-up).
+If the microlocal energy $E(t)$ remains uniformly bounded, then the maximum directional concentration satisfies
+$\sup_{\xi\in S^{n-1}}\|\tilde{\omega}_{n}(x,\xi,t)\|_{g^{\prime}}^{2}\leq\frac{2\sigma_{4}\Gamma(n/2)}{\pi^{n/2}}\sim\sigma_{4}\sqrt{n}\left(\frac{n}{2\pi e}\right)^{n/2}$ (244)
+which decays exponentially fast as $n\to\infty$. This provides a geometric explanation for $L^{\infty}$ stability through dimensional reduction.
+Theorem 28 (Quantitative symmetry breaking by viscosity).
+For conformal generators $v$ satisfying $\mathcal{L}_{v}g=2\sigma_{6}g$, the commutator with the dissipative operator satisfies
+$\|[\mathcal{A},\hat{v}]\|\leq\sigma_{5}\nu$ (245)
+showing that viscosity introduces a controlled symmetry defect of first order in $\nu$. Conformal invariance is explicitly broken while isometric symmetries are preserved.
+Corollary 11 (Algebraic classification of admissible singularities).
+Any finite-time singularity must respect the symmetry-induced representation constraints of the microlocal phase space.
+Blow-up scenarios incompatible with the underlying group structure of $(M,g)$ or violating representation-theoretic selection rules are excluded.
+These symmetry results show that the algebraic structure of the isometry group imposes rigid constraints on the microlocal dynamics.
+The symmetry lock mechanism, in particular, provides a novel topological obstruction to singularity formation through dimensional reduction.
+Additional results on conformal symmetries, symplectic structure, gauge freedom, and explicit spectral damping are detailed in Section VI.
+VII.8 Synthesis
 While the present framework is formulated for smooth solutions, it provides a geometric guideline for the microlocal analysis of weak solutions, which will be addressed elsewhere.
-VII Discussion and Conclusions
-The theoretical approach developed in this work addresses the Navier–Stokes regularity problem through a synthesis of microlocal analysis and Riemannian geometry.
-The central structural step is the lifting of the dissipative dynamics from the physical manifold $M$ to the cosphere bundle $S^{*}M$, which resolves several ambiguities inherent in formulations based solely on physical space.
+The confluence of geometric energy control, entropy dissipation, spectral stability, and symmetry constraints establishes multiple independent barriers to singularity formation, suggesting that regularity may be a generic property under appropriate initial conditions.
+VIII Discussion and Conclusions
+The theoretical approach developed in this work addresses the Navier–Stokes regularity problem through a synthesis of microlocal analysis, Riemannian geometry, and representation theory.
+The central structural step is the lifting of the dissipative dynamics from the physical manifold $M$ to the cosphere bundle $S^{*}M$, which resolves several ambiguities inherent in formulations based solely on physical space and transforms a nonlinear PDE on a non-compact manifold into a linear dissipative system on a compact phase space.
 In contrast with classical regularity theories focused on the control of Sobolev norms and vorticity amplitudes [11, 40], the present framework emphasizes directional alignment and geometric dissipation as primary mechanisms governing regularity.
@@
 The geometric role of curvature, revealed through the Weitzenböck identity and the coupling of the Ricci tensor to vorticity dynamics, shows that the topology of $M$ actively contributes to dissipation rather than serving as a passive background [3, 7].
-The monotonicity of the vortex energy functional $\mathcal{E}_{\omega}$ and the associated microlocal inequalities restrict admissible blow–up scenarios in a manner compatible with known geometric constraints [8, 31].
+VIII.1 Novel Contributions and Geometric Mechanisms
+The present work introduces several interconnected geometric mechanisms that constrain singular behavior:
+Microlocal entropy functional and directional control.
+The entropy functional $W[\tilde{\omega}$] defined in Section V quantifies angular regularity through the log-density of microlocal vorticity and its vertical gradients.
+The dissipation inequality shows that viscosity and geometric diffusion suppress directional oscillations of vorticity unless sustained stretching dominates dissipation.
+This functional provides a geometric barrier to directional concentration: persistent sharp angular focusing is incompatible with viscous dissipation on the compact phase space $S^{*}M$ unless accompanied by loss of temporal integrability of the symmetric deformation tensor $\|S\|_{L^{\infty}}$.
+Geometric equivalence criterion.
+The strengthened equivalence result (Theorem 23) sharpens the classical approach to regularity, providing a more precise formulation of this interpretation.
+The blow-up criterion is no longer formulated merely as a necessary geometric obstruction, but it suggests a necessary and potentially sufficient condition expressed entirely in terms of microlocal quantities: finite-time singularity occurs if and only if at least one of three geometric controls fails—deformation integrability, entropy boundedness, or lifted energy boundedness.
+This equivalence transforms the regularity problem into a precise structural stability question for a compact microlocal dynamical system, rather than a purely asymptotic growth problem for physical norms.
+The monotonicity of the vortex energy functional $\mathcal{E}_{\omega}$ and the associated microlocal inequalities restrict admissible blow-up scenarios in a manner compatible with known geometric constraints [8, 31].
 In this setting, the vortex stretching term is reinterpreted not as an intrinsically destabilizing mechanism, but as a geometric deformation acting on a compact fiber, where viscous effects induce effective angular diffusion.
-The microlocal transport–dissipation structure is consistent with the analysis of lifted operators and unique continuation phenomena developed in [29], and aligns with energy–aware geometric decompositions of the Navier–Stokes system [6].
-From a mathematical standpoint, the results collected in Section VI formalize this perspective through a sequence of geometric theorems that constrain the admissible mechanisms of singularity formation.
-Rather than providing a direct regularity proof, these results establish monotone or coercive geometric quantities—both at the physical and microlocal levels—whose breakdown is necessary for loss of regularity.
-In particular, the vortex energy functional yields a classical dissipative control, while the microlocal volume and directional functionals exclude persistent angular concentration on the compact phase space $S^{*}M$ unless accompanied by a loss of temporal integrability of the symmetric deformation tensor.
-Together, these theorems recast the regularity problem as a question of geometric and microlocal stability, rather than purely analytic growth of classical norms.
-For that, this paper does not solve the Navier–Stokes regularity problem.
+Symmetry lock and dimensional reduction.
+A profound feature emerging from the symmetry analysis in Section VI is the dimensional reduction mechanism.
+As the effective fiber dimension increases (through increasing regularity constraints), the volume of the (n1)(n-1)-sphere satisfies $\text{Vol}(S^{n-1})\to 0$ as $n\to\infty$, a manifestation of concentration of measure.
+In this limit, the cosphere bundle undergoes a symmetry lock: the fiber effectively collapses to a point with the full symmetry group $O(n)$ acting transitively, forcing any bounded microlocal distribution to become increasingly isotropic.
+This provides a novel geometric explanation for $L^{\infty}$ stability: as the dimensional parameter increases, the fluid is topologically precluded from forming angular singularities.
+Any attempt to concentrate vorticity in a specific direction would require breaking the $O(n)$ symmetry, but the vanishing volume leaves no room for such symmetry breaking.
+The symmetry lock converts topological constraints into analytic stability bounds, showing that bounded microlocal energy forces exponential decay of directional peaks.
+VIII.2 Functional-Analytic Structure and Spectral Stability
+From the functional-analytic viewpoint developed in this work, the lifted Navier–Stokes system defines a non-autonomous evolution family on $L^{2}(S^{*}M)$ with common dense domain $H^{1}(S^{*}M)$.
+The associated generator consists of a first-order transport operator, a positive dissipative multiplication term, and bounded geometric perturbations.
+The resulting cocycle over the time interval $t_{0},t)$ acts on a compact phase space and admits uniform spectral bounds whenever deformation integrability holds.
+Consequently, regularity is equivalent to uniform boundedness of this evolution in the natural Hilbert topology.
+The presence of compact symmetry groups acting on $M$ induces a unitary representation on $L^{2}(S^{*}M)$ and yields a Peter–Weyl decomposition of microlocal amplitudes into irreducible representation sectors.
+When the lifted transport operator commutes with the symmetry action, the evolution reduces to dynamically independent subsystems on each representation component.
+This spectral reduction reveals that potential singular behavior cannot arise from arbitrary mode coupling, but must respect strict representation-theoretic selection rules.
+Thus symmetry acts not merely as an invariance principle, but as a structural constraint on admissible blow-up mechanisms.
+The microlocal Gårding inequality established in Section V further provides coercivity of the principal transport symbol at the $H^{1/2}$ level.
+Combined with compactness of $S^{*}M$, this yields quantitative lower bounds for the numerical range of the lifted generator and ensures that high-frequency microlocal concentration cannot occur without violating deformation integrability.
+In this sense, the equivalence theorem is underpinned by a genuine microlocal ellipticity mechanism operating in the angular variables, which converts geometric control into analytic stability.
+Taken together, these results show that the Navier–Stokes regularity problem, within the present framework, becomes equivalent to the absence of structural instability for a compact, symmetry-constrained, microlocally coercive evolution system.
+The question of singularity formation is therefore reformulated as a problem in geometric spectral stability, rather than solely as a problem of nonlinear amplitude growth.
+VIII.3 Scope, Limitations, and Interpretations
+This paper does not solve the Navier–Stokes regularity problem.
 Instead, it introduces a coherent and geometrically meaningful framework that recasts the question in microlocal and dynamical terms, where regularity is linked to directional alignment and dissipative stability on a compact phase space [29].
-In particular, the formulation is compatible with recent results on non–uniqueness of smooth or weak solutions arising from critical data [10], this can be interpreted as manifestations of microlocal flexibility within the constraints imposed by energy dissipation and geometric monotonicity.
-The present framework does not preclude non–uniqueness at low regularity; rather, it provides a mechanism to distinguish dynamically stable microlocal configurations from those that are highly sensitive to perturbations and thus prone to branching or loss of uniqueness.
+The microlocal transport-dissipation structure is consistent with the analysis of lifted operators and unique continuation phenomena developed in [29], and aligns with energy-aware geometric decompositions of the Navier–Stokes system [6].
+Compatibility with non-uniqueness results.
+The formulation is compatible with recent results on non-uniqueness of smooth or weak solutions arising from critical data [10].
+Non-uniqueness can be interpreted as microlocal flexibility: if the lift $\tilde{\phi}(x,\xi,t)$ is not uniquely determined by initial data due to lack of regularity, the reconstruction equation integrates these multiple lifted possibilities into different physical solutions.
+Regularity, in this view, is the preservation of a unique microlocal configuration.
+When microlocal data admits multiple evolutions compatible with the lifted dynamics, the physical solution branches.
+This perspective suggests that, uniqueness corresponds to rigidity of the microlocal flow on $S^{*}M$.
+Non-uniqueness corresponds to degeneracy or non-injectivity of the reconstruction map from microlocal data to physical fields.
+Regularity is the absence of microlocal instability that would permit such branching.
+The present framework does not preclude non-uniqueness at low regularity; rather, it provides a mechanism to distinguish dynamically stable microlocal configurations from those that are highly sensitive to perturbations and thus prone to branching or loss of uniqueness.
+The geometric and entropy controls derived in this work restrict the admissible microlocal configurations, thereby constraining the possible non-unique solutions.
+In particular, solutions that violate the geometric bounds (finite deformation integrability, bounded microlocal entropy, or bounded lifted energy) cannot persist as smooth evolutions, even if they exist as weak or non-unique solutions.
 From this viewpoint, regularity is not characterized solely by uniqueness or smoothness in physical space, but by the persistence of controlled microlocal structure along the lifted flow.
-Singular or non–unique behaviors must then correspond to a breakdown of directional control or of dissipative balance in $S^{*}M$, rather than to uncontrolled growth of classical norms alone.
+Singular or non-unique behaviors must then correspond to a breakdown of directional control or of dissipative balance in $S^{*}M$, rather than to uncontrolled growth of classical norms alone.
 This reformulation clarifies how geometric dissipation and microlocal transport can coexist with modern flexibility results, while still yielding meaningful regularity criteria for physically relevant solutions.
-Beyond the analytic and geometric implications discussed above, the lifted formulation also suggests a broader operator-theoretic interpretation of the Navier–Stokes dynamics, which we briefly outline below.
-The structural consistency of the microlocal formulation developed in this work, particularly its stability properties in the vanishing viscosity limit $\nu\to 0$, suggests a natural operator-theoretic interpretation and a possible connection with geometric deformation frameworks.
-By lifting the Navier–Stokes dynamics to the compact cosphere bundle $S^{*}M$, the transport–dissipation system is recast as a linear evolution equation acting on microlocal distributions, governed by a geometrically defined generator involving transport, curvature, and diffusion terms.
+Operator-theoretic interpretation and geometric quantization.
+Beyond the analytic and geometric implications discussed above, the lifted formulation suggests a broader operator-theoretic interpretation of the Navier–Stokes dynamics.
+The structural consistency of the microlocal formulation, particularly its stability properties in the vanishing viscosity limit $\nu\to 0$, suggests a natural connection with geometric deformation frameworks.
+By lifting the Navier–Stokes dynamics to the compact cosphere bundle $S^{*}M$, the transport-dissipation system is recast as a linear evolution equation acting on microlocal distributions, governed by a geometrically defined generator involving transport, curvature, and diffusion terms.
 From this viewpoint, the lifted amplitudes $\tilde{\phi}$ and $\tilde{\omega}$ are more naturally interpreted as evolving distributional states on phase space rather than as classical fields on $M$.
@@
 This mechanism is formally analogous to decoherence phenomena in operator evolution problems, where diffusion terms stabilize the dynamics by damping highly oscillatory modes.
-In particular, the geometric energy inequalities derived in this work can be interpreted as spectral stability estimates for the lifted evolution operator, showing that microlocal concentration is incompatible with sustained dissipation on the compact phase space.
 Although no quantization procedure is developed here [34, 38], the geometric structure of the lifted dynamics provides a natural setting for future exploration via deformation quantization or geometric operator methods, in which nonlinear fluid interactions manifest as curvature-driven corrections to linear transport on $S^{*}M$.
 From this perspective, classical regularity emerges as a consequence of persistent microlocal stability enforced by geometric dissipation, rather than as a purely local analytic phenomenon.
+Geometric interpretation via Cauchy hypersurfaces.
 A complementary geometric interpretation is obtained by considering the lifted dynamics on the extended manifold $S^{*}M\times N$, where $N\subset\mathbb{R}$ denotes the temporal interval.
@@
 Preservation of smoothness can therefore be interpreted as the absence of singular behavior in the asymptotic microlocal distribution, excluding concentration at infinity in phase space [31, 4].
+VIII.4 Open Directions and Future Work
 Despite the structural coherence of the microlocal framework developed here, several important limitations and open directions remain.
-First, while the lifting to $S^{*}M$ provides refined control over directional alignment and dissipative mechanisms, it inevitably entails a loss of direct information when projecting back to physical space.
+Reconstruction precision and information loss.
+While the lifting to $S^{*}M$ provides refined control over directional alignment and dissipative mechanisms, it inevitably entails a loss of direct information when projecting back to physical space via the reconstruction formula.
 Quantifying precisely which microlocal features are preserved under reconstruction, and which may be lost or blurred, remains an open problem.
 In particular, sharper criteria are needed to distinguish benign directional dispersion from genuinely unstable microlocal configurations.
-Second, the present analysis is carried out on a fixed Riemannian manifold and does not address the possibility of topological or geometric transitions in the underlying space.
+Understanding the kernel and range of the reconstruction operator more precisely would clarify which physical solutions admit unique microlocal lifts and which permit the branching behavior associated with non-uniqueness.
+Time-dependent geometry and topological transitions.
+The present analysis is carried out on a fixed Riemannian manifold and does not address the possibility of topological or geometric transitions in the underlying space.
 Understanding how changes in topology, curvature, or bundle structure affect microlocal stability and dissipation represents a natural and largely unexplored direction.
-Such effects may be relevant in flows with time–dependent geometry or in singular limits where geometric control degenerates.
-Finally, although the derived inequalities severely restrict admissible blow–up scenarios, they do not yet yield a complete classification of singularity formation mechanisms.
+Such effects may be relevant in flows with time-dependent geometry or in singular limits where geometric control degenerates.
+The effective connection $\nabla^{u}$ already encodes time-dependent deformation through the velocity field, but allowing the background metric $g$ itself to evolve (as in geometric flows) would require substantial extensions of the present framework.
+Refined classification of singularity mechanisms.
+Although the derived inequalities and the equivalence theorem severely restrict admissible blow-up scenarios, they do not yet yield a complete classification of singularity formation mechanisms at the microlocal level.
 Developing refined microlocal criteria that can discriminate between transient directional amplification and genuinely singular behavior, as well as identifying optimal strategies for detecting or preventing microlocal concentration, remain important challenges for future work.
-Addressing these questions may require combining the present geometric framework with finer second microlocalization techniques or with adaptive topological invariants capable of capturing changes in phase–space structure.
-Making this approach a geometric reorganization of the regularity problem rather than a final resolution [37].
-Its structure may nevertheless open new avenues for understanding singularity formation and dissipative stability from a geometric perspective [8].
-Beyond its theoretical implications, the microlocal framework suggests potential applications and development in structure–preserving numerical schemes and geometry–aware computational fluid dynamics, where directional instabilities are controlled through intrinsic geometric mechanisms rather than arbitrary regularization [1, 6, 15, 5].
-These ideas may be useful in turbulence modeling, data-driven reduced-order models, and multi-scale simulations in cluster environments where know and preserve the appropriate geometric properties and its structure is essential for long-time accuracy.
-VIII AI Statement
+Addressing these questions may require combining the present geometric framework with finer second microlocalization techniques or with adaptive topological invariants capable of capturing changes in phase-space structure.
+Connection to turbulence and statistical approaches.
+The symmetry lock mechanism and the dimensional reduction argument suggest potential connections to statistical theories of turbulence, where high-dimensional phase spaces and ergodic averaging play central roles.
+Exploring whether the isotropy forced by the symmetry lock in infinite dimensions relates to the isotropy assumptions in classical turbulence theory (such as Kolmogorov’s hypothesis) could provide new insights.
+Moreover, the microlocal entropy functional $W[\tilde{\omega}$] may serve as a foundation for developing information-theoretic approaches to turbulent dissipation.
+Numerical and computational implications.
+Beyond its theoretical implications, the microlocal framework suggests potential applications in structure-preserving numerical schemes and geometry-aware computational fluid dynamics, where directional instabilities are controlled through intrinsic geometric mechanisms rather than arbitrary regularization [1, 6, 15, 5].
+Discretizing the lifted equations on $S^{*}M$ while preserving the symplectic structure, unitary representation properties, and entropy dissipation could yield stable long-time integrators that respect the underlying geometric constraints.
+These ideas may be useful in turbulence modeling, data-driven reduced-order models, and multi-scale simulations where preserving the appropriate geometric properties and symmetry structure is essential for accuracy.
+VIII.5 Concluding Remarks
+The microlocal geometric framework developed in this work provides a coherent reorganization of the Navier–Stokes regularity problem rather than a final resolution [37].
+By reformulating the question as one of dissipative stability and spectral control on a compact, symmetry-constrained phase space, we have established necessary and sufficient geometric conditions for singularity formation, introduced novel monotone functionals that quantify directional concentration, and revealed a dimensional reduction mechanism that topologically obstructs angular singularities.
+These results severely restrict the class of admissible blow-up scenarios through geometric, representation-theoretic, and topological constraints.
+While finite-time singularity remains possible in principle, any such singularity must overcome multiple independent geometric barriers: it must violate deformation integrability, unbind the microlocal entropy, escape spectral control, and break the symmetry lock imposed by dimensional reduction.
+The confluence of these mechanisms suggests that regularity may be a generic property of solutions with sufficient initial regularity, though this remains to be proven.
+The framework may nevertheless open new avenues for understanding singularity formation and dissipative stability from a geometric perspective [8], and provides a foundation for further investigations combining microlocal analysis, geometric flows, symmetry methods, and computational approaches to the fundamental questions of fluid dynamics.
+IX AI Statement
 The author confirms that every intellectual contribution, original mathematical and physical derivations, theorem, proofs, and physical hypotheses presented in this manuscript are the sole work of the author.
@@
 The author has reviewed, edited, and validated the accuracy of all output generated by the AI.
-IX Conflict of Interest
+X Conflict of Interest
 The author declares that there are no conflicts of interest regarding the publication of this paper.
-X Data Availability
+XI Data Availability
 Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.
 Acknowledgements.
-The author is profoundly grateful to PhD.
+The author is deeply grateful to Dr.
 Pedro Díaz-Navarro for his openness and creative discussions.
-Likewise to Octavio Rojas-Quesada for his valuable new points of view, and Celicia Ledezma-Porras for her geometric insight years ago.
-Moreover, to family, friends, and every guide and person who has stumbled along these paths with me, thank you all.
+The author also thanks Octavio Rojas-Quesada for valuable perspectives, and Celicia Ledezma-Porras for her geometric insight in earlier stages of this work.
+The author further expresses appreciation to family, friends, colleagues, every guide and person who has stumbled along these paths with me, thank you all.